### v3

# A Framework for Lorentz-Dirac Dynamics and Post-Newtonian Interaction of Radiating Point Charges

Abstract
We examine classical radiation reaction by combining the covariant Lorentz–Dirac formulation, its Landau–Lifshitz (LL) order reduction, and a post-Newtonian (PN) Hamiltonian treatment of interacting and radiating charges. After reviewing the LL reduction and its removal of runaway and preacceleration behavior, we verify energy balance in several relativistic single-particle scenarios by demonstrating agreement between the LL Larmor power and the loss of mechanical energy. We then construct an N-body framework based on the conservative Darwin Hamiltonian supplemented with the leading 1.5PN radiation–reaction term. Numerical simulations of charge-neutral binary systems of both symmetric and asymmetric mass configurations show orbital decay, circularization, and monotonic Hamiltonian decrease consistent with dipole radiative losses. The resulting framework provides a simple analogue of gravitational PN radiation reaction and a tractable system for studying dissipative and potentially chaotic electromagnetic dynamics.
I Introduction
The motion of an accelerating charge is influenced not only by externally applied electromagnetic fields, but also by the particle’s own field. The resulting radiation–reaction (self–force) problem has a century–long history and remains central to the modeling of high–energy astrophysical systems in which radiation losses shape the dynamics of relativistic plasmas. In pulsar and magnetar magnetospheres, black–hole magnetospheres, and gamma–ray burst outflows, charged particles are routinely accelerated to ultra–relativistic energies and radiate copiously through curvature, synchrotron, and inverse–Compton processes, thereby regulating both particle distributions and observed spectra [1, 2, 3, 4]. A systematic classical description of radiation reaction is also valuable as an analogue system for gravitational self–force and waveform modeling in compact binaries, where back–reaction of emitted gravitational waves drives inspiral and determines the phase evolution of signals observed by ground– and space–based detectors [5, 6].
In relativistic electrodynamics the Lorentz–Dirac (LD) equation supplements the Lorentz force with a covariant self–force term linear in the proper–time derivative of the acceleration, capturing the leading effect of radiation back–reaction on the worldline of a point charge [7, 8]. The LD equation admits complementary derivations: a decomposition of the electromagnetic field into half–retarded minus half–advanced parts that isolates its radiative component [10, 9], and Dirac’s world–tube momentum–balance argument based on the electromagnetic stress–energy tensor [7, 8]. Both approaches emphasize the split between bound and radiative degrees of freedom and lead to the same covariant structure. However, the LD equation is third order in time and exhibits well–known pathologies such as runaway solutions and preacceleration, reflecting the need to prescribe initial acceleration data in addition to position and velocity [11, 12, 13]. These issues are resolved by the reduction–of–order prescription of Landau and Lifshitz, which treats the self–force as a perturbation and replaces the derivative of the acceleration by the derivative of the external Lorentz force. The resulting Landau–Lifshitz (LL) equation is second order, causal, and has been justified rigorously in modern self–force treatments of extended charged bodies in the point–particle limit [14, 15, 16].
Although historically motivated by internal consistency of classical electrodynamics, radiation-reaction dynamics has acquired renewed practical importance. In high–energy astrophysics, self–forces govern the damping of particle orbits in pulsar winds, regulate curvature and synchrotron emission in strongly magnetized neutron stars, and influence nonthermal spectra in relativistic jets and gamma–ray bursts [1, 2, 3, 4]. In parallel, ultra-intense laser facilities now access electromagnetic fields comparable to those near compact objects, enabling laboratory studies of radiation reaction in the collision of relativistic electron beams with petawatt–class laser pulses [17]. Recent experiments have reported direct signatures of radiation–reaction–induced energy loss and high–energy photon emission, in good agreement with classical and semiclassical models based on LL–type dynamics [18, 19]. In this work we view such laser applications as a valuable secondary arena, while placing primary emphasis on the role of electromagnetic self–forces as a useful analogue for gravitational radiation reaction in high energy astrophysical scenarios.
The gravitational–wave community has developed a sophisticated post–Newtonian (PN), self–force, and effective–one–body (EOB) formalism to describe the conservative and dissipative dynamics of compact binaries. In PN and ADM Hamiltonian formulations, the conservative sector is encoded in a canonical Hamiltonian known through at least 3PN–4PN order [20, 21, 22], while dissipative radiation–reaction terms first enter at 2.5PN and 3.5PN order [23, 24, 25, 26]. The EOB formalism reorganizes this information into a resummed Hamiltonian and a nonconservative force added to the momentum equation, calibrated to PN fluxes and numerical–relativity waveforms [27, 28, 29, 30, 31, 32]. Effective field theory (EFT) approaches further show that such radiation–reaction forces arise naturally from a doubled–field nonconservative action principle [33, 34]. An electromagnetic analogue that mirrors the PN/EOB “Hamiltonian plus radiation–reaction” framework can provide useful intuition and a testbed for ideas in gravitational–wave theory [5, 6, 16]. While radiation reaction, Landau–Lifshitz dynamics, and post-Newtonian expansions have each been studied extensively, an explicit Hamiltonian-plus-radiation-reaction framework for interacting charged particles in electrodynamics—directly analogous to PN/EOB treatments of gravitational binaries—has remained largely unexplored.
In this paper we construct such an analogue by combining the covariant LD/LL description with a PN Hamiltonian treatment of interacting charges. We first revisit the LD and LL equations, giving concise derivations based on the bound/radiative field split and on Dirac’s world–tube momentum balance, and we review the appearance of runaway and preaccelerated solutions together with their removal by reduction of order [7, 10, 9, 8, 14, 15]. We then implement the LL equation in several relativistic single–particle scenarios (cyclotron motion and three representative external–field configurations) and verify energy balance by comparing the change in relativistic kinetic energy, the work done by external fields, and the covariant Larmor power. Motivated by PN and EOB literature, we subsequently construct an NN–body framework in which the conservative dynamics are generated by the Darwin Hamiltonian at 1PN order, while the leading dipole radiation–reaction effects enter as a nonconservative 1.5PN force in the momentum equation, derived via order reduction in close analogy with gravitational PN radiation reaction [23, 24, 26, 33]. Numerical simulations of charge neutral binary systems of equal–mass and asymmetric mass systems then demonstrate orbital decay, circularization, trajectory deformation, and monotonic Hamiltonian decrease consistent with dipole Larmor losses. Taken together, these results provide a useful electromagnetic analogue of PN gravitational radiation reaction and a tractable system for exploring dissipative and potentially chaotic dynamics relevant to high–energy astrophysical plasmas, gravitational–wave source modeling, and strong–field laser–plasma experiments.
Organization of the paper.
The structure of this paper is as follows. In Sec. I we introduce the radiation–reaction problem in classical electrodynamics, Section II reviews the Lorentz–Dirac equation and its formulation in terms of bound and radiative field components, while Sec. III presents the Landau–Lifshitz reduction–of–order procedure that yields a causal, second–order evolution equation. In Sec. IV we validate this reduced dynamics through numerical simulations of single–particle motion in static, time–dependent, and multidimensional electromagnetic fields. Section V introduces a post–Newtonian truncation of the Lorentz–Dirac force suitable for multi–particle systems, and Sec. VI develops the combined Darwin (1PN) conservative Hamiltonian and the leading 1.51.5PN dipole radiation–reaction term, together with the corresponding Hamilton equations. In Sec. VII, we present numerical simulations of binary charge-neutral systems subject to both long–range Coulomb interactions and dissipative radiation reaction, illustrating orbital decay, circularization, deformation, and the onset of non–integrable dynamics. We summarize the main results and outline future directions in Sec. VIII. Several technical details are collected in the Appendices: App. A reviews the runaway solutions of the Lorentz–Dirac equation, App. B summarizes electromagnetic field–tensor and metric conventions, App. C provides explicit expressions for the Darwin 1PN corrections, and App. D gives a more detailed discussion of the four–acceleration and the covariant Larmor power.
II Radiation Reaction and the Modified Lorentz–Dirac Equation
The Lorentz–Dirac equation provides the classical expression for the radiation-reaction force on a point charge and admits two standard derivations. The first, following Landau and Lifshitz [10, 8], identifies the self-field as the half-retarded minus half-advanced solution, isolating the radiative component of the electromagnetic field. The second, due to Dirac [7] and developed further in [9], derives the same equation from energy–momentum conservation in a world-tube surrounding the charge. Both approaches lead to the covariant Lorentz–Dirac equation,
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta},$ (1)
which contains third-order derivatives of the worldline. Here $F^{\alpha}_{\mathrm{ext}}$ denotes an externally applied (non-self) force acting on the particle. The term in parentheses, $\left(\delta^{\alpha}_{\beta}+\dfrac{u^{\alpha}u_{\beta}}{c^{2}}\right)$, acts as a projection operator orthogonal to the four–velocity. It ensures that the self–force satisfies the constraint $u_{\alpha}a^{\alpha}=0$, so that the total four–acceleration remains orthogonal to $u^{\alpha}$, consistent with the normalization condition $u_{\alpha}u^{\alpha}=-c^{2}$. As is well known, Eq. (1) admits unphysical runaway and preaccelerated solutions. To remedy this, Landau and Lifshitz proposed a reduction-of-order procedure in which $\dot{a}^{\alpha}$ is replaced by the proper-time variation of the external force, yielding the physically meaningful equation of motion
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}mc^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)F^{\beta}_{\mathrm{ext},\gamma}u^{\gamma},$ (2)
free of runaway solutions and preacceleration, and now regarded as the correct classical description of radiation reaction [13, 12, 11]. Before delving into the dynamics prescribed by the above equation, let us have a brief recap on the Dirac-Lorentz equation based on the Landau-Lifshitz approach and next, the Dirac approach.
II.1 Landau Lifshitz approach
The electromagnetic field generated by an accelerating point charge acts back upon the charge itself, producing what is known as the radiation-reaction or self-force. The classical and systematic approach to deriving this effect, following Landau and Lifshitz [10, 8] begins by decomposing the electromagnetic potential into retarded and advanced solutions of the wave equation,
$\square A^{\alpha}=-4\pi j^{\alpha},$ (3)
where $j^{\alpha}$ is the current of the point charge moving along a worldline $z^{\alpha}(\tau)$. Let $A^{\alpha}_{\mathrm{ret}}$ denote the physically causal field and $A^{\alpha}_{\mathrm{adv}}$ its time-reversed counterpart. Their symmetric and antisymmetric combinations,
$A^{\alpha}_{\mathrm{sym}}=\frac{1}{2}\left(A^{\alpha}_{\mathrm{ret}}+A^{\alpha}_{\mathrm{adv}}\right),\qquad A^{\alpha}_{\mathrm{rr}}=\frac{1}{2}\left(A^{\alpha}_{\mathrm{ret}}-A^{\alpha}_{\mathrm{adv}}\right),$ (4)
have distinct physical interpretations [7, 9]. The symmetric part represents the Coulomb-like near field bound to the particle, while the antisymmetric combination
$A^{\alpha}_{\mathrm{rr}}$ (5)
contains the radiative information responsible for energy loss and self-force. The corresponding electromagnetic field tensor is
$F^{\alpha\beta}_{\mathrm{rr}}=\partial^{\alpha}A^{\beta}_{\mathrm{rr}}-\partial^{\beta}A^{\alpha}_{\mathrm{rr}}.$ (6)
In the momentarily comoving Lorentz frame (MCLF), where the charged particle is instantaneously at rest, one finds the leading contribution to the radiation-reaction field,
$\bm{E}_{\mathrm{rr}}=\frac{2}{3}\,\frac{q}{4\pi\varepsilon_{0}c^{3}}\,\dot{\bm{a}},\qquad\bm{B}_{\mathrm{rr}}=\bm{0},$ (7)
where $\bm{a}$ is the instantaneous acceleration and $\dot{\bm{a}}$ its proper-time derivative [11, 12]. Substitution into the Lorentz force law gives the familiar nonrelativistic expression,
$\bm{F}_{\mathrm{rr}}=\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\dot{\bm{a}}.$ (8)
The covariant generalization of (8), uniquely consistent with $u_{\alpha}a^{\alpha}=0$, is
$F^{\alpha}_{\mathrm{rr}}=\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta}.$ (9)
Including an external force $F^{\alpha}_{\mathrm{ext}}$, the full equation of motion becomes
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta},$ (10)
which is the Lorentz–Dirac equation. As is well known, Eq. (10) involves third derivatives of the position and admits unphysical runaway and pre-accelerated solutions. A consistent physical formulation requires the reduction-of-order procedure of Landau and Lifshitz [10, 13], as we will discuss in Sec. III.
II.2 Dirac’s Approach
A second and conceptually distinct route to the Lorentz–Dirac equation is due to Dirac [7], based on the conservation of total energy–momentum for the combined system of particle and electromagnetic field. In this formulation, one analyzes the flux of electromagnetic stress–energy through a narrow world tube enclosing the particle’s worldline, and equates this to the change in the particle’s mechanical momentum. Modern geometric expositions of this method are given in [9, 8].
II.2.1 World-Tube Construction and Momentum Balance
Let $T^{\alpha\beta}_{\mathrm{em}}$ denote the electromagnetic stress–energy tensor, and surround the particle’s worldline $z^{\alpha}(\tau)$ with a world tube $\Sigma$ of small radius $r$. The flow of electromagnetic four-momentum across this surface is
$\frac{dP^{\alpha}_{\mathrm{em}}}{d\tau}=\int_{\Sigma}T^{\alpha\beta}_{\mathrm{em}}\,d\Sigma_{\beta},$ (11)
where $d\Sigma^{\beta}$ is the outward-directed hypersurface element. Because $T^{\alpha\beta}_{\mathrm{em}}$ is divergenceless in the vacuum exterior of the worldline, the result is independent of the precise choice of tube.
The retarded electromagnetic field naturally decomposes into a near-field bound component and a radiative component [9]. Evaluating (11) yields
$\frac{dP^{\alpha}_{\mathrm{em}}}{d\tau}=\frac{q^{2}}{8\pi\varepsilon_{0}rc^{2}}\,a^{\alpha}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\frac{a^{2}}{c^{2}}\,u^{\alpha},$ (12)
where $u^{\alpha}$ and $a^{\alpha}=du^{\alpha}/d\tau$ are the four-velocity and four-acceleration of the particle, and $a^{2}=a_{\beta}a^{\beta}$. The first term diverges as $r\to 0$ and is interpreted as contributing to the particle’s inertial mass (electromagnetic mass). The second term, finite and directed along $u^{\alpha}$, corresponds to the radiated energy–momentum.
II.2.2 Mechanical Momentum and Mass Renormalization
The total momentum of the particle-plus-field system must be conserved in the absence of exterbnal forces:
$\frac{d}{d\tau}\!\left(P^{\alpha}_{\mathrm{mech}}+P^{\alpha}_{\mathrm{em}}\right)=F^{\alpha}_{\mathrm{ext}}.$ (13)
The divergent term in (12) is absorbed via mass renormalization,
$m=m_{0}+\frac{q^{2}}{8\pi\varepsilon_{0}rc^{2}}\,$ (14)
and consistency with $u_{\alpha}u^{\alpha}=-1$ requires the mechanical momentum to take the form
$P^{\alpha}_{\mathrm{mech}}=m_{0}u^{\alpha}-\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}\,c^{3}}\,a^{\alpha}.$ (15)
The minus sign in $P^{\alpha}_{\rm mech}$ is fixed by the requirement $u_{\alpha}u^{\alpha}=-c^{2}$: because $\dot{P}^{\alpha}_{\rm em}\propto\frac{2}{3}\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\frac{a^{2}}{c^{2}}u^{\alpha}$, total momentum conservation demands $P_{\rm mech}\!\cdot a=-\frac{2}{3}\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}a^{2}$ [8].
Substituting (12) and (15) into (13) produces the covariant Lorentz–Dirac equation,
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\left(\dot{a}^{\alpha}-\frac{a^{2}u^{\alpha}}{c^{2}}\right),$ (16)
where $F^{\alpha}_{\mathrm{ext}}=qF^{\alpha\beta}_{\mathrm{ext}}u_{\beta}$ is the external Lorentz force. Because the four-acceleration is orthogonal to the four-velocity,
$u_{\alpha}a^{\alpha}=0,$ (17)
differentiating this orthogonality relation with respect to proper time $\tau$ yields
$\frac{d}{d\tau}(u_{\alpha}a^{\alpha})=\dot{a}_{\alpha}u^{\alpha}+a_{\alpha}a^{\alpha}=0.$ (18)
Therefore,
$a^{2}\equiv a_{\alpha}a^{\alpha}=-\,\dot{a}_{\alpha}u^{\alpha}.$ (19)
We now use this identity to rewrite the self-force term in Eq. (16). Substituting into Eq. (16) and projecting orthogonally to the four-velocity gives the standard covariant Lorentz–Dirac equation,
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta}.$ (20)
Equation (20) is third order in the particle trajectory, since it contains $\dot{a}^{\alpha}=da^{\alpha}/d\tau$. This leads to runaway and preaccelerated solutions. The physically meaningful equation governing radiation reaction is obtained by applying the reduction-of-order prescription of Landau and Lifshitz [10, 13], replacing $\dot{a}^{\alpha}$ by the proper-time variation of the external field.
III Resolution via Reduction of Order
The Lorentz–Dirac equation,
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta},$ (21)
is third order in the worldline $z^{\alpha}(\tau)$, since it involves the derivative of the acceleration $\dot{a}^{\alpha}$. The requirement to specify initial position, velocity, and acceleration leads to well-known difficulties: runaway solutions and preacceleration. These features are already clearly visible in the nonrelativistic limit [8]. In the nonrelativistic regime, Eq. (21) reduces to
$a(t)-t_{0}\,\dot{a}(t)=\frac{1}{m}F_{\mathrm{ext}}(t),\qquad t_{0}\equiv\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}m}\$(22)
Considering a force that is switched on abruptly and held constant,
$F_{\mathrm{ext}}(t)=f\,\theta(t),$ (23)
where $\theta(t)$ is the Heaviside step function. Solving Eq. (22) gives
$a(t)=e^{t/t_{0}}\!\left[b-\frac{f}{m}\bigl(1-e^{-t/t_{0}}\bigr)\theta(t)\right].$ (24)
For a generic constant $b$, the exponential factor dominates for $t\gg t_{0}$,
$a(t)\sim b\,e^{t/t_{0}},$ (25)
so that even in the presence of a constant applied force the acceleration grows without bound. These unphysical runaway solutions occur whenever the third-derivative term is retained. They can be eliminated only by choosing the constant
$b=\frac{f}{m},$
which removes the runaway term in (24). With this choice, we find
$a(t)=\frac{f}{m}\!\left[\theta(-t)\,e^{t/t_{0}}+\theta(t)\right].$ (26)
Equation (26) shows that the particle begins to accelerate before the force is applied. Although $F_{\mathrm{ext}}(t)=0$ for $t<0$, the acceleration is nonzero for $t<0$:
$a(t)=\frac{f}{m}e^{t/t_{0}},\qquad t<0.$
This phenomenon is known as preacceleration. The particle “anticipates” the force a time $t_{0}$ in advance. For an electron,
$t_{0}=\frac{2}{3}\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}m}\approx 6\times 10^{-24}\,\mathrm{s},$
so preacceleration occurs only on extremely small timescales; nevertheless, its appearance indicates a conceptual inconsistency in treating the charge as a strict point particle.
These difficulties are resolved by treating the radiation-reaction term as a small correction and performing reduction of order [10, 13]. This yields the modified Dirac-Lorentz equations, which is second order, causal, and free of runaway behaviour. Equation (20) is third order in time and possesses unphysical runaway and preaccelerated solutions. Treating $q^{2}/m$ as small and replacing $\dot{a}^{\alpha}$ using the leading-order equation of motion gives
$ma^{\alpha}=F^{\alpha}_{\text{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}m}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)F^{\beta}_{\text{ext},\gamma}u^{\gamma},$ (27)
the Landau–Lifshitz form, which is second-order and free of pathologies. We first consider
$ma^{\alpha}$ $=F^{\alpha}_{\rm ext}+t_{0}\,(\delta^{\alpha}_{\ \beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}})\,F_{\rm ext}^{\beta}{}_{,\gamma}\,u^{\gamma},$ (28)
where $u^{\alpha}$ is the four-velocity, $a^{\alpha}=du^{\alpha}/d\tau$ the four-acceleration, and $t_{0}$ the small radiation-reaction timescale. For an electromagnetic external field we have
$F^{\alpha}_{\rm ext}=q\,F_{\rm ext}^{\alpha\mu}\,u_{\mu},$ (29)
with $F_{\rm ext}^{\alpha\mu}$ the field tensor. Computing the derivative of $F_{\rm ext}^{\beta}$ along the worldline:
$F_{\rm ext}^{\beta}{}_{,\gamma}u^{\gamma}$ $=q\left[(\partial_{\gamma}F_{\rm ext}^{\beta\mu})\,u_{\mu}u^{\gamma}+F_{\rm ext}^{\beta\mu}\,(\partial_{\gamma}u_{\mu})\,u^{\gamma}\right].$ (30)
Since $(\partial_{\gamma}u_{\mu})\,u^{\gamma}=du_{\mu}/d\tau=a_{\mu}$, this becomes
$F_{\rm ext}^{\beta}{}_{,\gamma}u^{\gamma}$ $=q\,F_{\rm ext}^{\beta\mu}{}_{,\gamma}\,u_{\mu}u^{\gamma}+q\,F_{\rm ext}^{\beta\mu}\,a_{\mu}.$ (31)
$ma^{\alpha}$ $=q\,F_{\rm ext}^{\alpha\mu}\,u_{\mu}+qt_{0}\,(\delta^{\alpha}{}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}})\,F_{\rm ext}^{\beta\mu}{}_{,\gamma}\,u_{\mu}u^{\gamma}$
$\quad+qt_{0}\,(\delta^{\alpha}{}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}})\,F_{\rm ext}^{\beta\mu}\,a_{\mu}.$ (32)
Inside the $O(t_{0})$ term we replace $a_{\mu}$ by its leading-order value from the Lorentz force:
$a_{\mu}$ $\approx\frac{q}{m}\,F_{{\rm ext}\,\mu\nu}\,u^{\nu}.$ (33)
Thus,
$F_{\rm ext}^{\beta\mu}\,a_{\mu}$ $\approx\frac{q}{m}\,F_{\rm ext}^{\beta\mu}\,F_{{\rm ext}\,\mu\nu}\,u^{\nu}.$ (34)
Putting everything together (one term is zero by antisymmetry):
$ma^{\alpha}$ $=q\,F_{\rm ext}^{\alpha}{}_{\mu}\,u^{\mu}+qt_{0}\,F_{\rm ext}^{\alpha}{}_{\mu,\nu}\,u^{\mu}u^{\nu}$
$\quad+\frac{q^{2}t_{0}}{m}\,(\delta^{\alpha}{}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}})\,F_{\rm ext}^{\beta}{}_{\mu}\,F_{\rm ext}^{\mu}{}_{\nu}\,u^{\nu}.$ (35)
This is the desired reduced-order Landau–Lifshitz form of the Lorentz–Dirac equation and will be a central equation for our single particle simulations of radiation reaction. Expanding it out in full
$ma^{\alpha}=q\,F_{\rm ext}^{\alpha}{}_{\mu}\,u^{\mu}+q\,t_{0}\!\left[F_{\rm ext}^{\alpha}{}_{\mu,\nu}\,u^{\mu}u^{\nu}+\frac{q}{m}\,F_{\rm ext}^{\alpha}{}_{\mu}\,F_{\rm ext}^{\mu}{}_{\nu}\,u^{\nu}+\frac{q}{mc^{2}}\,u^{\alpha}\!\left(u_{\beta}\,F_{\rm ext}^{\beta}{}_{\mu}\,F_{\rm ext}^{\mu}{}_{\nu}\,u^{\nu}\right)\right],$ (36)
where
$t_{0}=\frac{1}{4\pi\epsilon_{0}}\,\frac{2q^{\,2}}{3mc^{3}}.$ (37)
IV Simulation results at single particle level
(a) Trajectory in 3D.
(b) $r(\tau)$
(c) $x(\tau)$
(d) $y(\tau)$
(e) $z(\tau)$
Figure 1: Relativistic trajectory of a charged particle in a uniform magnetic field $\mathbf{B}=B\,\hat{\mathbf{z}}$, computed using the order–reduced Lorentz–Dirac equations. (a) Three–dimensional trajectory with initial (green) position. (b–e) Radius $r(\tau)$ measured from the relativistic guiding center and the spatial components $x(\tau)$, $y(\tau)$, and $z(\tau)$ at select time intervals. The motion shows damped cyclotron oscillations in the xxyy plane, while $z(\tau)$ stays constant for the chosen initial data. Parameters described in the main text.
In this section we numerically solve the order–reduced Lorentz–Dirac equations (Eq. 36) for a single charged particle in a uniform magnetic field, followed by three examples cases of increasing complexity. Unless stated otherwise, we choose the charge, mass, and permittivity to be $q=m=\varepsilon_{0}=1$. The radiation–reaction timescale $\tau_{0}=\frac{1}{4\pi\varepsilon_{0}}\!\left(\frac{2q^{2}}{3mc^{3}}\right)$ is evaluated using $c=30$ for the cyclotron example to enhance the dissipative effects over the integration interval. The particle is initially located at $\mathbf{r}_{0}=(0,0,0)$ with initial three–velocity $\mathbf{v}_{0}=(5,5,0)$, corresponding to a relativistic cyclotron orbit in the xxyy plane. We impose a uniform magnetic field $\mathbf{B}=(0,0,B_{c})$ with $B_{c}=10$, and no external electric field. The dynamical variables evolved are $(t(\tau),x(\tau),y(\tau),z(\tau))$ together with the four–velocity $u^{\mu}(\tau)$, integrated from proper time $\tau=0$ up to $\tau=\tau_{\mathrm{f}}=150$. These parameter choices produce long, slowly decaying cyclotron trajectories that illustrate the radiation–reaction–induced drift and the expected damping of the transverse motion, while the longitudinal motion remains constant due to the chosen initial conditions. As a consistency check, we verified that the instantaneous radiated power predicted by the Larmor formula matches the loss of kinetic energy obtained directly from the numerical solution of the order–reduced equations. Using the expression
$P_{\rm rad}(\tau)=\frac{2}{3}\,\frac{q^{2}}{4\pi\epsilon_{0}\,m^{2}c^{3}}\,\frac{dp_{\mu}}{d\tau}\frac{dp^{\mu}}{d\tau},$
with the proper–acceleration invariant $a^{2}(\tau)=a^{\mu}a_{\mu}$, we computed the total radiated energy $E_{\rm rad}=\int_{\tau_{\min}}^{\tau_{\max}}P_{\rm rad}(\tau)\,d\tau.$ For the parameter choices, the numerical integration yields
$E_{\rm rad}=1.51272,\qquad\Delta E_{\rm kin}=E_{\rm kin}(\tau_{\min})-E_{\rm kin}(\tau_{\max})=1.55527,$
corresponding to a fractional agreement $E_{\rm rad}/\Delta E_{\rm kin}=0.97.$ The small mismatch is expected and arises from finite–step integration error and the fact that radiation reaction is implemented through the order–reduced Landau–Lifshitz form rather than the full Lorentz–Dirac equation. For reference, the radiation–reaction timescale used in the simulation is
$\tau_{0}=\frac{1}{4\pi\varepsilon_{0}}\left(\frac{2q^{2}}{3mc^{3}}\right)=1.96488\times 10^{-6},$
which is much smaller than the dynamical timescale of the cyclotron orbit; this ensures that radiation damping accumulates slowly and provides a good numerical test of energy balance. The characteristic dynamical timescale of the motion is set by the cyclotron period. For a particle of charge $q$ and mass $m$ in a uniform magnetic field of magnitude $B_{c}$, the coordinate–time cyclotron frequency is $\omega_{\rm c}=qB_{c}/\gamma_{0}m$, so that
$T_{\rm c}^{(t)}=\frac{2\pi}{\omega_{\rm c}}=\frac{2\pi\gamma m}{qB_{c}}.$
With our choices $q=m=1$ and $B_{c}=10$ this gives
$T_{\rm c}^{(\tau)}\simeq 0.646534,$
while the corresponding coordinate–time period between successive peaks (as measured from the relativistically corrected center of the cyclotron orbit) is $T_{\rm c}^{(t)}=\gamma_{0}T_{\rm c}^{(\tau)}\simeq 0.646578$ for the initial Lorentz factor $\gamma_{0}\simeq 1.029$. Comparing with the radiation–reaction timescale
$\tau_{0}=\frac{1}{4\pi\varepsilon_{0}}\left(\frac{2q^{2}}{3mc^{3}}\right)=1.96488\times 10^{-6},$
we see that the radiation–reaction timescale is many orders of magnitude smaller than the orbital timescale, so that dissipation accumulates slowly over many revolutions. To quantify the radiative damping of the cyclotron orbit, we extract an effective decay timescale directly from the numerical trajectory. From the solution $x^{\mu}(\tau)$ we compute the instantaneous cyclotron radius $r(\tau)$ relative to the guiding center, and fit the resulting data to the exponential form
$r(\tau)\simeq A\,e^{-\beta\tau}+C,$ (38)
using a least–squares nonlinear model. The damping rate is then identified as $\beta$, giving the numerical timescale
$\tau_{\rm damp}^{\rm sim}\equiv\beta^{-1}.$ (39)
For comparison, the non-relativistic Landau–Lifshitz prediction for the cyclotron damping time,
$\tau_{\rm damp}^{\rm th}=\frac{1}{\tau_{0}\,\omega_{c}^{2}},\qquad\tau_{0}=\frac{2q^{2}}{3mc^{3}(4\pi\varepsilon_{0})},\qquad\omega_{c}=\frac{qB_{c}}{m},$ (40)
provides analytic benchmark. Evaluating these expressions for the parameters of the simulation yields $\tau_{\rm damp}^{\rm sim}\approx 4.40\times 10^{3}$ and $\tau_{\rm damp}^{\rm th}\approx 5.09\times 10^{3}$, in good quantitative agreement given the small-damping regime and the use of the non-relativistic theoretical estimate.
Case 1: Constant transverse $\mathbf{E}$ and $\mathbf{B}$ fields
(a) 3D trajectory
(b) $x(\tau)$
(c) $y(\tau)$
(d) $z(\tau)$
Case 2: Two Oscillating Electric and Magnetic Fields
(e) 3D trajectory
(f) $x(\tau)$
(g) $y(\tau)$
(h) $z(\tau)$
Case 3: Two crossed Electromagnetic Waves
(i) 3D trajectory
(j) $x(\tau)$
(k) $y(\tau)$
(l) $z(\tau)$
Figure 2: More examples: Landau–Lifshitz trajectories for three representative external-field configurations. Each row shows (left to right) the full 3D worldline projection followed by the coordinate evolution $x(\tau)$, $y(\tau)$, and $z(\tau)$.
IV.1 More examples
To illustrate the structure of radiation–reacting worldlines predicted by the Landau–Lifshitz (LL) equation, we integrate the dynamics for three progressively more complex electromagnetic backgrounds, see Fig. 2 and Table 1. In all simulations the four–velocity is initialized relativistically as
$u^{\mu}=(\gamma c,\,\gamma v_{x},\,\gamma v_{y},\,\gamma v_{z}),$
with initial speeds chosen sufficiently small compared to the simulation light speed $c$. For the constant–field configurations we set $(v_{x},v_{y},v_{z})=(0,0,0)$. For the oscillatory and crossed–wave cases, symmetry is broken with small seed velocities
$v_{x}=c/100,\qquad v_{y}=c/1000,\qquad v_{z}=c/10,$
ensuring that the trajectory samples the full spatial structure of the fields. The time–dependent examples use amplitudes
$E_{0}=0.12\,c,\qquad B_{0}=43.$
Different numerical values of $c$ are used for stability in each configuration: $c=30$ for the cyclotron case, $c=5$ for Case 1, and $c=25$ for Cases 2 and 3.
The Cyclotron case (with $c=30$).
We begin with pure cyclotron motion in a uniform magnetic field $\mathbf{B}=(0,0,B_{0})$ and no electric field. This serves as a good test of LL damping in the absence of electric acceleration, and provides a reference point for the energy–balance diagnostics.
Case 1: Constant transverse fields (with $c=5$).
A uniform electric field $\mathbf{E}=(0,0,E_{0})$ together with a constant magnetic field $\mathbf{B}=(1,0,0)$ produces the standard transverse cycloidal trajectory with an $\mathbf{E}\!\times\!\mathbf{B}$ drift. This configuration serves as a test for validating LL damping in static external fields.
Case 2: Oscillating electric and magnetic fields (with $c=25$).
To probe LL radiation reaction in a time–dependent oscillating background, we impose independently oscillating fields
$\mathbf{E}(t)=\{E_{0}\sin[(\mathbf{k}\!\cdot\!\mathbf{r})-\omega t],0,0\},\qquad\mathbf{B}(t)=\{0,B_{0}\sin[(\mathbf{k}\!\cdot\!\mathbf{r})-\omega t],0\},$
with $\omega=c|\mathbf{k}|$. Because the amplitudes are unrelated, this configuration is not a plane electromagnetic wave For the oscillating-field setup we adopt the amplitudes $E_{0}=0.12\,c$ for the electric field and $B_{0}=43$ for the magnetic field.
Case 3: Two crossed electromagnetic waves (with $c=25$).
We next superpose two oscillatory fields with wavevectors $\mathbf{k}_{1}=(0,0,1/25)$ and $\mathbf{k}_{2}=(0,1/25,0)$ and corresponding frequencies $\omega_{1,2}=c|\mathbf{k}_{1,2}|$. Using the same amplitudes $E_{0}$ and $B_{0}$ as in Case 2, the resulting field has genuine three–dimensional interference, producing LL trajectories with multiscale drifts, sharp oscillations, and direction reversals characteristic of strong–field radiation–reacting motion. For each configuration we compute: (i) the relativistic kinetic energy
$K(\tau)=mc^{2}(\gamma(\tau)-1),$
(ii) the work done by the external fields,
$W_{\rm ext}(\tau)=\int_{\tau_{0}}^{\tau}q\,\mathbf{E}_{\rm ext}\!\cdot\!\mathbf{v}\,d\sigma,$
and (iii) the radiated energy,
$E_{\rm rad}(\tau)=\int_{\tau_{0}}^{\tau}P_{\rm cov}(\sigma)\,d\sigma,$
where $P_{\rm cov}$ is the covariant LL Larmor power. Energy balance requires
$\Delta K=W_{\rm ext}-E_{\rm rad},$
and we quantify this balance using the parameter
$R=\frac{|\Delta K-W_{\rm ext}|}{E_{\rm rad}}.$
Across the four simulations the LL energy balance agrees at a satisfactory numerical level:
$R_{\rm cyclotron}\simeq 1.03,\qquad R_{\rm Case\,1}\simeq 0.95,\qquad R_{\rm Case\,2}\simeq 1.00,\qquad R_{\rm Case\,3}\simeq 1.09.$
Even in the multidimensional, strongly nonlinear crossed–wave configuration (Case 3), the energy consistency remains within $\sim 8\%$, confirming that the Landau–Lifshitz order–reduced dynamics is energetically reliable across static, time–dependent, and fully multidimensional forcing environments.
Table 1: Energy–balance diagnostics for the cyclotron orbit and the three external–field configurations studied in Sec. III. Here $\Delta K=|K(t_{f})-K(0)|$, $W_{\rm ext}=\left|\int_{0}^{t_{f}}d\tau\,q\,\mathbf{E}\!\cdot\!\mathbf{v}\right|$, $E_{\rm rad}$ is the covariant radiated energy, and $R\equiv|\Delta K-W_{\rm ext}|/E_{\rm rad}$ tests the LL energy–balance relation.
Case $\Delta K$ $W_{\rm ext}$ $E_{\rm rad}$ $R$
Case 0: The Cyclotron case 1.55459 0 1.51272 1.02768
Case 1: Constant transverse $E$ and $B$ 0.00583564 0.005916 0.0000841766 0.954583
Case 2: Oscillating electric & magnetic fields 0.472655 0.296402 0.175510 1.00423
Case 3: Two crossed electromagnetic waves 39.5982 43.4668 3.55509 1.08818
V Post-Newtonian Truncation of the Lorentz–Dirac Equation
We now consider the more involved problem of a system of point charges interacting through their electromagnetic fields. In the near zone, where typical velocities satisfy $v\ll c$, the electromagnetic interaction can be organized in a post-Newtonian (PN) expansion in powers of $v/c$. The Newtonian (0PN) dynamics is governed by the instantaneous Coulomb force, while the first nontrivial corrections arise at 1PN order. These corrections are conservative: after renormalizing the divergent self-energy of each point charge into a physical mass, the energy is conserved up to 1PN order. Dissipative effects associated with electromagnetic radiation first appear at 1.5PN order. These arise from the radiative part of the electromagnetic field, which is expressed in terms of the electric dipole moment of the charge distribution,
$\mathbf{d}(t)=\sum_{a}q_{a}\,\mathbf{x}_{a}(t).$ (41)
Starting from our covariant Lorentz–Dirac equation (16) which we repeat here for clarity
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\left(\dot{a}^{\alpha}-\frac{a^{2}u^{\alpha}}{c^{2}}\right),$ (42)
we can are now in a position to systematically truncate our covariant expression above to the desired 1.5PN order. Specializing to electromagnetic external forces, the covariant momentum–balance equation for a charged particle reads
$\frac{dp^{\mu}_{a}}{d\tau}=\frac{q_{a}}{m_{a}}\,F^{\mu\nu}_{\mathrm{ext}}\!\big[x_{a}(\tau)\big]\,p_{a,\nu}+\frac{1}{4\pi\varepsilon_{0}}\frac{2q_{a}^{2}}{3m_{a}c^{3}}\frac{d^{2}p^{\mu}_{a}}{d\tau^{2}}-\frac{1}{4\pi\varepsilon_{0}}\frac{2q_{a}^{2}}{3m_{a}^{3}c^{5}}\!\left(\frac{dp_{a,\nu}}{d\tau}\frac{dp_{a}^{\nu}}{d\tau}\right)\!p^{\mu}_{a},$ (43)
where $a$ is the particle index. We observe that the self–interaction terms involve higher derivatives of the four–momentum. These can be split into two physically distinct contributions. The first term,
$F^{\mu}_{\mathrm{Schott}}=\frac{1}{4\pi\varepsilon_{0}}\frac{2q_{a}^{2}}{3m_{a}c^{3}}\,\frac{d^{2}p^{\mu}_{a}}{d\tau^{2}},$ (44)
is a total proper–time derivative of acceleration and is known as the Schott term. It describes a reversible exchange of energy–momentum between the particle and its near (bound) electromagnetic field.
The second term,
$F^{\mu}_{\mathrm{rad}}=-\frac{1}{4\pi\varepsilon_{0}}\frac{2q_{a}^{2}}{3m_{a}^{3}c^{5}}\!\left(\frac{dp_{a,\nu}}{d\tau}\frac{dp_{a}^{\nu}}{d\tau}\right)\!p^{\mu}_{a},$ (45)
represents the irreversible loss of energy–momentum carried away by electromagnetic radiation and is therefore called the radiation–reaction term. The total self–force is the sum of both contributions,
$F^{\mu}_{\mathrm{self}}=F^{\mu}_{\mathrm{Schott}}+F^{\mu}_{\mathrm{rad}}.$ (46)
To reach the non–relativistic limit, we expand to lowest order in $v/c$. Focusing on the spatial components, we have
$\mathbf{F}_{\mathrm{Schott}}=\frac{1}{4\pi\varepsilon_{0}}\frac{2q_{a}^{2}}{3c^{3}}\,\dot{\mathbf{a}}_{a}+\mathcal{O}\!\left(\frac{1}{c^{5}}\right),\qquad\mathbf{F}_{\mathrm{rad}}=\mathcal{O}\!\left(\frac{1}{c^{5}}\right).$ (47)
Therefore, to leading order (1.5 PN) only the Schott term contributes, and Eq. (43) reduces to the familiar Abraham–Lorentz equation,
$m_{a}\,\frac{d\mathbf{v}_{a}}{dt}=\frac{1}{4\pi\varepsilon_{0}}\,\frac{2q_{a}^{2}}{3c^{3}}\,\frac{d^{2}\mathbf{v}_{a}}{dt^{2}}+\mathbf{F}_{\mathrm{ext}},$ (48)
which correctly reproduces the non–relativistic limit of the radiation–reaction force. As we noted, the characteristic radiation–reaction timescale,
$t_{a}=\frac{1}{4\pi\varepsilon_{0}}\,\frac{2q_{a}^{2}}{3m_{a}c^{3}},$ (49)
is extremely small for elementary charges ($\tau_{e}\sim 6\times 10^{-24}\,\mathrm{s}$). Equation (48) can then be expressed compactly as
$m_{a}\left(\frac{d\mathbf{v}_{a}}{dt}-\tau_{a}\,\frac{d^{2}\mathbf{v}_{a}}{dt^{2}}\right)=\mathbf{F}_{\mathrm{ext}},$ (50)
which is the standard non–relativistic Abraham–Lorentz form. Here, the first term represents the usual acceleration due to external forces, while the second term encodes the leading–order radiation–reaction correction, arising from the self–interaction of the accelerated charge. A near-zone expansion of the Liénard–Wiechert potentials also shows that the 1.51.5PN contribution to the vector potential is spatially uniform,
$\mathbf{A}_{1.5\mathrm{PN}}(t)=-\,\frac{1}{4\pi\varepsilon_{0}}\,\frac{1}{c^{3}}\,\ddot{\mathbf{d}}(t),$ (51)
so that the corresponding magnetic field vanishes. The associated electric field is
$\mathbf{E}_{1.5\mathrm{PN}}(t)=\frac{1}{4\pi\varepsilon_{0}}\,\frac{2}{3c^{3}}\,\dddot{\mathbf{d}}(t).$ (52)
The radiation-reaction force acting on charge $q_{a}$ at position $\mathbf{x}_{a}$ is therefore
$\mathbf{F}_{a}^{(1.5\mathrm{PN})}=\frac{q_{a}}{4\pi\varepsilon_{0}}\,\frac{2}{3c^{3}}\,\dddot{\mathbf{d}}(t).$ (53)
This term breaks time-reversal symmetry and leads to energy dissipation. The mechanical energy of the system satisfies
$\frac{dE}{dt}=-\frac{1}{4\pi\varepsilon_{0}}\,\frac{2}{3c^{3}}\,\big|\ddot{\mathbf{d}}(t)\big|^{2},$ (54)
which is exactly the Larmor power radiated by an accelerated charge distribution. To leading order in $v/c$, the temporal components of Eqs. (44)–(45) yield
$F^{0}_{\mathrm{Schott}}$ $=\frac{1}{4\pi\varepsilon_{0}}\frac{2q_{a}^{2}}{3c^{4}}\big(\dot{\mathbf{v}}_{a}\!\cdot\!\mathbf{v}_{a}+v_{a}^{2}\big)+\mathcal{O}\!\left(\frac{1}{c^{6}}\right),$ (55)
$F^{0}_{\mathrm{rad}}$ $=-\frac{1}{4\pi\varepsilon_{0}}\frac{2q_{a}^{2}}{3c^{4}}\,v_{a}^{2}+\mathcal{O}\!\left(\frac{1}{c^{6}}\right),$ (56)
which govern the evolution of the particle’s energy. The sum of these two contributions represents the total temporal self–force, and only $F^{0}_{\mathrm{Schott}}$ contributes at leading order in $v/c$. The corresponding energy balance law reads
$\frac{dE_{a}}{dt}=\frac{1}{4\pi\epsilon_{0}}\,\frac{2}{3c^{3}}\left[\left(\dddot{\mathbf{d}}_{a}\cdot\dot{\mathbf{d}}_{a}+\ddot{\mathbf{d}}_{a}^{\,2}\right)-\ddot{\mathbf{d}}_{a}^{\,2}\right]+\mathbf{F}_{\rm ext}\!\cdot\!\mathbf{v}_{a}.$ (57)
where the first term inside the brackets arises from the Schott contribution and the second from the radiative loss term. The Schott term represents a total time derivative of the near–field energy, describing reversible energy exchange between the charge and its bound electromagnetic field. In contrast, $F^{0}_{\mathrm{rad}}$ corresponds to the irreversible radiation of energy through electromagnetic waves, reproducing the Larmor power at order $1/c^{3}$. In particular, the Schott term, being a total derivative, produces the post–Newtonian energy correction $E_{1.5\mathrm{PN}}$, which accounts for the instantaneous exchange of energy between the particle and its self–field
VI The N-body framework: Darwin Hamiltonian, Hamilton’s equations and 1.5PN radiation reaction
(a) Trajectories.
(b) Darwin Hamiltonian evolution.
Figure 3: Two oppositely charged blobs, each containing ten particles, evolved with the conservative Darwin Hamiltonian at 1PN order. Panel (a) shows the three–dimensional trajectories (blue: positive charges; black: negative charges), and panel (b) the corresponding evolution of the Darwin Hamiltonian $H(\tau)$.
In gravitational two–body dynamics it is standard, both in the post-Newtonian (PN) formalism and in the Effective-One-Body (EOB) approach, to decompose the motion into a conservative Hamiltonian sector and a separate dissipative sector describing radiation reaction [27, 30, 29] . The conservative dynamics admit a fully canonical formulation in the ADM Hamiltonian approach up to at least 3PN and 4PN order [20, 21, 22], while the leading dissipative terms enter at 2.5PN and 3.5PN order [23, 24, 25, 26]. These radiation-reaction contributions cannot be generated from a Hamiltonian, and are therefore incorporated as a nonconservative force $F^{\mathrm{RR}}_{i}$ added to the canonical momentum equation. The function $F^{\mathrm{RR}}_{i}$ is fixed by the balance equations equating the mechanical losses to the gravitational-wave energy and angular-momentum fluxes [26]. The same methodology is implemented in the EOB formalism [27, 28, 29]: a resummed Hamiltonian $H_{\mathrm{EOB}}(q,p)$ generates the conservative sector, while an EOB radiation-reaction force $\mathcal{F}_{i}^{\mathrm{RR}}$ is appended to the momentum equation and calibrated to PN fluxes and numerical-relativity waveforms [30, 31, 32]. This canonical plus non-Hamiltonian structure is further supported by effective field theory (EFT) treatments, where the PN radiation-reaction terms arise from a doubled-field nonconservative action principle [33, 34]. Motivated by this well-developed approach, we adopt the same strategy for electromagnetic systems: we use the well known Darwin Hamiltonian to generate the conservative 1PN dynamics of charged particles and supplement the canonical momentum equation with the leading 1.5PN dipole radiation-reaction force obtained by order reduction.
We will now implement the Post Newtonian truncated version of the radiation reaction and self force developed in the last section into the basic conservative 1PN Hamiltonian dynamics of a system of charges developed by Charles Galton Darwin. We will work in SI units. Particle labels are $a,b,c\in\{1,\dots,N\}$, and three–vectors are denoted in bold, e.g. $\mathbf{x}_{a},\mathbf{p}_{a},\mathbf{v}_{a}$. The Coulomb constant is
$k=\frac{1}{4\pi\varepsilon_{0}}.$
For each unordered pair $(a,b)$ we define the relative quantities
$\mathbf{r}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b},\qquad r_{ab}=|\mathbf{r}_{ab}|,\qquad\hat{\mathbf{r}}_{ab}=\frac{\mathbf{r}_{ab}}{r_{ab}}.$
We expand the conservative dynamics consistently to 1PN order, $\mathcal{O}(1/c^{2})$, and include the leading dissipative 1.5PN radiation–reaction force, $\mathcal{O}(1/c^{3})$, as a non-Hamiltonian perturbation via order reduction. Although the radiation–reaction force enters at 1.5PN order, its consistent evaluation requires only the conservative equations of motion through 1PN; higher-order conservative corrections would contribute solely beyond the accuracy considered here. The conservative dynamics up to $\mathcal{O}(1/c^{2})$ is encoded in the Darwin Hamiltonian
$H_{\rm D}$ $=\sum_{a}\left(\frac{\mathbf{p}_{a}^{2}}{2m_{a}}-\frac{\mathbf{p}_{a}^{4}}{8m_{a}^{3}c^{2}}\right)+\frac{1}{2}\sum_{a\neq b}k\frac{q_{a}q_{b}}{r_{ab}}$
$\qquad-\frac{1}{2}\sum_{a\neq b}\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}}\frac{1}{m_{a}m_{b}c^{2}}\frac{1}{r_{ab}}\left[\mathbf{p}_{a}\!\cdot\!\mathbf{p}_{b}+\frac{(\mathbf{p}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}}\right].$ (58)
The factor $\tfrac{1}{2}$ in the pair sums avoids double counting of unordered pairs. The kinetic term arises from the expansion $\sqrt{m^{2}c^{4}+c^{2}\mathbf{p}^{2}}\approx mc^{2}+\mathbf{p}^{2}/(2m)-\mathbf{p}^{4}/(8m^{3}c^{2})+\dots$, with the rest energy $mc^{2}$ omitted. The velocity–dependent interaction is the standard Darwin term obtained by expanding the Liénard–Wiechert potentials and performing a canonical transformation to $(\mathbf{x},\mathbf{p})$. Hamilton’s equations are
$\dot{\mathbf{x}}_{a}=\frac{\partial H_{\rm D}}{\partial\mathbf{p}_{a}},\qquad\dot{\mathbf{p}}_{a}=-\frac{\partial H_{\rm D}}{\partial\mathbf{x}_{a}}.$
We evaluate the derivatives term by term. For
$H_{\rm kin}=\sum_{a}\left(\frac{\mathbf{p}_{a}^{2}}{2m_{a}}-\frac{\mathbf{p}_{a}^{4}}{8m_{a}^{3}c^{2}}\right),$
we obtain
$\frac{\partial H_{\rm kin}}{\partial\mathbf{p}_{a}}$ $=\frac{\mathbf{p}_{a}}{m_{a}}-\frac{\mathbf{p}_{a}^{2}\,\mathbf{p}_{a}}{2m_{a}^{3}c^{2}},$ (59)
$-\frac{\partial H_{\rm kin}}{\partial\mathbf{x}_{a}}$ $=\mathbf{0}.$ (60)
The Coulomb piece is
$H_{\rm C}=\frac{1}{2}\sum_{a\neq b}k\frac{q_{a}q_{b}}{r_{ab}}.$
Differentiating with respect to $\mathbf{x}_{a}$ and using $\partial(1/r_{ab})/\partial\mathbf{x}_{a}=-\mathbf{r}_{ab}/r_{ab}^{3}=-\hat{\mathbf{r}}_{ab}/r_{ab}^{2}$, we find
$\frac{\partial H_{\rm C}}{\partial\mathbf{x}_{a}}=-\sum_{b\neq a}k\frac{q_{a}q_{b}}{r_{ab}^{2}}\,\hat{\mathbf{r}}_{ab},$
so that
$\dot{\mathbf{p}}_{a}\big|_{\rm Coulomb}=-\frac{\partial H_{\rm C}}{\partial\mathbf{x}_{a}}=\sum_{b\neq a}k\frac{q_{a}q_{b}}{r_{ab}^{2}}\,\hat{\mathbf{r}}_{ab}.$ (61)
This is the usual Coulomb force $\mathbf{F}_{ab}=q_{a}\mathbf{E}_{b}$ with $\mathbf{E}_{b}=kq_{b}\hat{\mathbf{r}}_{ab}/r_{ab}^{2}$. Focusing on the darwin Interaction at 1PN, let us introduce
$\mathcal{P}_{ab}:=\mathbf{p}_{a}\!\cdot\!\mathbf{p}_{b}+\frac{(\mathbf{p}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}},$
and thus we write the interaction as
$H_{\rm D,int}=-\frac{1}{2}\sum_{a\neq b}\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}}\frac{1}{m_{a}m_{b}c^{2}}\frac{1}{r_{ab}}\,\mathcal{P}_{ab}.$
We have
$\frac{\partial\mathcal{P}_{ab}}{\partial\mathbf{p}_{a}}=\mathbf{p}_{b}+\frac{\mathbf{r}_{ab}}{r_{ab}^{2}}(\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab}),$
and therefore
$\frac{\partial H_{\rm D,int}}{\partial\mathbf{p}_{a}}=-\sum_{b\neq a}\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}}\frac{1}{m_{a}m_{b}c^{2}}\frac{1}{r_{ab}}\left[\mathbf{p}_{b}+\mathbf{r}_{ab}\frac{\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab}}{r_{ab}^{2}}\right].$
Adding the kinetic piece gives the canonical velocity equation  [^1] For completeness, let us note that inverting perturbatively, we can write $\mathbf{p}_{a}=m_{a}\mathbf{v}_{a}+\delta\mathbf{p}_{a}$ and solve to $\mathcal{O}(1/c^{2})$. Replacing $\mathbf{p}_{b}\to m_{b}\mathbf{v}_{b}$ on the right–hand side of (64) of main text yields $\mathbf{p}_{a}=m_{a}\mathbf{v}_{a}+\frac{m_{a}v_{a}^{2}}{2c^{2}}\,\mathbf{v}_{a}+\sum_{b\neq a}\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}c^{2}r_{ab}}\left[\mathbf{v}_{b}+\mathbf{r}_{ab}\frac{\mathbf{v}_{b}\!\cdot\!\mathbf{r}_{ab}}{r_{ab}^{2}}\right]+O\!\left(\frac{1}{c^{4}}\right).$
$\dot{\mathbf{x}}_{a}=\frac{\mathbf{p}_{a}}{m_{a}}-\frac{\mathbf{p}_{a}^{2}\,\mathbf{p}_{a}}{2m_{a}^{3}c^{2}}-\sum_{b\neq a}\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}}\frac{1}{m_{a}m_{b}c^{2}}\frac{1}{r_{ab}}\left[\mathbf{p}_{b}+\mathbf{r}_{ab}\frac{\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab}}{r_{ab}^{2}}\right].$ (62)
Differentiating $H_{\rm D,int}$ with respect to $\mathbf{x}_{a}$ is straightforward but algebraically lengthy, since derivatives act on both $1/r_{ab}$ and the projection factor $(\mathbf{p}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab})/r_{ab}^{2}$. After simplification, and using $\mathbf{v}_{a}\approx\mathbf{p}_{a}/m_{a}$ inside 1PN terms, the Darwin correction to $\dot{\mathbf{p}}_{a}$ takes the standard velocity–dependent form (please see Appendix Section C for details),
$\dot{\mathbf{p}}_{a}\Big|_{\rm 1PN}=k\sum_{b\neq a}\frac{q_{a}q_{b}}{2c^{2}r_{ab}^{2}}\Big[\,\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})-\hat{\mathbf{r}}_{ab}\big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\big)\Big].$ (63)
Combining (61) and (63) gives the full conservative momentum equation,
$\dot{\mathbf{p}}_{a}\big|_{\rm cons}=\sum_{b\neq a}k\frac{q_{a}q_{b}}{r_{ab}^{2}}\,\hat{\mathbf{r}}_{ab}+k\sum_{b\neq a}\frac{q_{a}q_{b}}{2c^{2}r_{ab}^{2}}\Big[\,\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})-\hat{\mathbf{r}}_{ab}\big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\big)\Big],$
where one may use $\mathbf{v}\!\to\!\mathbf{p}/m$ inside it to the desired PN order. Equation (62) yields $\mathbf{v}_{a}=\dot{\mathbf{x}}_{a}$ in terms of $\mathbf{p}_{a}$. To 1PN,
$\mathbf{v}_{a}$ $=\frac{\mathbf{p}_{a}}{m_{a}}-\frac{\mathbf{p}_{a}^{2}\,\mathbf{p}_{a}}{2m_{a}^{3}c^{2}}-\sum_{b\neq a}\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}}\frac{1}{m_{a}m_{b}c^{2}}\frac{1}{r_{ab}}\left[\mathbf{p}_{b}+\mathbf{r}_{ab}\frac{\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab}}{r_{ab}^{2}}\right]+O\!\left(\frac{1}{c^{4}}\right).$ (64)
The leading dissipative contribution arises from dipole radiation reaction. The near–zone RR force can be written as
$\mathbf{F}_{a}^{(1.5\mathrm{PN})}(t)=k\,\frac{2q_{a}}{3c^{3}}\,\dddot{\mathbf{d}}(t),\qquad\mathbf{d}(t)=\sum_{b}q_{b}\mathbf{x}_{b}(t),$ (65)
and is treated by order reduction, expressing $\dddot{\mathbf{d}}$ in terms of positions and velocities via the Newtonian dynamics. The Newtonian (Coulomb) acceleration of particle $b$ is
$\mathbf{a}_{b}^{(N)}=\frac{1}{m_{b}}\mathbf{F}_{b}^{(N)}=\frac{1}{4\pi\varepsilon_{0}m_{b}}\sum_{c\neq b}q_{b}q_{c}\frac{\mathbf{r}_{bc}}{r_{bc}^{3}}.$ (66)
(Here $\mathbf{F}_{bc}=q_{b}\mathbf{E}_{c}$ with $\mathbf{E}_{c}=kq_{c}\mathbf{r}_{bc}/r_{bc}^{3}$.) Differentiating in time, with $\mathbf{v}_{bc}=\mathbf{v}_{b}-\mathbf{v}_{c}$ and
$\frac{d}{dt}\left(\frac{\mathbf{r}}{r^{3}}\right)=\frac{1}{r^{3}}\Big[\mathbf{v}-3\hat{\mathbf{r}}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v})\Big],$
we obtain
$\dot{\mathbf{a}}_{b}^{(N)}$ $=\frac{1}{4\pi\varepsilon_{0}m_{b}}\sum_{c\neq b}q_{b}q_{c}\frac{1}{r_{bc}^{3}}\Big[\mathbf{v}_{bc}-3\hat{\mathbf{r}}_{bc}(\hat{\mathbf{r}}_{bc}\!\cdot\!\mathbf{v}_{bc})\Big].$ (67)
Multiplying (67) by $q_{b}$ and summing over $b$ gives
$\dddot{\mathbf{d}}(t)$ $=\sum_{b}q_{b}\dot{\mathbf{a}}_{b}^{(N)}$
$=\frac{1}{4\pi\varepsilon_{0}}\sum_{b}\sum_{c\neq b}\frac{q_{b}^{2}q_{c}}{m_{b}r_{bc}^{3}}\Big[\mathbf{v}_{b}-\mathbf{v}_{c}-3\hat{\mathbf{r}}_{bc}\big(\hat{\mathbf{r}}_{bc}\!\cdot\!(\mathbf{v}_{b}-\mathbf{v}_{c})\big)\Big].$ (68)
Defining
$\mathbf{S}:=\sum_{b}\sum_{c\neq b}\frac{q_{b}^{2}q_{c}}{m_{b}r_{bc}^{3}}\Big[\mathbf{v}_{b}-\mathbf{v}_{c}-3\hat{\mathbf{r}}_{bc}\big(\hat{\mathbf{r}}_{bc}\!\cdot\!(\mathbf{v}_{b}-\mathbf{v}_{c})\big)\Big],$
Eq. (68) simply reads
$\dddot{\mathbf{d}}(t)=\frac{1}{4\pi\varepsilon_{0}}\,\mathbf{S}.$
Substituting (68) into (65), and using $k=1/(4\pi\varepsilon_{0})$, we obtain
$\mathbf{F}_{a}^{(1.5\mathrm{PN})}(t)$ $=k\,\frac{2q_{a}}{3c^{3}}\,\dddot{\mathbf{d}}(t)=\frac{1}{4\pi\varepsilon_{0}}\frac{2q_{a}}{3c^{3}}\cdot\frac{1}{4\pi\varepsilon_{0}}\,\mathbf{S}$
$=\frac{1}{(4\pi\varepsilon_{0})^{2}}\frac{2q_{a}}{3c^{3}}\,\sum_{b}\sum_{c\neq b}\frac{q_{b}^{2}q_{c}}{m_{b}r_{bc}^{3}}\Big[\mathbf{v}_{b}-\mathbf{v}_{c}-3\hat{\mathbf{r}}_{bc}\big(\hat{\mathbf{r}}_{bc}\!\cdot\!(\mathbf{v}_{b}-\mathbf{v}_{c})\big)\Big].$ (69)
At canonical level we replace $\mathbf{v}_{b}\mapsto\mathbf{p}_{b}/m_{b}$ inside the bracket (consistent with 1.5PN accuracy), leading to the final implementable form
$\mathbf{F}_{a}^{(1.5\mathrm{PN})}(t)=\frac{1}{(4\pi\varepsilon_{0})^{2}}\frac{2q_{a}}{3c^{3}}\sum_{b}\sum_{c\neq b}\frac{q_{b}^{2}q_{c}}{m_{b}r_{bc}^{3}}\Big[\tfrac{\mathbf{p}_{b}}{m_{b}}-\tfrac{\mathbf{p}_{c}}{m_{c}}-3\hat{\mathbf{r}}_{bc}\big(\hat{\mathbf{r}}_{bc}\!\cdot\!(\tfrac{\mathbf{p}_{b}}{m_{b}}-\tfrac{\mathbf{p}_{c}}{m_{c}})\big)\Big].$ (70)
The mechanical power supplied by RR is then
$P_{\rm mech}^{\rm(RR)}$ $=\sum_{a}\mathbf{F}_{a}^{(1.5\rm PN)}\!\cdot\!\mathbf{v}_{a}=\frac{2}{3(4\pi\varepsilon_{0})c^{3}}\,\dddot{\mathbf{d}}\!\cdot\!\dot{\mathbf{d}}$
$=\frac{d}{dt}\!\left(\frac{\dot{\mathbf{d}}\!\cdot\!\ddot{\mathbf{d}}}{6\pi\varepsilon_{0}c^{3}}\right)-\frac{\ddot{\mathbf{d}}\!\cdot\!\ddot{\mathbf{d}}}{6\pi\varepsilon_{0}c^{3}}=\frac{d}{dt}(\text{Schott})-P_{\rm rad},$ (71)
where $P_{\rm rad}=\dfrac{1}{6\pi\varepsilon_{0}c^{3}}\,\ddot{\mathbf{d}}^{2}$ is the dipole Larmor power and the total derivative represents the Schott term. Thus (70) is fully consistent with electromagnetic energy balance. Collecting the results, the 1.5PN equations of motion in canonical variables are
$\dot{\mathbf{x}}_{a}$ $=\frac{\partial H_{\rm D}}{\partial\mathbf{p}_{a}}=\frac{\mathbf{p}_{a}}{m_{a}}-\frac{\mathbf{p}_{a}^{2}\,\mathbf{p}_{a}}{2m_{a}^{3}c^{2}}-\sum_{b\neq a}\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}}\frac{1}{m_{a}m_{b}c^{2}}\frac{1}{r_{ab}}\left[\mathbf{p}_{b}+\mathbf{r}_{ab}\frac{\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab}}{r_{ab}^{2}}\right],$
$\dot{\mathbf{p}}_{a}$ $=-\frac{\partial H_{\rm D}}{\partial\mathbf{x}_{a}}+\mathbf{F}_{a}^{(1.5\mathrm{PN})}(\mathbf{x},\mathbf{p}),$ (72)
with $\partial H_{\rm D}/\partial\mathbf{x}_{a}$ given explicitly by (61) and (63), and the 1.5PN radiation–reaction force $\mathbf{F}_{a}^{(1.5\mathrm{PN})}$ given by (70). As a stringent validation of the conservative sector of our framework, we consider a system of two spatially separated three–dimensional charge blobs of opposite sign, evolved using the Darwin Hamiltonian alone, with all radiation–reaction terms explicitly switched off. Each blob consists of multiple point charges initialized within a compact spherical region, with the two charge clouds separated by a distance large compared to their individual sizes.
Figure 3(a) displays the resulting trajectories. Positively charged particles are shown in blue and negatively charged particles in black, with filled markers indicating the initial positions. The evolution is fully conservative, with the velocity–dependent Darwin interaction producing nontrivial many–body dynamics while preserving the overall phase–space structure.
The corresponding Hamiltonian $H(\tau)$ is shown in Fig. 3(b). As expected for purely conservative dynamics, no secular drift is observed. The small bounded fluctuations visible in $H(\tau)$ arise from unavoidable numerical stiffness during close encounters between particles and provide a diagnostic of numerical accuracy. The absence of any systematic energy loss confirms that the Darwin sector has been implemented consistently and provides a reliable starting point for the inclusion of dissipative radiation–reaction effects in the next section.
VII Simulation results: the Charge Neutral Binary
We now present representative solutions of the dissipative 1.5PN accurate NN–body framework Eq. (72). Throughout this section we use $\tau$ to denote the ordinary coordinate time (not the proper time). The conservative interactions are evolved with the Darwin Hamiltonian, retaining all terms through 1PN order, $\mathcal{O}(1/c^{2})$, while dissipation is incorporated as a non-Hamiltonian Landau–Lifshitz radiation–reaction force at leading 1.5PN order, $\mathcal{O}(1/c^{3})$, implemented via order reduction. This canonical-plus-dissipative framework allows us to test how radiative losses modify otherwise integrable Coulombic motion, producing secular inspiral and circularization. The examples below progress from a symmetric equal-mass neutral binary to an extreme-mass-ratio “hydrogen-like” configuration and finally to an initially eccentric orbit that circularizes under radiation reaction; in each case we monitor trajectories and the Hamiltonian $H(\tau)$ to quantify the cumulative energy loss.
(a) 3D trajectories with initial (green) and final (red) markers.
(b) Parametric projection in the xxyy plane.
(c) $x(\tau)$ vs. $\tau$.
(d) $y(\tau)$ vs. $\tau$.
(e) $z(\tau)$ vs. $\tau$.
(f) Hamiltonian $H(\tau)$ showing radiative energy loss.
Figure 4: Charge neutral binary of same mass starting from diametrically opposite locations with equal and opposite momenta, with radiation reaction from our truncated post Newtonian approach: (a) Full 3D trajectories; (b) parametric xxyy plane; (c–e) components $x(\tau)$, $y(\tau)$, $z(\tau)$; (f) Hamiltonian $H(\tau)$, decreasing due to radiation. The initial particle positions are shown in green, and the final positions at $\tau=\tau_{f}$ are shown in red.
(a) 3D trajectories with initial (green) and final (red) markers.
(b) Parametric projection in the xxyy plane.
(c) $x(\tau)$ vs. $\tau$.
(d) $y(\tau)$ vs. $\tau$.
(e) $z(\tau)$ vs. $\tau$.
(f) Hamiltonian $H(\tau)$ showing radiative energy loss.
Figure 5: Charge neutral binary of extreme mass ratio, one heavy and one light (with circular orbit initial conditions), with radiation reaction from our truncated post Newtonian approach: (a) Full 3D trajectories; (b) parametric xxyy plane; (c–e) components $x(\tau)$, $y(\tau)$, $z(\tau)$; (f) Hamiltonian $H(\tau)$, decreasing due to radiation. The initial particle positions are shown in green, and the final positions at $\tau=\tau_{f}$ are shown in red.
(a) 3D trajectories with initial (green) and final (red) markers.
(b) Parametric projection in the xxyy plane.
(c) $x(\tau)$ vs. $\tau$.
(d) $y(\tau)$ vs. $\tau$.
(e) $z(\tau)$ vs. $\tau$.
(f) Hamiltonian $H(\tau)$ showing radiative energy loss.
Figure 6: Charge neutral binary, one heavy and one light, with elliptic orbit initial conditions. Radiation reaction from our truncated post Newtonian approach shows clear signature of circularization: (a) Full 3D trajectories; (b) parametric xxyy plane; (c–e) components $x(\tau)$, $y(\tau)$, $z(\tau)$; (f) Hamiltonian $H(\tau)$, decreasing due to radiation. The initial particle positions are shown in green, and the final positions at $\tau=\tau_{f}$ are shown in red.
VII.1 Charge neutral binary of same mass
To validate our construction of the conservative plus dissipative system governed by Eq. (72), we first consider a symmetric charge-neutral bound configuration consisting of two point charges of equal mass $m=1$ and opposite charge $\pm q$, interacting via an attractive Coulomb potential V(r)=k/rV(r)=-k/r. The particles are initially placed at diametrically opposite positions,
$\mathbf{r}_{1}(0)=(1,0,0),\qquad\mathbf{r}_{2}(0)=(-1,0,0),$
so that the interparticle separation is $r=2R=2$. For a circular orbit the centripetal balance condition,
$\frac{mv^{2}}{R}=\frac{\alpha}{(2R)^{2}},$
requires a tangential speed $v=\tfrac{1}{2}$ for $m=k=R=1$. We therefore choose the initial momenta
$\mathbf{p}_{1}(0)=(0,\,0.5,\,0),\qquad\mathbf{p}_{2}(0)=(0,\,-0.5,\,0),$
which supply the equal and opposite tangential velocities needed to sustain the circular motion. The corresponding total energy,
$E_{\rm tot}=mv^{2}-\frac{k}{2R}=-\,\frac{1}{4},$
is negative as expected for a bound nonrelativistic Coulomb orbit. These initial data are chosen so that the nonrelativistic Coulomb dynamics would support a circular orbit of radius unity. When the Landau–Lifshitz self–force is included, however, the system no longer admits an exact periodic solution: radiation reaction steadily removes mechanical energy, producing a gradual reduction in orbital radius and an associated inward spiral, as shown in Fig. 4. To make the dissipative dynamics visible on numerically tractable time scales, we deliberately choose $c$ such that the leading radiation– reaction force, scaling as $\mathcal{O}(1/c^{3})$, is suppressed by only $\sim 10^{-2}$ relative to the conservative 1PN corrections $\mathcal{O}(1/c^{2})$. This choice does not affect the formal consistency of the PN expansion, but simply rescales the inspiral time.
The six–panel figure displays the full three-dimensional worldlines, the parametric projection in the xxyy plane, and the individual coordinate functions $x(\tau)$, $y(\tau)$, and $z(\tau)$. The final panel shows the monotonic decrease of the Hamiltonian $H(\tau)$, which directly reflects the radiative energy loss predicted by our order-reduced evolution. For clarity, the initial particle positions are rendered in green, while their final locations at $\tau=\tau_{f}$ are shown in red, making the cumulative orbit shrinkage visually explicit. This two-body example provides a clean test of the LL radiation–reaction scheme and illustrates how even a perfectly symmetric bound configuration evolves irreversibly once electromagnetic self-interaction is taken into account.
VII.2 Charge neutral binary of extreme mass ratio
As a second application of our N–body framework Eq. (72), we consider an unequal–mass, oppositely charged system designed to mimic a classical “hydrogen–like” configuration. For a pair of point charges $(q_{1},m_{1})$ and $(q_{2},m_{2})$ interacting through the Coulomb potential, a circular orbit may be constructed by imposing the usual centripetal balance in the relative two–body system. Denoting by $\mathbf{r}=\mathbf{x}_{2}-\mathbf{x}_{1}$ the separation vector, the magnitude of the Coulomb force is $F_{C}=k\,|q_{1}q_{2}|/r^{2}$ with $k=1/(4\pi\varepsilon_{0})$. In the centre–of–mass frame the relative coordinate evolves as a particle of reduced mass $\mu=m_{1}m_{2}/(m_{1}+m_{2})$ in this central potential. A circular orbit of radius $r_{0}$ requires
$\frac{\mu v_{\rm rel}^{2}}{r_{0}}=k\,\frac{|q_{1}q_{2}|}{r_{0}^{2}},\qquad\Rightarrow\qquad v_{\rm rel}=\sqrt{\,k\,\frac{|q_{1}q_{2}|}{\mu r_{0}}\,}$, (73)
where $v_{\rm rel}$ is the relative orbital speed. One convenient choice of initial data is to place the two charges on the xx–axis at
$\mathbf{x}_{1}(0)=-\frac{m_{2}}{m_{1}+m_{2}}\,(r_{0},0,0),\qquad\mathbf{x}_{2}(0)=\frac{m_{1}}{m_{1}+m_{2}}\,(r_{0},0,0),$ (74)
with momenta
$\mathbf{p}_{1}(0)=-\mathbf{p}_{2}(0),\qquad|\mathbf{p}_{2}(0)|=\mu~v_{\rm rel},$ (75)
and the velocity of particle 2 chosen along the +y+y direction to produce a counterclockwise orbit in the $xy$ plane. In the heavy–nucleus limit $m_{1}\gg m_{2}$ these expressions reduce to the familiar hydrogenic initial data $\mathbf{x}_{1}(0)=(0,0,0)$, $\mathbf{x}_{2}(0)=(r_{0},0,0)$, and $\mathbf{p}_{2}(0)=(0,\,\sqrt{m_{2}k|q_{1}q_{2}|/r_{0}},\,0)$. We choose charges $q_{1}=+1$, $q_{2}=-1$ and masses $m_{1}\gg m_{2}$ so that particle 1 acts as a heavy Coulomb center, while particle 2 executes a bound orbit. This initialization guarantees that, in the absence of radiation reaction, the system would remain on a circular Coulomb orbit. When LL radiation reaction is included, however, the light particle gradually spirals inward as it continuously loses energy to electromagnetic radiation. Once again, for numerical illustration, we choose the speed of light $c$ to be moderately small so that the radiation–reaction force, which enters at $\mathcal{O}(1/c^{3})$, is only suppressed by a factor $\sim 10^{-3}$ relative to the conservative 1PN terms; this exaggerates dissipative effects and allows the inspiral to be resolved on accessible time scales without altering the underlying PN ordering.
Figure 5 shows the resulting inspiral. Panel (a) displays the full 3D trajectories, with green and red markers denoting the initial and final positions. The parametric projection in the xxyy plane (b) makes the secular inward drift clearly visible. The coordinate components $x(\tau)$, $y(\tau)$, and $z(\tau)$ plotted in panels (c)–(e) exhibit both the fast orbital oscillations and the slow amplitude decay associated with radiative backreaction. Finally, the Hamiltonian $H(\tau)$ in panel (f) shows a clean and monotonic decrease, confirming that the truncated post–Newtonian LL scheme captures the expected secular energy loss. This hydrogen–like example illustrates how the LL self–force naturally produces inspiral dynamics in a bound Coulomb system, providing a classical analogue of radiation–reaction–driven orbital decay.
VII.3 Charge neutral binary of comparable mass showing circularization of initial elliptic orbits
In Fig. 6 we examine a two–body configuration demosntrating the radiative circularization predicted by our truncated post-Newtonian scheme Eq. (72). We consider oppositely charged particles with masses $(m_{1},m_{2})=(5,1)$ initialized in the center–of–mass frame at a separation $r_{0}=1$,
$\mathbf{r}_{1}(0)=(0,0,0),\qquad\mathbf{r}_{2}(0)=\frac{m_{1}}{m_{1}+m_{2}}(r_{0},0,0),$
and assign equal–magnitude and opposite momenta $\mathbf{p}_{1}(0)=-\mathbf{p}_{2}(0)$ with $|\mathbf{p}_{1}(0)|=\mu\,v_{\rm rel}$,
$v_{\rm rel}=\sqrt{\frac{k\,|q_{1}q_{2}|}{\mu\,r_{0}}},\qquad\mu=\frac{m_{1}m_{2}}{m_{1}+m_{2}},$
so that, in the absence of radiation reaction, the orbit is elliptic. When evolved in our scheme governed by Eq. (72), the orbit gradually shrinks and circularizes, as seen in the parametric xxyy projection in Fig. 6(b). Panels (c)–(e) display the individual coordinate components $x(\tau)$, $y(\tau)$, and $z(\tau)$, showing the expected modulation of the radial amplitude and no out–of–plane drift. The Hamiltonian $H(\tau)$ in Fig. 6(f) decreases monotonically, matching the cumulative Larmor power by construction. As with earlier examples, the initial particle positions are marked in green and their final locations at $\tau=\tau_{f}$ are indicated in red.
VII.4 Comments on the Painlevé universal structures in binary black hole dynamics
The circularization observed in Fig. 6 highlights a robust feature of dissipative Coulomb dynamics. Although the charge–neutral binary studied here is neither integrable nor restricted to a single effective degree of freedom, radiation reaction systematically drives initially eccentric orbits toward quasi–circular inspiral while producing a monotonic decrease of the Hamiltonian. This behavior closely parallels the circularization of gravitational binaries under radiation reaction and suggests a form of universality in dissipative long–range interactions. In simpler analytic settings, it has been shown that the Landau–Lifshitz equations for a radiating charge in a Coulomb potential admit an exact reduction to the Painlevé–II equation, providing a rare example where radiation–reaction–driven inspiral can be described in closed form [35]. Related order–reduced formulations have also been explored for radiating spinning particles in external magnetic fields [36]. Motivated in part by such results, Painlevé transcendents have been proposed as underlying universal structures of inspiral and merger dynamics in binary black hole coalescence [37, 38, 39]. The framework developed here complements these analytic and asymptotic approaches in a relativistic electromagnetic system with radiation reaction: rather than relying on integrability in limiting situations.
VIII Conclusion and Future Directions
In this work we have examined the classical theory of radiation reaction motivated by post–Newtonian dynamics of binary systems in general relativity. Starting from the Lorentz–Dirac equation, we briefly reviewed its derivations from both the bound/radiative field split and Dirac’s world–tube momentum–balance construction, and traced the appearance of runaway and preaccelerated solutions to the third–order structure of the equation [7, 10, 9, 8, 15, 14, 16]. Implementing the Landau–Lifshitz reduction of order, we obtained a second–order, causal evolution equation and tested it in a set of single–particle simulations in static, time–dependent, and fully multidimensional electromagnetic fields. For each configuration we verified energy balance by comparing the change in relativistic kinetic energy, the work done by external fields, and the covariant Larmor power, finding agreement at the few–percent level across all cases.
Building on this single–particle results, we developed a PN Hamiltonian framework for interacting and radiating charges by combining the conservative Darwin Hamiltonian at 1PN order with the leading 1.5PN dipole radiation–reaction force obtained via order reduction. In this canonical-plus-dissipative scheme, the conservative sector is generated by a standard Hamiltonian while radiation reaction enters only through a nonconservative force in the momentum equation, in direct analogy with gravitational PN and effective–one–body (EOB) treatments of compact binaries [20, 21, 22, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Numerical experiments with equal–mass and hydrogen–like two–body systems show orbit circularization, inward inspiral, and monotonic decrease of the Hamiltonian, with the dissipative power accurately tracking the dipole Larmor flux. These examples provide fully reproducible examples in which dissipative electromagnetic dynamics can be explored at PN accuracy.
The formalism developed here is directly relevant to a broad class of high–energy astrophysical environments in which radiation losses regulate charged–particle dynamics. Examples include pulsar and magnetar magnetospheres, black–hole magnetospheres, and gamma–ray burst outflows, where curvature and synchrotron losses shape both particle distributions and observed spectra [1, 2, 3, 4].
Several immediate extensions suggest themselves. On the electromagnetic side, it will be important to move beyond the two–body examples considered here and analyze many–particle systems in which long–range Coulomb forces and radiation reaction compete to produce collective phenomena such as clustering, evaporation, or dissipation–induced chaos. The formalism can also be generalized to spatially structured and turbulent background fields, more directly mimicking magnetospheres, jets, and laser–driven plasmas. On the gravitational side, our results motivate a systematic use of electromagnetic analog models to probe issues of radiation–reaction–driven inspiral, waveform sensitivity, and nonconservative phase–space structure in a framework that is simpler than full general relativity yet shares the same Hamiltonian plus radiation–reaction framework as PN, self–force, and EOB descriptions of compact binaries [6, 26, 29, 30, 31].
Overall, the framework presented here provides a simple system in which the dynamics of radiating and interacting point charges can be studied with the same conceptual tools used in modern gravitational–wave theory. We expect that further development of these electromagnetic analogs will aid the interpretation of high–energy astrophysical observations, inform the construction of nonconservative actions for self–interacting systems, and help bridge the gap between classical radiation reaction, semiclassical corrections, and fully relativistic models of compact objects and strong–field plasmas.
IX Acknowledgments
RS thanks Suryateja Gavva, Anuradha Gupta, Sashwat Tanay and Leo C Stein for discussions on related problems in binary black hole dynamics. R.S is supported by DST INSPIRE Faculty fellowship, India (Grant No.IFA19-PH231). All authors acknowledge support from NFSG and OPERA Research Grant from Birla Institute of Technology and Science, Pilani (Hyderabad Campus).
X Data Availability
The data that supports the findings of this study are available within the article.
Appendix A Runaway solution
Starting from the reduced one-dimensional Lorentz–Dirac equation,
$a(t)-t_{0}\,\dot{a}(t)=\frac{1}{m}F_{\mathrm{ext}}(t),\qquad t_{0}=\frac{2}{3}\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}m},$ (76)
we treat it as a first–order linear differential equation for the acceleration $a(t)$. Multiplying by the integrating factor $e^{-t/t_{0}}$ gives
$\frac{d}{dt}\!\left(e^{-t/t_{0}}a(t)\right)=-\frac{e^{-t/t_{0}}}{mt_{0}}\,F_{\mathrm{ext}}(t),$
which upon integration yields
$a(t)=e^{t/t_{0}}\!\left[a(0)-\frac{1}{mt_{0}}\!\int_{0}^{t}e^{-t^{\prime}/t_{0}}F_{\mathrm{ext}}(t^{\prime})\,dt^{\prime}\right].$
For a force that is switched on abruptly and then held constant, $F_{\mathrm{ext}}(t)=f\,\theta(t)$, the integral can be evaluated explicitly to give
$a(t)=e^{t/t_{0}}\left[b-\frac{f}{m}\left(1-e^{-t/t_{0}}\right)\theta(t)\right],$ (77)
where $b=a(0)$ is an integration constant and $\theta(t)$ is the Heaviside step function. The exponentially growing factor $e^{t/t_{0}}$ corresponds to the unphysical runaway mode, which is removed by the reduction–of–order procedure or by imposing suitable initial conditions that suppress it.
Appendix B Electromagnetic Field Tensor and Metric Conventions
The electromagnetic field tensor $F_{\mu\nu}$ compactly encodes the electric and magnetic fields in relativistic form. For a metric signature $\eta_{\mu\nu}=\mathrm{diag}(-,+,+,+)$, the components of $F_{\mu\nu}$ are defined in terms of the scalar potential $\phi$ and the vector potential $\mathbf{A}$ as
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}.$ (78)
Raising both indices using $\eta^{\mu\nu}$ gives $F^{\mu\nu}=\eta^{\mu\alpha}\eta^{\nu\beta}F_{\alpha\beta}$.
With the above metric convention, the explicit matrix representation of $F^{\mu\nu}$ is
$F^{\mu\nu}=\begin{pmatrix}0&E_{1}/c&E_{2}/c&E_{3}/c\\ -E_{1}/c&0&B_{3}&-B_{2}\\ -E_{2}/c&-B_{3}&0&B_{1}\\ -E_{3}/c&B_{2}&-B_{1}&0\end{pmatrix}$, (79)
where the three electric field components $(E_{1},E_{2},E_{3})$ and magnetic field components $(B_{1},B_{2},B_{3})$ represent the six independent degrees of freedom of the tensor.
Under a gauge transformation of the four–potential,
$A_{\mu}\;\to\;A_{\mu}^{\prime}=A_{\mu}+\partial_{\mu}\theta,$ (80)
the field tensor transforms as
$F_{\mu\nu}^{\prime}=\partial_{\mu}A_{\nu}^{\prime}-\partial_{\nu}A_{\mu}^{\prime}=F_{\mu\nu}-(\partial_{\mu}\partial_{\nu}\theta-\partial_{\nu}\partial_{\mu}\theta)=F_{\mu\nu},$ (81)
confirming that $F_{\mu\nu}$ is gauge invariant.
Sign conventions: The signs of the matrix elements depend on the metric convention. In texts that use $\eta_{\mu\nu}=(+,-,-,-)$ (e.g. Jackson, Classical Electrodynamics, 3rd ed., 1998), the definition of $A_{\mu}$ and the components of $F_{\mu\nu}$ differ by an overall minus sign.
Appendix C Details on the Darwin force correction at 1PN
In this section, we show the detailed steps leading to Eq. (63) of the main text. We start from the Darwin interaction Hamiltonian, along with the useful notations as presented below:
$H_{D,\mathrm{int}}=-\frac{1}{2}\sum_{a\neq b}\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}m_{a}m_{b}c^{2}}\;\frac{1}{r_{ab}}\,\mathcal{P}_{ab},\qquad\mathbf{r}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b},\quad r_{ab}=|\mathbf{r}_{ab}|$ (82)
where
$\mathcal{P}_{ab}:=\mathbf{p}_{a}\!\cdot\!\mathbf{p}_{b}+\frac{(\mathbf{p}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}}.$ (83)
Using the symmetry under $a\leftrightarrow b$ we can write
$H_{D,\mathrm{int}}=-\sum_{a<b}H_{ab},\qquad H_{ab}:=\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}m_{a}m_{b}c^{2}}\;\frac{1}{r_{ab}}\,\mathcal{P}_{ab}.$ (84)
Within 1PN accuracy we may replace $\mathbf{p}_{a}\simeq m_{a}\mathbf{v}_{a}$ and $\mathbf{p}_{b}\simeq m_{b}\mathbf{v}_{b}$, which yields
$\mathcal{P}_{ab}=m_{a}m_{b}\left[\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+\frac{(\mathbf{v}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{v}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}}\right].$ (85)
The masses thus cancel in the full expression, and with $k=1/(4\pi\varepsilon_{0})$ we obtain
$H_{ab}=-\frac{k\,q_{a}q_{b}}{2c^{2}}\,\frac{1}{r_{ab}}\left[\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+\frac{(\mathbf{v}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{v}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}}\right].$ (86)
The Darwin contribution to the canonical equation for $\mathbf{p}_{a}$ is thus
$\dot{\mathbf{p}}_{a}\Big|_{1\mathrm{PN}}=-\,\frac{\partial H_{D,\mathrm{int}}}{\partial\mathbf{x}_{a}}=-\sum_{b\neq a}\frac{\partial H_{ab}}{\partial\mathbf{x}_{a}}.$ (87)
Since $H_{ab}$ depends on $\mathbf{x}_{a}$ only through $\mathbf{r}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b}$, we may replace
$\frac{\partial}{\partial\mathbf{x}_{a}}\;=\;\frac{\partial}{\partial\mathbf{r}_{ab}}\equiv\bm{\nabla}_{\mathbf{r}},\qquad\mathbf{r}\equiv\mathbf{r}_{ab},\quad r\equiv r_{ab},\quad\hat{\mathbf{r}}\equiv\hat{\mathbf{r}}_{ab}=\frac{\mathbf{r}}{r}.$ (88)
For a single pair $(a,b)$, Eq. (86) becomes
$H_{ab}=-A\Big[S_{1}(\mathbf{r})+S_{2}(\mathbf{r})\Big],\qquad A:=\frac{k\,q_{a}q_{b}}{2c^{2}},$ (89)
with the associated quantities
$S_{1}(\mathbf{r}):=\frac{\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}}{r},\qquad S_{2}(\mathbf{r}):=\frac{(\mathbf{v}_{a}\!\cdot\!\mathbf{r})(\mathbf{v}_{b}\!\cdot\!\mathbf{r})}{r^{3}}.$ (90)
We can now write
$\dot{\mathbf{p}}_{a}^{(ab)}\Big|_{1\mathrm{PN}}=-\frac{\partial H_{ab}}{\partial\mathbf{x}_{a}}=-\bm{\nabla}_{\mathbf{r}}H_{ab}=A\,\bm{\nabla}_{\mathbf{r}}\big[S_{1}(\mathbf{r})+S_{2}(\mathbf{r})\big].$ (91)
Since $\mathbf{v}_{a}$ and $\mathbf{v}_{b}$ are independent of $\mathbf{r}$, we have
$\bm{\nabla}_{\mathbf{r}}S_{1}=(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b})\,\bm{\nabla}_{\mathbf{r}}\!\left(\frac{1}{r}\right)=(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b})\left(-\frac{\hat{\mathbf{r}}}{r^{2}}\right)=-\frac{\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}}{r^{2}}\,\hat{\mathbf{r}}$. (92)
We now define two intermediate scalars
$\alpha:=\mathbf{v}_{a}\!\cdot\!\mathbf{r},\qquad\beta:=\mathbf{v}_{b}\!\cdot\!\mathbf{r}.$ (93)
Then $S_{2}=\alpha\beta/r^{3}$ and
$\bm{\nabla}_{\mathbf{r}}S_{2}=\frac{1}{r^{3}}\bm{\nabla}_{\mathbf{r}}(\alpha\beta)+\alpha\beta\,\bm{\nabla}_{\mathbf{r}}\!\left(\frac{1}{r^{3}}\right).$ (94)
Because $\bm{\nabla}_{\mathbf{r}}\alpha=\mathbf{v}_{a}$ and $\bm{\nabla}_{\mathbf{r}}\beta=\mathbf{v}_{b}$, we have
$\bm{\nabla}_{\mathbf{r}}(\alpha\beta)=\beta\,\mathbf{v}_{a}+\alpha\,\mathbf{v}_{b}=(\mathbf{v}_{b}\!\cdot\!\mathbf{r})\,\mathbf{v}_{a}+(\mathbf{v}_{a}\!\cdot\!\mathbf{r})\,\mathbf{v}_{b}.$ (95)
Using the standard result
$\bm{\nabla}_{\mathbf{r}}\!\left(\frac{1}{r^{3}}\right)=-3\,\frac{\mathbf{r}}{r^{5}}=-\frac{3}{r^{4}}\,\hat{\mathbf{r}}$, (96)
we thus obtain
$\bm{\nabla}_{\mathbf{r}}S_{2}=\frac{1}{r^{3}}\Big[(\mathbf{v}_{b}\!\cdot\!\mathbf{r})\,\mathbf{v}_{a}+(\mathbf{v}_{a}\!\cdot\!\mathbf{r})\,\mathbf{v}_{b}\Big]-3\,\frac{(\mathbf{v}_{a}\!\cdot\!\mathbf{r})(\mathbf{v}_{b}\!\cdot\!\mathbf{r})}{r^{5}}\,\mathbf{r}.$ (97)
In order to reach Eq. (63) of the main text, it is convenient to express the scalar products with $\hat{\mathbf{r}}$:
$\mathbf{v}_{a}\!\cdot\!\mathbf{r}=r\,(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a}),\qquad\mathbf{v}_{b}\!\cdot\!\mathbf{r}=r\,(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b}).$ (98)
Using these, the previous expression becomes
$\bm{\nabla}_{\mathbf{r}}S_{2}$ $=\frac{1}{r^{3}}\Big[r(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})\,\mathbf{v}_{a}+r(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})\,\mathbf{v}_{b}\Big]-3\,\frac{r^{2}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})}{r^{5}}\,\mathbf{r}$
$=\frac{1}{r^{2}}\Big[\mathbf{v}_{a}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})-3\,\hat{\mathbf{r}}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})\Big].$ (99)
Adding Eqs. (92) and (99) we find
$\bm{\nabla}_{\mathbf{r}}\big(S_{1}+S_{2}\big)$ $=-\frac{\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}}{r^{2}}\,\hat{\mathbf{r}}$
$\quad+\frac{1}{r^{2}}\Big[\mathbf{v}_{a}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})-3\,\hat{\mathbf{r}}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})\Big]$
$=\frac{1}{r^{2}}\Big[\mathbf{v}_{a}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})-\hat{\mathbf{r}}\big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})\big)\Big].$
Therefore the Darwin contribution to $\dot{\mathbf{p}}_{a}$ from particle $b$ is
$\dot{\mathbf{p}}_{a}^{(ab)}\Big|_{1\mathrm{PN}}=\frac{k\,q_{a}q_{b}}{2c^{2}r_{ab}^{2}}\Big[\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})-\hat{\mathbf{r}}_{ab}\big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\big)\Big].$ (100)
Summing over $b\neq a$ gives the full Darwin 1PN correction
$\dot{\mathbf{p}}_{a}\Big|_{1\mathrm{PN}}=k\sum_{b\neq a}\frac{q_{a}q_{b}}{2c^{2}r_{ab}^{2}}\Big[\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})-\hat{\mathbf{r}}_{ab}\big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\big)\Big],$ (101)
which is Eq. (63) in the main text. The Darwin 1PN correction has a simple physical interpretation. The terms proportional to $\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})$ and $\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})$ represent the magnetostatic, or current–current, interaction between moving charges: each particle experiences a velocity–dependent force generated by the motion of the other charge. The remaining contribution, proportional to $\hat{\mathbf{r}}_{ab}\big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\big)$, ensures that the interaction has the unique Poincaré–consistent form required at order $v^{2}/c^{2}$.
The velocity–dependent force derived in Eq. (63) of the main text is the classical, spin–independent component of the Darwin interaction. It represents the leading relativistic correction to Coulomb’s law obtained by expanding the full electromagnetic interaction between two charges to order $v^{2}/c^{2}$. This correction encodes conservative retardation effects in the near zone and admits three equivalent interpretations.
First, starting from classical electrodynamics, the retarded scalar and vector potentials of a point charge $q_{b}$ moving with velocity $\mathbf{v}_{b}$ are
$\Phi_{b}(\mathbf{x},t)=k\,\frac{q_{b}}{R_{b}-\mathbf{R}_{b}\!\cdot\!\bm{\beta}_{b}},\qquad\mathbf{A}_{b}(\mathbf{x},t)=\frac{k\,q_{b}}{c}\,\frac{\mathbf{v}_{b}}{R_{b}-\mathbf{R}_{b}\!\cdot\!\bm{\beta}_{b}},$ (102)
where $\mathbf{R}_{b}=\mathbf{x}-\mathbf{x}_{b}(t_{\rm ret})$ and $\bm{\beta}_{b}=\mathbf{v}_{b}/c$. Expanding in the near zone for small velocities and small retardation, and retaining terms through order $v^{2}/c^{2}$, one finds schematically
$\frac{1}{R_{b}-\mathbf{R}_{b}\!\cdot\!\bm{\beta}_{b}}=\frac{1}{r_{ab}}\left[1+\frac{\mathbf{v}_{b}\!\cdot\!\mathbf{r}_{ab}}{c\,r_{ab}}+\frac{v_{b}^{2}}{2c^{2}}-\frac{(\mathbf{v}_{b}\!\cdot\!\mathbf{r}_{ab})^{2}}{2c^{2}r_{ab}^{2}}+\mathcal{O}(c^{-3})\right].$ (103)
Constructing the electric field $\mathbf{E}_{b}=-\nabla\Phi_{b}-\partial\mathbf{A}_{b}/\partial t$, the field of particle $b$ evaluated at the position of particle $a$ contains, in addition to the Coulomb term, a velocity–dependent 1PN contribution proportional to
$\frac{kq_{b}}{c^{2}r_{ab}^{2}}\Big[\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})+\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})-\hat{\mathbf{r}}_{ab}(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b})-3\hat{\mathbf{r}}_{ab}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\Big].$
Multiplication by $q_{a}$ reproduces exactly the Darwin force in Eq. (63). Thus, the Darwin interaction corresponds to the $v^{2}/c^{2}$ expansion of the full retarded electromagnetic interaction when radiation emission is neglected.
Second, the same interaction arises naturally from an effective–action viewpoint. After integrating out the electromagnetic field (e.g. in Coulomb gauge) and expanding the near–zone effective action to order $v^{2}/c^{2}$, one obtains an instantaneous current–current interaction,
$S=-\sum_{a}m_{a}c^{2}\!\int\!dt\,\sqrt{1-\frac{v_{a}^{2}}{c^{2}}}+\frac{1}{8\pi k}\!\int\!d^{3}x\,d^{3}x^{\prime}\frac{\mathbf{J}(\mathbf{x},t)\!\cdot\!\mathbf{J}(\mathbf{x}^{\prime},t)}{|\mathbf{x}-\mathbf{x}^{\prime}|}+\cdots,$ (104)
where $\mathbf{J}$ is the charge current. Expanding the matter action to 1PN order and substituting the point–particle current $\mathbf{J}(\mathbf{x},t)=\sum_{a}q_{a}\mathbf{v}_{a}\,\delta^{3}(\mathbf{x}-\mathbf{x}_{a})$ yields the Darwin Lagrangian
$L_{\rm Darwin}=\frac{k}{2c^{2}}\sum_{a\neq b}\frac{q_{a}q_{b}}{r_{ab}}\Big[\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+(\mathbf{v}_{a}\!\cdot\!\hat{\mathbf{r}}_{ab})(\mathbf{v}_{b}\!\cdot\!\hat{\mathbf{r}}_{ab})\Big].$ (105)
The corresponding Hamiltonian (via a Legendre transform) coincides with $H_{D,\mathrm{int}}$, and differentiation with respect to $\mathbf{x}_{a}$ reproduces Eq. (63), making explicit that the Darwin term originates purely from the instantaneous, non–radiative sector of the electromagnetic field.
Third, in quantum electrodynamics the spin–independent part of the Breit Hamiltonian is
$H_{\rm Breit}=\frac{kq_{a}q_{b}}{2m_{a}m_{b}c^{2}r_{ab}}\left[\mathbf{p}_{a}\!\cdot\!\mathbf{p}_{b}+\frac{(\mathbf{p}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}}\right].$ (106)
which has exactly the same operator structure as the classical Darwin Hamiltonian. In the $\hbar\!\to\!0$ limit, identifying $\mathbf{p}_{a}=m_{a}\mathbf{v}_{a}$ reduces this expression to the interaction (86), and hence to the same equations of motion.
In all three approaches, the resulting velocity–dependent force takes the universal form
$\dot{\mathbf{p}}_{a}\Big|_{1\mathrm{PN}}\propto\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})-\hat{\mathbf{r}}_{ab}\Big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\Big),$ (107)
demonstrating that the Darwin interaction is the complete conservative correction to the classical two–body problem at order $v^{2}/c^{2}$, coinciding with the standard Darwin/Breit interaction in both classical and quantum electrodynamics.
Appendix D Closer look into the Four–acceleration and the Covariant Larmor Power
In this appendix we derive the explicit form of the four–acceleration $a^{\mu}=du^{\mu}/d\tau$ in terms of the ordinary velocity $\mathbf{v}$ and acceleration $\mathbf{a}=d\mathbf{v}/dt$, and we show how the covariant expression for the radiated power reduces to the standard Liénard generalization of the Larmor formula. The four–velocity of a particle with ordinary velocity $\mathbf{v}$ is
$u^{\mu}=\gamma(c,\mathbf{v}),\qquad\gamma=\frac{1}{\sqrt{1-v^{2}/c^{2}}}.$ (108)
Since $d\tau=dt/\gamma$, the derivative with respect to proper time may be written as
$\frac{d}{d\tau}=\gamma\frac{d}{dt}.$ (109)
We first compute the time component. Because $u^{0}=\gamma c$, we obtain
$a^{0}=\frac{du^{0}}{d\tau}=\gamma\frac{d}{dt}(\gamma c)=\gamma c\frac{d\gamma}{dt}.$ (110)
Using
$\frac{d\gamma}{dt}=\gamma^{3}\frac{\mathbf{v}\cdot\mathbf{a}}{c^{2}},$ (111)
we find
$a^{0}=\gamma^{4}\frac{\mathbf{v}\cdot\mathbf{a}}{c}.$ (112)
For the spatial components, begin with $\mathbf{u}=\gamma\mathbf{v}$. Then
$\mathbf{a}^{(4)}$ $=\frac{d\mathbf{u}}{d\tau}=\gamma\frac{d}{dt}(\gamma\mathbf{v})=\gamma(\gamma\mathbf{a}+\mathbf{v}\,\dot{\gamma}),$ (113)
$\dot{\gamma}$ $=\gamma^{3}\frac{\mathbf{v}\cdot\mathbf{a}}{c^{2}}.$
This gives
$\mathbf{a}^{(4)}=\gamma^{2}\mathbf{a}+\gamma^{4}\frac{\mathbf{v}\cdot\mathbf{a}}{c^{2}}\mathbf{v}.$ (114)
Combining Eqs. (112) and (114), the four–acceleration takes the explicit form
$a^{\mu}=\left(\gamma^{4}\frac{\mathbf{v}\cdot\mathbf{a}}{c},\;\gamma^{2}\mathbf{a}+\gamma^{4}\frac{\mathbf{v}\cdot\mathbf{a}}{c^{2}}\mathbf{v}\right).$ (115)
Using the metric signature (,+,+,+)(-,+,+,+),
$a_{\mu}a^{\mu}=-(a^{0})^{2}+|\mathbf{a}^{(4)}|^{2}.$ (116)
Substituting Eq. (115) and simplifying, one obtains the standard invariant form
$a_{\mu}a^{\mu}=\gamma^{6}\!\left(a^{2}-\frac{(\mathbf{v}\times\mathbf{a})^{2}}{c^{2}}\right)=\gamma^{4}\!\left(a^{2}+\gamma^{2}\frac{(\mathbf{v}\cdot\mathbf{a})^{2}}{c^{2}}\right).$ (117)
Both expressions are equivalent via $(\mathbf{v}\times\mathbf{a})^{2}=v^{2}a^{2}-(\mathbf{v}\cdot\mathbf{a})^{2}$. The covariant form of the radiated power for a particle with rest mass $m$ and charge $q$ is
$P=-\frac{2}{3}\frac{q^{2}}{4\pi\epsilon_{0}m^{2}c^{3}}\left(\frac{dp_{\mu}}{d\tau}\frac{dp^{\mu}}{d\tau}\right),\qquad p^{\mu}=mu^{\mu}.$ (118)
Since $dp^{\mu}/d\tau=ma^{\mu}$, the factors of $m$ cancel and we obtain
$P=-\frac{2}{3}\frac{q^{2}}{4\pi\epsilon_{0}c^{3}}\,a_{\mu}a^{\mu}.$ (119)
Substituting Eq. (117), the radiated power becomes
$P=\frac{q^{2}}{6\pi\epsilon_{0}c^{3}}\,\gamma^{6}\!\left(a^{2}-\frac{(\mathbf{v}\times\mathbf{a})^{2}}{c^{2}}\right)\\$ (120)
or equivalently,
$P=\frac{q^{2}}{6\pi\epsilon_{0}c^{3}}\,\gamma^{4}\!\left(a^{2}+\gamma^{2}\frac{(\mathbf{v}\cdot\mathbf{a})^{2}}{c^{2}}\right).$ (121)
In the nonrelativistic limit $\gamma\to 1$ and $v\ll c$, both forms reduce to the familiar Larmor result,
$P=\frac{q^{2}a^{2}}{6\pi\epsilon_{0}c^{3}}.$ (122)