### v5

# Post-Newtonian Dynamics of Radiating Charges: Canonical Formulation and Binary Inspiral Laws

Abstract
We construct an explicit electromagnetic analogue of the post-Newtonian Hamiltonian framework widely used in gravitational-wave physics. Starting from the Lorentz–Dirac equation and implementing Landau–Lifshitz order reduction, we derive the near-zone 1.51.5PN dipole radiation–reaction force and combine it with the Darwin Hamiltonian through 11PN order to obtain a closed canonical $N$-body phase-space system. The equations are directly implementable, strictly conservative when dissipation is switched off, and exhibit monotonic energy loss, secular inspiral, and circularization when radiation reaction is included. For binaries we derive analytic circular and eccentric inspiral laws, including 11PN conservative corrections to the dipole-driven inspiral and merger time. Extending to charged compact binaries in Einstein–Maxwell theory, we combine the 22PN ADM-type conservative Hamiltonian with leading 1.51.5PN dipole dissipation and gravitational quadrupole flux, obtaining gauge-invariant energy–frequency relations, closed-form circular inspiral laws, and a dipole–quadrupole crossover scale that separates electromagnetic and gravitational flux dominated inspirals.
I Introduction
Radiation reaction—the backreaction of emitted radiation on the motion of its source—is a unifying element of classical field theories, ranging from electrodynamics to the dissipative inspiral of compact binaries in general relativity [26, 22, 6, 16]. In gravitational-wave physics it is routinely incorporated within a canonical “Hamiltonian plus radiation–reaction” paradigm, implemented in PN/ADM and EOB frameworks [23, 24, 25, 27, 28, 29, 30, 31, 32] and complemented by nonconservative action/EFT formulations [33, 34]. In electrodynamics, a local self-force law arises from Dirac’s world-tube momentum balance and equivalent radiative-field decompositions, leading to the covariant Lorentz–Dirac (LD) equation [7, 9, 8, 10]. Its third-order character admits runaway and preaccelerated solutions, and the standard resolution is Landau–Lifshitz (LL) reduction of order, which yields a causal second-order dynamics perturbatively equivalent to LD [10, 13, 11, 12, 15]. These structural analogies motivate the main goal of this work: to construct an explicit, directly implementable electromagnetic analogue of the PN canonical paradigm and to use it for analytic and numerical studies for dissipative long-range dynamics.
We begin by adopting the order-reduced LL dynamics and taking its nonrelativistic near-zone limit, where dissipation first enters at absolute 1.51.5PN order and is governed by the electric dipole moment, yielding a dipole radiation–reaction force consistent with Larmor energy balance (up to the Schott term) [10, 8, 11]. We then build an $N$-body phase-space system in direct analogy with PN gravity: a conservative Hamiltonian sector truncated at Darwin (11PN) order plus an explicit non-Hamiltonian dipole radiation–reaction force at 1.51.5PN order, expressed entirely in canonical variables. This construction is designed to be operational (ready for numerical integration) and diagnostically transparent: in the conservative limit it preserves the Darwin Hamiltonian without secular drift, while with radiation reaction it exhibits monotonic energy loss, secular inspiral, and robust circularization. Specializing to $N=2$ we recover the characteristic dipole structure proportional to the charge-to-mass asymmetry and the exact suppression of the 1.51.5PN radiation reaction when $q_{1}/m_{1}=q_{2}/m_{2}$, in agreement with post-Coulombian analyses [35]. Using the same dissipative input, we derive analytic inspiral laws for circular and eccentric binaries, including secular evolution and closed-form integrals for $a(e)$ and $t(e)$ at leading order, and we validate these results against direct integrations of the full canonical equations.
To connect this purely electromagnetic system to relativistic charged compact binaries, we also examine inspiral laws in Einstein–Maxwell theory by combining the 22PN ADM-type center-of-mass Hamiltonian with leading 1.51.5PN dipole dissipation, following recent results for charged black-hole binaries [36]. This yields circular energy–frequency relations in terms of an effective coupling and analytic circular inspiral laws through 22PN conservative order. Including gravitational quadrupole flux in the balance law then identifies a dipole–quadrupole crossover scale that depends only on charge-to-mass asymmetry, interpolating between the dipole scaling $\dot{\Omega}\propto\Omega^{3}$ and the GR quadrupole scaling $\dot{\Omega}\propto\Omega^{11/3}$ [26]. We estimate the dipole–quadrupole crossover frequency as $f_{\rm cross}\simeq 2.2\times 10^{3}\,\mathrm{Hz}\,(M_{\odot}/M)\,|\eta_{2}-\eta_{1}|^{3}/(1-\eta_{1}\eta_{2})$, where $M$ is the total mass and $\eta_{A}\equiv\sqrt{k_{E}/G}\,(q_{A}/m_{A})$ denotes the dimensionless charge-to-mass ratio of body $A$. For a $60\,M_{\odot}$ binary this reduces to $f_{\rm cross}\simeq 36\,\mathrm{Hz}\,|\eta_{2}-\eta_{1}|^{3}/(1-\eta_{1}\eta_{2})$. The cubic dependence on the charge-to-mass asymmetry implies that for $|\eta_{A}|\ll 1$ the crossover lies far below the ground-based detector band, so the inspiral remains quadrupole dominated; dipole radiation effects enter the LIGO/Virgo band only if the charge-to-mass asymmetry $|\eta_{2}-\eta_{1}|$ is of order unity, i.e. in highly asymmetric and near-extremal configurations. It is also noteworthy that order-reduced electromagnetic radiation-reaction dynamics admits special analytic structures in certain limits, including exact reductions to Painlevé transcendents in Coulombic settings [37, 38]. Such features have motivated recent conjectures on universality in binary black-hole coalescence [39, 40, 41]. The present framework provides an explicit relativistic realization of dipole-driven inspiral within an explicit post-Newtonian structure.
Organization of the paper.
In Sec. II we review radiation reaction for point charges, deriving the Lorentz–Dirac equation, implementing the Landau–Lifshitz order-reduction procedure, and extracting the 1.51.5PN near-zone dipole radiation–reaction force. Section III combines this dissipative sector with the Darwin Hamiltonian to construct a closed 11PN+1.5+1.5PN canonical $N$-body phase-space system. In Sec. IV we specialize to $N=2$, derive the explicit dipole radiation–reaction structure for binaries, and connect with existing post-Coulombian and Einstein–Maxwell results. Sections V and VI develop analytic circular and eccentric inspiral laws at leading dipole order, while Sec. VII incorporates 11PN conservative corrections into the eccentric evolution. We then extend the framework to relativistic charged binaries in Einstein–Maxwell theory: Sec. VIII presents the ADM-type center-of-mass Hamiltonian through 22PN order together with the leading 1.51.5PN radiation reaction. Section IX analyzes the dipole–quadrupole crossover scale, and Secs. X and XI derive the circular inspiral laws through 22PN conservative order, first in the dipole-dominated regime and then including both dipole and quadrupole radiation.
Technical details and validation material are collected in the Appendices: Appendix A discusses runaway solutions of the Lorentz–Dirac equation; Appendix B presents single-particle consistency tests of the Landau–Lifshitz dynamics; Appendix C provides details of the 11PN Darwin force correction; and Appendix D summarizes numerical simulations of the charge-neutral binary.
II Radiation reaction: Lorentz–Dirac, Landau–Lifshitz reduction, and the near-zone limit
Radiation reaction for a point charge may be derived either by isolating the radiative field via the half–retarded minus half–advanced decomposition [10, 8] or by Dirac’s world–tube momentum balance [7, 9]. Both routes lead to the covariant Lorentz–Dirac (LD) equation
$ma^{\alpha}=F^{\alpha}_{\rm ext}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta},$ (1)
where $u^{\alpha}$ is the four–velocity, $a^{\alpha}\!=\!du^{\alpha}/d\tau$, and the projector enforces $u_{\alpha}a^{\alpha}=0$ (with $u_{\alpha}u^{\alpha}=-c^{2}$). Equation (1) is third order and admits runaway and preaccelerated solutions; this is already evident in the nonrelativistic limit,
$a(t)-t_{0}\,\dot{a}(t)=\frac{1}{m}F_{\rm ext}(t),\qquad t_{0}=\frac{1}{4\pi\varepsilon_{0}}\,\frac{2q^{2}}{3mc^{3}},$ (2)
for which a step force $F_{\rm ext}(t)=f\,\theta(t)$ yields
$a(t)=e^{t/t_{0}}\!\left[b-\frac{f}{m}\bigl(1-e^{-t/t_{0}}\bigr)\theta(t)\right],\qquad a(t)=\frac{f}{m}\left[\theta(-t)e^{t/t_{0}}+\theta(t)\right]\ \ (b=f/m),$ (3)
showing, respectively, runaway growth for generic $b$ and preacceleration for the tuned (no–runaway) choice.
The physical dynamics is obtained by the Landau–Lifshitz (LL) reduction of order [10, 13], treating the self–force perturbatively and replacing $\dot{a}^{\alpha}$ by the proper–time derivative of the leading Lorentz–force acceleration. This gives the second–order LL equation
$ma^{\alpha}=F^{\alpha}_{\rm ext}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}mc^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)F^{\beta}_{{\rm ext},\gamma}u^{\gamma},$ (4)
which is causal and free of runaway/preacceleration pathologies [13, 12, 11]. For electromagnetic forcing $F^{\alpha}_{\rm ext}=q\,F^{\alpha}{}_{{\rm ext}\,\mu}u^{\mu}$, one convenient explicit form is
$ma^{\alpha}$ $=q\,F^{\alpha}{}_{{\rm ext}\,\mu}\,u^{\mu}+q\,t_{0}\,F^{\alpha}{}_{{\rm ext}\,\mu,\nu}\,u^{\mu}u^{\nu}+\frac{q^{2}t_{0}}{m}\left(\delta^{\alpha}{}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)F^{\beta}{}_{{\rm ext}\,\mu}\,F^{\mu}{}_{{\rm ext}\,\nu}\,u^{\nu},$ (5)
with $t_{0}$ given in (2) (one intermediate term vanishes by antisymmetry of $F_{\mu\nu}$). In Dirac’s world–tube picture, the electromagnetic momentum flux through a tube $\Sigma$ of radius $r$,
$\frac{dP^{\alpha}_{\rm em}}{d\tau}=\int_{\Sigma}T^{\alpha\beta}_{\rm em}\,d\Sigma_{\beta},$ (6)
splits into a divergent near–field piece and a finite radiative piece,
$\frac{dP^{\alpha}_{\rm em}}{d\tau}=\frac{q^{2}}{8\pi\varepsilon_{0}rc^{2}}\,a^{\alpha}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\frac{a^{2}}{c^{2}}\,u^{\alpha},$ (7)
where the $1/r$ term is absorbed into the renormalized mass and the finite term encodes radiated energy–momentum. The equivalent “Dirac” form of the LD equation,
$ma^{\alpha}=F^{\alpha}_{\rm ext}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\left(\dot{a}^{\alpha}-\frac{a^{2}}{c^{2}}u^{\alpha}\right),$ (8)
reduces to (1) upon using
$a^{2}\equiv a_{\alpha}a^{\alpha}=-\dot{a}_{\alpha}u^{\alpha}.$ (9)
In the near zone ($v\ll c$), dissipative electromagnetic radiation reaction first enters at 1.51.5PN order and is controlled by the electric dipole moment
$\mathbf{d}(t)=\sum_{a}q_{a}\,\mathbf{x}_{a}(t).$ (10)
To leading order, the 1.51.5PN near–zone field gives the dipole radiation–reaction force
$\mathbf{F}_{a}^{(1.5{\rm PN})}=\frac{q_{a}}{4\pi\varepsilon_{0}}\frac{2}{3c^{3}}\,\dddot{\mathbf{d}}(t),$ (11)
and the associated mechanical energy loss equals the dipole (Larmor) power,
$\frac{dE}{dt}=-\frac{1}{4\pi\varepsilon_{0}}\frac{2}{3c^{3}}\,|\ddot{\mathbf{d}}(t)|^{2},$ (12)
up to a total time derivative (Schott energy). In our single–particle tests we verify the covariant energy balance relation
$\Delta K=W_{\rm ext}-E_{\rm rad},$ (13)
and in the remainder of this paper we adopt the order–reduced LL dynamics as the dissipative input for the post–Newtonian many–body framework developed below.
III The N-body Darwin–radiation-reaction system
We construct the electromagnetic NN–body dynamics in direct analogy with the PN/EOB treatment of compact binaries: a conservative Hamiltonian sector through 11PN ($\mathcal{O}(c^{-2})$), generated by the Darwin Hamiltonian, augmented by a non-Hamiltonian 1.51.5PN dipole radiation–reaction force ($\mathcal{O}(c^{-3})$). We work in SI units, with $k=(4\pi\varepsilon_{0})^{-1}$ and $\mathbf{r}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b},\;r_{ab}=|\mathbf{r}_{ab}|,\;\hat{\mathbf{r}}_{ab}=\mathbf{r}_{ab}/r_{ab}.$
The conservative dynamics is generated by the Darwin Hamiltonian
$H_{\rm D}$ $=\sum_{a}\!\left(\frac{\mathbf{p}_{a}^{2}}{2m_{a}}-\frac{\mathbf{p}_{a}^{4}}{8m_{a}^{3}c^{2}}\right)+\frac{1}{2}\sum_{a\neq b}k\,\frac{q_{a}q_{b}}{r_{ab}}$
$\quad-\frac{1}{2}\sum_{a\neq b}k\,\frac{q_{a}q_{b}}{2m_{a}m_{b}c^{2}}\,\frac{1}{r_{ab}}\!\left[\mathbf{p}_{a}\!\cdot\!\mathbf{p}_{b}+\frac{(\mathbf{p}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}}\right].$ (14)
The leading dissipative contribution is the dipole radiation–reaction force, evaluated by order reduction and expressed entirely in canonical variables,
$\mathbf{F}_{a}^{(1.5\mathrm{PN})}(t)=\frac{1}{(4\pi\varepsilon_{0})^{2}}\frac{2q_{a}}{3c^{3}}\sum_{b}\sum_{c\neq b}\frac{q_{b}^{2}q_{c}}{m_{b}r_{bc}^{3}}\Big[\tfrac{\mathbf{p}_{b}}{m_{b}}-\tfrac{\mathbf{p}_{c}}{m_{c}}$
$-3\hat{\mathbf{r}}_{bc}\big(\hat{\mathbf{r}}_{bc}\!\cdot\!(\tfrac{\mathbf{p}_{b}}{m_{b}}-\tfrac{\mathbf{p}_{c}}{m_{c}})\big)\Big],$ (15)
which is fully consistent with electromagnetic energy balance.
Collecting results, the closed 11PN+1.5+1.5PN equations of motion in canonical phase space are
$\dot{\mathbf{x}}_{a}$ $=\frac{\partial H_{\rm D}}{\partial\mathbf{p}_{a}}=\frac{\mathbf{p}_{a}}{m_{a}}-\frac{\mathbf{p}_{a}^{2}\,\mathbf{p}_{a}}{2m_{a}^{3}c^{2}}$
$\quad-\sum_{b\neq a}\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}}\frac{1}{m_{a}m_{b}c^{2}}\frac{1}{r_{ab}}\left[\mathbf{p}_{b}+\mathbf{r}_{ab}\frac{\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab}}{r_{ab}^{2}}\right],$
$\dot{\mathbf{p}}_{a}$ $=-\frac{\partial H_{\rm D}}{\partial\mathbf{x}_{a}}+\mathbf{F}_{a}^{(1.5\mathrm{PN})}(\mathbf{x},\mathbf{p}),$ (16)
with the conservative force written explicitly as
$-\frac{\partial H_{\rm D}}{\partial\mathbf{x}_{a}}$ $=\sum_{b\neq a}k\,\frac{q_{a}q_{b}}{r_{ab}^{2}}\,\hat{\mathbf{r}}_{ab}$ (17)
$\quad+k\sum_{b\neq a}\frac{q_{a}q_{b}}{2c^{2}r_{ab}^{2}}\Bigg[\frac{\mathbf{p}_{a}}{m_{a}}\left(\hat{\mathbf{r}}_{ab}\!\cdot\!\frac{\mathbf{p}_{b}}{m_{b}}\right)+\frac{\mathbf{p}_{b}}{m_{b}}\left(\hat{\mathbf{r}}_{ab}\!\cdot\!\frac{\mathbf{p}_{a}}{m_{a}}\right)$
$\qquad\qquad-\hat{\mathbf{r}}_{ab}\left(\frac{\mathbf{p}_{a}\!\cdot\!\mathbf{p}_{b}}{m_{a}m_{b}}+3\left(\hat{\mathbf{r}}_{ab}\!\cdot\!\frac{\mathbf{p}_{a}}{m_{a}}\right)\left(\hat{\mathbf{r}}_{ab}\!\cdot\!\frac{\mathbf{p}_{b}}{m_{b}}\right)\right)\Bigg].$
Equations (16)–(17), together with (15), constitute the fully explicit and implementable 11PN+1.5+1.5PN phase–space system used in our simulations. When $\mathbf{F}_{a}^{(1.5\mathrm{PN})}$ is switched off, the evolution is strictly conservative and exhibits no secular drift in $H_{\rm D}$, providing a stringent validation of the Darwin sector prior to the inclusion of dissipative radiation reaction.
IV Specializing to charged binaries: $N=2$
We now present a brief analysis of the electromagnetic radiation–reaction term used Eq.(15) in our $N$-body framework, specialized to $N=2$ and connect it explicitly to both the Einstein–Maxwell post–Newtonian system of Placidi et al. [36] and the analysis of Kunze and Spohn [35]. We start from the many-body electromagnetic radiation–reaction force at 1.51.5PN order, written in canonical variables, Eq. (15) and specialize to a binary system with particle labels $a,b,c\in\{1,2\}$ and focus on the radiation–reaction force acting on particle $1$. The double sum contains exactly two nonvanishing contributions, corresponding to the ordered pairs $(b,c)=(1,2)$ and $(b,c)=(2,1)$.
For $(b,c)=(1,2)$ we obtain
$\mathbf{F}_{1}^{(12)}=\frac{2q_{1}}{3(4\pi\varepsilon_{0})^{2}c^{3}}\frac{q_{1}^{2}q_{2}}{m_{1}r_{12}^{3}}\Bigl[\frac{\mathbf{p}_{1}}{m_{1}}-\frac{\mathbf{p}_{2}}{m_{2}}-3\hat{\mathbf{r}}_{12}\bigl(\hat{\mathbf{r}}_{12}\!\cdot(\tfrac{\mathbf{p}_{1}}{m_{1}}-\tfrac{\mathbf{p}_{2}}{m_{2}})\bigr)\Bigr].$ (18)
For the second contribution $(b,c)=(2,1)$ we use $\hat{\mathbf{r}}_{21}=-\hat{\mathbf{r}}_{12}$ and $\mathbf{p}_{2}/m_{2}-\mathbf{p}_{1}/m_{1}=-(\mathbf{p}_{1}/m_{1}-\mathbf{p}_{2}/m_{2})$, which leads to
$\mathbf{F}_{1}^{(21)}=-\frac{2q_{1}}{3(4\pi\varepsilon_{0})^{2}c^{3}}\frac{q_{2}^{2}q_{1}}{m_{2}r_{12}^{3}}\Bigl[\frac{\mathbf{p}_{1}}{m_{1}}-\frac{\mathbf{p}_{2}}{m_{2}}-3\hat{\mathbf{r}}_{12}\bigl(\hat{\mathbf{r}}_{12}\!\cdot(\tfrac{\mathbf{p}_{1}}{m_{1}}-\tfrac{\mathbf{p}_{2}}{m_{2}})\bigr)\Bigr].$ (19)
Adding the two contributions and defining the relative velocity $\mathbf{v}\equiv\mathbf{v}_{1}-\mathbf{v}_{2}=\mathbf{p}_{1}/m_{1}-\mathbf{p}_{2}/m_{2}$, we find
$\mathbf{a}^{\mathrm{RR}}_{1}=\frac{2q_{1}^{2}q_{2}}{3(4\pi\varepsilon_{0})^{2}c^{3}}\frac{1}{m_{1}r^{3}}\left(\frac{q_{1}}{m_{1}}-\frac{q_{2}}{m_{2}}\right)\left[\mathbf{v}-3\hat{\mathbf{r}}\big(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}\big)\right].$ (20)
where $\hat{\mathbf{r}}\equiv\hat{\mathbf{r}}_{12}$. Equation (20) exhibits the characteristic electromagnetic dipole radiation–reaction structure and is proportional to the difference of charge-to-mass ratios. Consequently, when $q_{1}/m_{1}=q_{2}/m_{2}$ the electromagnetic dipole moment is constant and the 1.51.5PN radiation–reaction force vanishes identically. This behavior is observed in the numerical simulations and in agreement with the result of Kunze and Spohn, [35]. On a related note, Placidi et al.  [36] recently derived the 1.51.5PN dissipative acceleration for charged binaries in Einstein–Maxwell theory using Hadamard regularization. Taking the limit $G\to 0$ which isolates the purely electromagnetic contributions, we obtain perfect agreement with their results. The more general formula in Eq. (15) holds for an N-body system of point charges. At leading dissipative order (1.5PN), the radiation-reaction acceleration for the two-body electromagnetic system reduces in relative coordinates to the compact form
[…]
where
$\mathcal{K}=\frac{q_{1}q_{2}(m_{1}q_{2}-m_{2}q_{1})^{2}}{24\pi^{2}c^{3}\epsilon_{0}^{2}m_{1}^{2}m_{2}^{2}}.$ (22)
$|\mathbf{r}|, 𝐫^=𝐫$$\hat{\mathbf{r}}$$\mathbf${r}/r, and $\mathbf{v}=\dot{\mathbf{r}}$. Apart from the electromagnetic prefactor $\mathcal{K}$, the tensorial structure is characteristic of dipole radiation reaction and is the electromagnetic analogue of the quadrupolar Burke–Thorne term in gravity. The relative orbital energy and angular momentum are
$E=\frac{\mu v^{2}}{2}-\frac{\alpha}{r},\qquad\mathbf{L}=\mu\,\mathbf{r}\times\mathbf{v},$ (23)
with reduced mass $\mu=m_{1}m_{2}/(m_{1}+m_{2})$ and Coulomb parameter $\alpha=k\,q_{1}q_{2}$. The dissipative evolution follows from
$\dot{E}=\mu\,\mathbf{v}\cdot\mathbf{a}_{\rm RR},\qquad\dot{\mathbf{L}}=\mu\,\mathbf{r}\times\mathbf{a}_{\rm RR}.$ (24)
Using (21),
$\dot{E}=\mu\,\mathcal{K}\frac{v^{2}-3v_{r}^{2}}{r^{3}}$ (25)
$\dot{L}=\mathcal{K}\frac{L}{r^{3}}$ (26)
where $v_{r}=\hat{\mathbf{r}}\cdot\mathbf{v}$.
V Circular inspiral with 11PN conservative sector
We consider a binary of masses $m_{1},m_{2}$ and charges $q_{1},q_{2}$ interacting electromagnetically. Conservative dynamics is retained through 11PN ($\mathcal{O}(c^{-2})$), while dissipation is driven by the leading electric–dipole flux at absolute order $\mathcal{O}(c^{-3})$. Higher multipoles and relativistic corrections to the dipole formula, which enter at $\mathcal{O}(c^{-5})$, are consistently neglected. This truncation isolates the 11PN conservative correction to the dipole-driven inspiral.
Reducing the Darwin Hamiltonian to the center-of-mass frame ($\mathbf{P}=0$), the relative Hamiltonian through 11PN is
$H$ $=\frac{p_{r}^{2}}{2\mu}+\frac{L^{2}}{2\mu r^{2}}+\frac{\alpha}{r}$
$\quad+\frac{1}{c^{2}}\left[-\frac{1-3\nu}{8\mu^{3}}\left(p_{r}^{2}+\frac{L^{2}}{r^{2}}\right)^{2}+\frac{\alpha}{2\mu Mr}\left(2p_{r}^{2}+\frac{L^{2}}{r^{2}}\right)\right],$ (27)
where $M=m_{1}+m_{2},\;\mu=m_{1}m_{2}/M,\;\nu=\mu/M,$ and $\alpha=q_{1}q_{2}/(4\pi\varepsilon_{0}).$
For circular motion ($p_{r}=0$), one obtains the generalized Kepler law
$\Omega^{2}(r)=\frac{-\alpha}{\mu r^{3}}\left[1+\frac{\alpha}{2\mu c^{2}r}(1-2\nu)\right]+\mathcal{O}(c^{-4}),$ (28)
and the binding energy as a function of frequency
$E(\Omega)=-\frac{1}{2}\mu^{1/3}(-\alpha)^{2/3}\Omega^{2/3}\left[1-\frac{1+\nu}{12c^{2}}\left(\frac{-\alpha}{\mu}\right)^{2/3}\Omega^{2/3}\right]+\mathcal{O}(c^{-4}).$ (29)
The leading dipole flux gives
$\dot{E}=\mu\mathcal{K}\frac{\Omega^{2}}{r},\qquad\mathcal{K}=\frac{q_{1}q_{2}(m_{2}q_{1}-m_{1}q_{2})^{2}}{24\pi^{2}\varepsilon_{0}^{2}c^{3}m_{1}^{2}m_{2}^{2}}.$ (30)
Using $\dot{E}=(dE/d\Omega)\dot{\Omega}$ and (29) yields the 11PN-corrected dipole chirp equation
$\dot{\Omega}=A\,\Omega^{3}\left[1+\frac{B}{c^{2}}\Omega^{2/3}\right]+\mathcal{O}(c^{-4}),$ (31)
with
$A=\frac{3\kappa\mu}{-\alpha},\qquad B=\frac{2-\nu}{6}\left(\frac{-\alpha}{\mu}\right)^{2/3}.$ (32)
Integration gives the time-to-coalescence
$t_{c}-t=\frac{1}{2A}\Omega^{-2}\left[1-\frac{3B}{2c^{2}}\Omega^{2/3}\right]+\mathcal{O}(c^{-4}),$ (33)
and the frequency evolution
$\Omega(t)=(2A\tau)^{-1/2}\left[1-\frac{3B}{4c^{2}}(2A\tau)^{-1/3}\right]+\mathcal{O}(c^{-4}),\qquad\tau=t_{c}-t.$ (34)
The orbital phase $\dot{\Phi}=\Omega$ becomes
$\Phi(t)=\Phi_{c}-2(2A)^{-1/2}\tau^{1/2}+\frac{9B}{2c^{2}}(2A)^{-5/6}\tau^{1/6}+\mathcal{O}(c^{-4}),$ (35)
or equivalently
$\Phi(\Omega)=\Phi_{c}-\frac{1}{A}\Omega^{-1}+\frac{3B}{c^{2}A}\Omega^{-1/3}+\mathcal{O}(c^{-4}).$ (36)
At leading order, $\dot{\Omega}_{\rm EM}\propto\Omega^{3}$ and $\Phi_{\rm EM}(\Omega)\propto\Omega^{-1},$ in contrast with quadrupole-driven gravitational inspiral, $\dot{\Omega}_{\rm GW}\propto\Omega^{11/3}$ and $\Phi_{\rm GW}(\Omega)\propto\Omega^{-5/3}.$ Both theories exhibit a relative $\mathcal{O}(\Omega^{2/3}/c^{2})$ correction, but with distinct theory-dependent coefficients (here proportional to $2-\nu$).
VI Eccentric inspiral at leading order
(a) 3D trajectories with initial (green) and final (red) markers.
(b) Parametric projection in the xxyy plane.
(c) Hamiltonian $H(\tau)$ showing radiative energy loss.
Figure 1: Charge neutral binary, one heavy and one light, with elliptic orbit initial conditions. Radiation reaction from the truncated post Newtonian approach shows clear signature of circularization: (a) Full 3D trajectories; (b) parametric xxyy plane; (c) Hamiltonian $H(\tau)$, decreasing due to radiation. The initial particle positions are shown in green, and the final positions at $\tau=\tau_{f}$ are shown in red.
We treat the secular inspiral of an eccentric electromagnetic binary driven by leading dipole radiation reaction (absolute order $\mathcal{O}(c^{-3})$), while retaining conservative motion at the Coulombic (Kepler) level. In relative variables the energy and angular momentum are
$E=\frac{\mu v^{2}}{2}-\frac{\alpha}{r},\qquad\mathbf{L}=\mu\,\mathbf{r}\times\mathbf{v},$ (37)
where $\mu=m_{1}m_{2}/(m_{1}+m_{2})$ and $\alpha\equiv q_{1}q_{2}/(4\pi\varepsilon_{0})$. The dissipative evolution is obtained from the radiation–reaction acceleration $\mathbf{a}_{\rm RR}$ via
$\dot{E}=\mu\,\mathbf{v}\!\cdot\!\mathbf{a}_{\rm RR},\qquad\dot{\mathbf{L}}=\mu\,\mathbf{r}\times\mathbf{a}_{\rm RR}.$ (38)
Using the compact dipole RR form (cf. Eq. (21)) one finds the instantaneous losses
$\dot{E}=\mu\,\mathcal{K}\,\frac{v^{2}-3v_{r}^{2}}{r^{3}},$ (39)
$\dot{L}=\mathcal{K}\,\frac{L}{r^{3}},$ (40)
where $v_{r}=\hat{\mathbf{r}}\!\cdot\!\mathbf{v}$ and $\mathcal{K}$ is the (constant) electromagnetic dipole prefactor defined in the circular case.
At leading conservative order the motion is Keplerian,
$r=\frac{a(1-e^{2})}{1+e\cos f},$ (41)
with orbital elements satisfying
$E=-\frac{\alpha}{2a},\qquad L^{2}=\mu\alpha a(1-e^{2}),$ (42)
and velocity invariants
$v^{2}=\frac{\alpha}{\mu}\!\left(\frac{2}{r}-\frac{1}{a}\right),\qquad v_{r}^{2}=\frac{\alpha}{\mu a(1-e^{2})}\,e^{2}\sin^{2}f.$ (43)
Orbit averaging, $\langle X\rangle=(1/T)\int_{0}^{2\pi}X(f)(dt/df)\,df$ with $T=2\pi\sqrt{a^{3}\mu/\alpha}$, yields the exact secular fluxes
$\langle\dot{E}\rangle=\frac{\alpha\mathcal{K}\,(e^{2}+2)}{2a^{4}(1-e^{2})^{5/2}},\qquad\langle\dot{L}\rangle=\frac{\mathcal{K}\,\sqrt{\alpha\mu}}{a^{5/2}(1-e^{2})}.$ (44)
Using $E=-\alpha/(2a)$ and differentiating $L^{2}=\mu\alpha a(1-e^{2})$ then gives the evolution equations for dipole-driven electromagnetic inspiral,
$\langle\dot{a}\rangle=\frac{(e^{2}+2)\mathcal{K}}{a^{2}(1-e^{2})^{5/2}},\qquad\langle\dot{e}\rangle=\frac{3e\,\mathcal{K}}{2a^{3}(1-e^{2})^{3/2}}.$ (45)
In particular, $e=0$ is a marginal fixed point ($\langle\dot{e}\rangle\propto e$), and for inspiral ($\mathcal{K}<0$) the system both shrinks and circularizes.
Taking the ratio of (45) gives an exactly integrable semi-major-axis–eccentricity relation,
$\frac{de}{da}=\frac{3e(1-e^{2})}{2a(e^{2}+2)}\quad\Longrightarrow\quad a(e)=a_{0}\frac{e^{4/3}}{1-e^{2}}\,\frac{1-e_{0}^{2}}{e_{0}^{4/3}},$ (46)
which is notably simpler than the quadrupole-driven gravitational analogue. Substituting $a(e)$ into $\langle\dot{e}\rangle$ yields an elementary time–eccentricity integral and the exact closed form
$t(e)-t_{0}=\frac{2}{3\mathcal{K}}a_{0}^{3}\frac{(1-e_{0}^{2})^{3}}{e_{0}^{4}}\left[\frac{2-e^{2}}{\sqrt{1-e^{2}}}-\frac{2-e_{0}^{2}}{\sqrt{1-e_{0}^{2}}}\right],$ (47)
from which the merger time (defined by $e\to 0$ and hence $a\to 0$) follows as
$t_{\rm c}-t_{0}=\frac{2}{3\mathcal{K}}a_{0}^{3}\frac{(1-e_{0}^{2})^{3}}{e_{0}^{4}}\left[2-\frac{2-e_{0}^{2}}{\sqrt{1-e_{0}^{2}}}\right].$ (48)
For inspiral ($\mathcal{K}<0$), the bracket is negative so that $t_{\rm c}-t_{0}>0$. In the circular limit $e_{0}\to 0$ one finds $t_{\rm c}-t_{0}=a_{0}^{3}/(6|\mathcal{K}|)+O(e_{0}^{2})$, while as $e_{0}\to 1^{-}$ the merger time scales as $t_{\rm c}-t_{0}\propto(1-e_{0}^{2})^{5/2}$, showing that highly eccentric electromagnetic binaries merge parametrically faster. Relative to quadrupole-driven gravitational radiation, the dipole system has the characteristic weaker $a$-scaling in (45) and the distinctive near-circular invariant $a\propto e^{4/3}$ (cf. gravity case’ $a\propto e^{12/19}$), directly reflecting the lower multipolar order of the radiation channel. To validate the analytic eccentric inspiral in the N body framework Eq. (16), we performed direct numerical integration of the full two-body phase-space equations including exact Coulomb interactions, conservative Darwin corrections at $\mathcal{O}(c^{-2})$, and leading dipole radiation reaction at $\mathcal{O}(c^{-3})$. Starting from bound Keplerian ellipses with moderate eccentricity, the motion remains strictly planar and exhibits monotonic decrease of the Darwin Hamiltonian, confirming consistent radiative energy loss. Secular orbital elements extracted from successive turning points show $\dot{a}<0$ and $\dot{e}<0$ for $\mathcal{K}<0$, demonstrating simultaneous shrinkage and circularization in agreement with Eq. (45). Furthermore, the predicted Kepler-level invariant $\mathcal{I}=a(1-e^{2})/e^{4/3}$ is preserved to sub-percent accuracy, with relative drift $\Delta\mathcal{I}/\mathcal{I}_{0}\lesssim 4\times 10^{-3}$ over $\sim 10^{2}$ orbits (The small secular drift is consistent with residual $\mathcal{O}(1/c^{2})$ conservative corrections, which should show up in the simulations by the choice $c=20$.). The small systematic deviation is consistent with the inclusion of $\mathcal{O}(c^{-2})$ conservative Darwin corrections, which perturb the pure Kepler mapping between $(E,L)$ and $(a,e)$. The inspiral timescale also follows the analytic dipole scaling, $t_{c}-t_{0}\propto a_{0}^{3}$ in the near-circular limit and $\propto(1-e_{0}^{2})^{5/2}$ for highly eccentric initial data. Overall, the simulations provide good support for the eccentric dipole driven inspiral and the phase-space radiation-reaction implementation described by Eq. (16).
VII Eccentric inspiral with 11PN conservative corrections and leading dipole radiation
We derive the secular inspiral of an eccentric electromagnetic binary including conservative dynamics through 11PN ($\mathcal{O}(c^{-2})$) while retaining the leading electric–dipole radiation channel (absolute $\mathcal{O}(c^{-3})$). The resulting evolution equations for $(a,e)$ are expanded consistently through absolute $\mathcal{O}(c^{-5})$ (i.e. Newtonian dipole flux plus 11PN corrections entering via the conservative mapping between $(E,L)$ and $(a,e)$).
The relative Hamiltonian in the center-of-mass frame is
$H=\frac{p_{r}^{2}}{2\mu}+\frac{L^{2}}{2\mu r^{2}}+\frac{\alpha}{r}+\frac{1}{c^{2}}\left[-\frac{1-3\nu}{8\mu^{3}}\left(p_{r}^{2}+\frac{L^{2}}{r^{2}}\right)^{2}+\frac{\alpha}{2\mu Mr}\left(2p_{r}^{2}+\frac{L^{2}}{r^{2}}\right)\right],$ (49)
with $M=m_{1}+m_{2}$, $\mu=m_{1}m_{2}/M$, $\nu=\mu/M$, and $\alpha=q_{1}q_{2}/(4\pi\varepsilon_{0})$. For bound motion, imposing the turning-point conditions at $r_{p}=a(1-e)$ and $r_{a}=a(1+e)$ yields the invariants
$E(a,e)$ $=-\frac{\alpha}{2a}-\frac{\alpha^{2}(1-\nu)}{8\mu a^{2}c^{2}}+\mathcal{O}(c^{-4}),$
$L^{2}(a,e)$ $=\mu\alpha a(1-e^{2})+\frac{\alpha^{2}}{2c^{2}}\left(1+e^{2}(1-2\nu)\right)+\mathcal{O}(c^{-4}).$ (50)
Dissipation is driven by the leading dipole radiation reaction, giving the orbit-averaged fluxes (Coulombic averaging)
$\langle\dot{E}\rangle=\frac{\alpha\mathcal{K}(e^{2}+2)}{2a^{4}(1-e^{2})^{5/2}},\qquad\langle\dot{L}\rangle=\frac{\mathcal{K}\sqrt{\alpha\mu}}{a^{5/2}(1-e^{2})}.$ (51)
Since $E=E(a,e)$ and $L=L(a,e)$, the secular evolution follows from $(\dot{E},\dot{L})^{T}=J(a,e)\,(\dot{a},\dot{e})^{T}$ with $J_{ij}=\partial_{(a,e)}(E,L)$. Inverting $J$ consistently through $\mathcal{O}(c^{-2})$ yields the 11PN-corrected dipole-driven inspiral laws
$\dot{a}$ $=\frac{\mathcal{K}(e^{2}+2)}{a^{2}(1-e^{2})^{5/2}}\left[1+\frac{\alpha(1-\nu)}{2\mu ac^{2}}\right]+\mathcal{O}(c^{-5}),$
$\dot{e}$ $=\frac{3e\mathcal{K}}{2a^{3}(1-e^{2})^{3/2}}\left[1+\frac{\alpha(1-2\nu)}{2\mu ac^{2}}\right]+\mathcal{O}(c^{-5}).$ (52)
Equation (52) reduces to the dipole (Coulombic) system in the $c\to\infty$ limit, with $e=0$ a marginal fixed point and (inspiral) $\mathcal{K}<0$ implying both $\dot{a}<0$ and $\dot{e}<0$ for $e>0$.
For later use it is convenient to introduce the 11PN-corrected mean motion
$n^{2}=\frac{\alpha}{\mu a^{3}}\left[1+\frac{\alpha}{2\mu ac^{2}}(1-2\nu)\right],$ (53)
so that $\dot{n}=-(3/2)n\,\dot{a}/a$ gives an eccentric dipole “chirp” of the form
$\dot{n}=A\,n^{3}\,\frac{e^{2}+2}{(1-e^{2})^{5/2}}\left[1+x\,\Gamma(e,\nu)\right]+\mathcal{O}(c^{-5}),$ (54)
with
$A=\frac{3\mathcal{K}\mu}{-\alpha},\qquad x=\frac{1}{c^{2}}\left(\frac{-\alpha}{\mu}\right)^{2/3}n^{2/3},$ (55)
and $\Gamma(e,\nu)$ an analytic function determined by the 11PN terms in (52) together with (53). The orbital phase follows from $\dot{\Phi}=n$, i.e. $d\Phi/dn=n/\dot{n}$, so that the leading dipole structure implies $\Phi\sim\tau^{1/2}$ (with $\tau=t_{c}-t$) rather than the quadrupole gravitational scaling $\Phi\sim\tau^{5/8}$.
Finally, the electromagnetic dipole hierarchy is systematically “multipole-lowered” relative to quadrupole-driven gravitational inspiral: at leading order $\dot{n}\propto n^{3}$ with an eccentric enhancement $(1-e^{2})^{-5/2}$, whereas in GR $\dot{n}_{\rm GR}\propto n^{11/3}$ with $(1-e^{2})^{-7/2}$, reflecting the higher multipolar order of the radiation channel.
VIII Einstein–Maxwell binaries: ADM-type CoM Hamiltonian through 2PN and leading 1.5PN radiation reaction
Compact objects are generally expected to be electrically neutral in astrophysical environments, as standard electromagnetic processes tend to rapidly neutralize any macroscopic charge imbalance. Nevertheless, a small residual charge cannot be excluded a priori, particularly in scenarios involving additional hidden $U(1)$ sectors or exotic charge assignments beyond standard electromagnetism. Such charges, if sufficiently long-lived, would modify both the conservative two-body dynamics and the radiation channel during inspiral, leading to potentially observable corrections to the gravitational-wave phase evolution. Recent work has established the conservative dynamics of charged black-hole binaries in Einstein–Maxwell theory through 22PN order, including gauge-invariant quantities such as the binding energy, periastron advance, and scattering angle [36]. Building on these results, and incorporating the leading 1.51.5PN dipolar radiation-reaction sector discussed in Sec. IV, we now extend the inspiral analysis to relativistic charged compact binaries within a canonical framework.
From an observational standpoint, electric charge modifies the inspiral through two distinct mechanisms. First, it alters the conservative two-body dynamics via an effective coupling, thereby shifting the binding energy and the energy–frequency relation at 11PN and 22PN order. Second, it introduces dipolar radiation at absolute 1.51.5PN order, generating a characteristic $\dot{\Omega}\propto\Omega^{3}$ scaling (Eq. (31)) that competes with the standard gravitational quadrupole flux $\dot{\Omega}\propto\Omega^{11/3}$. The relative importance of these channels determines whether the binary inspiral is charge dipole- or gravitational quadrupole-dominated and thus impacts inspiral dynamics.
In the following we summarize the ADM-type center-of-mass Hamiltonian through 22PN order and incorporate the leading 1.51.5PN dipole radiation-reaction force, providing the explicit phase-space structure required for gauge-invariant inspiral laws and for assessing the detectability of charge-induced deviations in GW observations. After performing the harmonic-to-ADM contact transformation and reducing to the center-of-mass (CoM) frame, the two-body dynamics can be written in terms of the dimensionless variables
$P^{2}\equiv\frac{p^{2}}{\mu^{2}},\qquad P_{R}\equiv\frac{p_{r}}{\mu},\qquad R\equiv\frac{r}{M},$ (56)
where $M=m_{1}+m_{2}$, $\mu=m_{1}m_{2}/M$, $\nu=\mu/M$, and $X_{12}=(m_{1}-m_{2})/M$. We also introduce the charge-to-mass ratios
$\eta_{A}\equiv\frac{q_{A}}{\sqrt{G}\,m_{A}},\qquad s\equiv 1-\eta_{1}\eta_{2},$ (57)
so that the Newtonian interaction is controlled by an effective coupling $G_{\rm eff}=Gs$. In these variables the ADM-type CoM Hamiltonian admits the post-Newtonian expansion
$\frac{H_{\rm CoM}^{\rm ADM-type}}{\mu}=-\frac{c^{2}}{\nu}+H_{\rm N}+\frac{1}{c^{2}}H_{1{\rm PN}}+\frac{1}{c^{4}}H_{2{\rm PN}},$ (58)
with Newtonian, 1PN, and 2PN contributions
$H_{\rm N}=\frac{P^{2}}{2}-\frac{G}{R}\,s,$ (59)
$H_{1{\rm PN}}=$ $\;\frac{P^{4}}{8}(3\nu-1)-\frac{GM}{2R}\Big[3P^{2}+\nu s\,(P^{2}+P_{R}^{2})\Big]$
$\;+\frac{G^{2}}{4R^{2}}\Big[2+\eta_{1}^{2}-4\eta_{1}\eta_{2}+\eta_{2}^{2}+X_{12}(\eta_{1}^{2}-\eta_{2}^{2})\Big],$ (60)
and
$H_{2{\rm PN}}=$ $\;\frac{P^{6}}{16}(1-5\nu+5\nu^{2})$
$+\frac{G^{3}}{4R^{3}}\Bigg[\eta_{1}\eta_{2}-\eta_{1}^{2}-\eta_{2}^{2}-2+X_{12}(\eta_{2}^{2}-\eta_{1}^{2})+\nu\Big(23\eta_{1}\eta_{2}-2\eta_{1}^{2}-2\eta_{2}^{2}-3\eta_{1}^{2}\eta_{2}^{2}-\eta_{1}^{3}\eta_{2}^{3}-15\Big)\Bigg]$
$+\frac{G^{2}}{8R^{2}}\Bigg[X_{12}(\eta_{2}^{2}-\eta_{1}^{2})(P^{2}-2P_{R}^{2})(1-\nu)+2P_{R}^{2}(\eta_{1}^{2}+\eta_{2}^{2}-2)+P^{2}(\eta_{1}^{2}+\eta_{2}^{2}+22)$
$\hskip 73.97733pt+\nu\Big(P^{2}(58+7\eta_{1}^{2}-46\eta_{1}\eta_{2}+7\eta_{2}^{2}+2\eta_{1}^{2}\eta_{2}^{2})-P_{R}^{2}(32+10\eta_{1}^{2}+10\eta_{2}^{2}+4\eta_{1}\eta_{2})\Big)\Bigg]$
$+\frac{G}{8R}\Bigg[5P^{4}+2\nu P^{2}\Big(s\,P_{R}^{2}-P^{2}(11+\eta_{1}\eta_{2})\Big)-\nu^{2}s\,(3P^{4}+2P^{2}P_{R}^{2}+3P_{R}^{4})\Bigg]$
$+\frac{G}{4R^{3}}s\,(1-12\nu)+\frac{P_{R}^{2}}{2R^{2}}-\frac{P^{2}}{4R^{2}}+\frac{3\nu}{R^{2}}(2P_{R}^{2}-P^{2})-\frac{G\nu}{4R^{2}}(2P^{2}+P_{R}^{2})s+\frac{\nu P^{2}}{4R}(P^{2}-P_{R}^{2}).$ (61)
The Hamiltonian is defined up to canonical transformations. The residual freedom is encoded in three combinations $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$:
$H_{2{\rm PN}}^{\rm contact}=$ $\mathbf{A}\Bigg[\frac{P^{2}}{MR}(P_{R}^{2}-P^{2})+\frac{G}{MR^{2}}(1-\eta_{1}\eta_{2})(2P_{R}^{2}+P^{2})\Bigg]$
$+\mathbf{B}\Bigg[\frac{2P_{R}^{2}-P^{2}}{M^{2}R^{2}}+\frac{G}{M^{2}R^{3}}(1-\eta_{1}\eta_{2})\Bigg]$
$+\mathbf{C}\Bigg[\frac{3GP_{R}^{2}}{MR^{2}}(1-\eta_{1}\eta_{2})+\frac{3P_{R}^{2}}{MR}(P_{R}^{2}-P^{2})\Bigg].$ (62)
The coefficients are
$\mathbf{A}$ $=\frac{1}{M^{3}}\Big(m_{2}^{3}A_{1}-m_{1}m_{2}^{2}(A_{2}+A_{3})+m_{1}^{2}m_{2}(A_{4}+A_{5})-m_{1}^{3}A_{6}\Big),$ (63)
$\mathbf{B}$ $=\frac{1}{M}(-m_{1}B_{1}+m_{2}B_{2}),$ (64)
$\mathbf{C}$ $=\frac{1}{M^{3}}\Big(m_{2}^{3}C_{1}-m_{1}m_{2}^{2}C_{2}+m_{1}^{2}m_{2}C_{3}-m_{1}^{3}C_{4}\Big).$ (65)
For circular orbits we set $P_{R}=0$ and write
$P^{2}=\frac{p^{2}}{\mu^{2}}=\frac{L^{2}}{\mu^{2}r^{2}}=\frac{j^{2}}{R^{2}},\qquad j\equiv\frac{L}{\mu M}.$ (66)
Defining the (dimensionless) binding energy per unit $\mu$ by
$E_{b}\equiv\frac{H-Mc^{2}}{\mu}=H_{\rm N}+\frac{1}{c^{2}}H_{1{\rm PN}}+\frac{1}{c^{4}}H_{2{\rm PN}},$ (67)
we impose the circularity condition at fixed $j$, $\left.\partial_{R}E_{b}(R,j)\right|_{j}=0$, solve perturbatively for $j=j(R)$ through 2PN order, and compute the orbital frequency from Hamilton’s equation,
$\Omega=\dot{\phi}=\frac{\partial H}{\partial L}\quad\Longrightarrow\quad M\Omega=\frac{\partial E_{b}(R,j)}{\partial j}\Big|_{j=j(R)}.$ (68)
Introducing the gauge-invariant frequency parameter
$x_{q}\equiv\left(\frac{GM\Omega}{c^{3}}\right)^{2/3}s^{2/3},$ (69)
and eliminating $R$ in favor of $x_{q}$, the circular binding energy admits the 2PN expansion
$E_{b}(x_{q})=-\frac{x_{q}}{2}+\frac{x_{q}^{2}}{24}\,s^{-2}\,\mathcal{E}_{1{\rm PN}}+\frac{x_{q}^{3}}{48}\,s^{-4}\,\mathcal{E}_{2{\rm PN}}+\mathcal{O}(x_{q}^{4}),$ (70)
with
$\mathcal{E}_{1{\rm PN}}=$  $\,9+\nu+2(2+X_{12})\eta_{1}^{2}-2(3+\nu)\eta_{1}\eta_{2}-2(2-X_{12})\eta_{2}^{2}+2(1+\nu)\eta_{1}^{2}\eta_{2}^{2},$ (71)
and
$\mathcal{E}_{2{\rm PN}}=$  $\;81-57\nu+\nu^{2}-2(81-84\nu+2\nu^{2})\eta_{1}\eta_{2}-4(1-2\nu+X_{12})\eta_{1}^{4}-4(1+X_{12}-2\nu)\eta_{2}^{4}$
$+4\eta_{1}^{3}\eta_{2}\,[9+13\nu+X_{12}(9+\nu)]+4\eta_{1}\eta_{2}^{3}\,[15-11\nu-X_{12}(15+\nu)]+(90-254\nu+6\nu^{2})\eta_{1}^{2}\eta_{2}^{2}$
$-\eta_{1}^{3}\eta_{2}^{3}(9-76\nu+2\nu^{2})+2\eta_{1}^{4}\eta_{2}^{2}(2\nu-19+9X_{12})+2\eta_{1}^{2}\eta_{2}^{4}(2\nu-19-9X_{12})+(1-25\nu+\nu^{2})\eta_{1}^{4}\eta_{2}^{4}.$ (72)
By construction, $E_{b}(x_{q})$ is invariant under canonical/contact transformations and therefore independent of the specific ADM-type gauge coefficients entering the Hamiltonian. Finally, the leading dissipative (1.5PN) radiation-reaction acceleration in Einstein–Maxwell theory can be written in SI units by restoring the electromagnetic coupling $k\equiv(4\pi\varepsilon_{0})^{-1}$ (equivalently $q_{\rm geom}=q_{\rm SI}/\sqrt{4\pi\varepsilon_{0}}$). The relative dissipative acceleration $\mathbf{a}_{\rm rel}^{\rm RR}\equiv\mathbf{a}_{1}^{\rm RR}-\mathbf{a}_{2}^{\rm RR}$ becomes
$\mathbf{a}_{\rm rel}^{\rm RR}=-\frac{2}{3c^{3}r^{3}}\,k\,\frac{(m_{1}q_{2}-m_{2}q_{1})^{2}}{m_{1}^{2}m_{2}^{2}}\left(Gm_{1}m_{2}-k\,q_{1}q_{2}\right)\mathbf{T},$ (73)
with the universal tensor structure
$\mathbf{T}=\mathbf{v}-3(\hat{\mathbf{r}}\!\cdot\!\mathbf{v})\,\hat{\mathbf{r}},\qquad\hat{\mathbf{r}}\equiv\frac{\mathbf{r}}{r},$ (74)
where $\mathbf{r}$ and $\mathbf{v}$ are the relative separation and velocity.
IX Dipole–Quadrupole Crossover
The dipole–quadrupole transition is defined by equality of the orbit-averaged luminosities,
$\mathcal{F}_{\rm dip}=\mathcal{F}_{\rm quad}.$ (75)
For circular orbits written in terms of the gauge-invariant parameter
$x_{q}\equiv\left(\frac{GM\Omega}{c^{3}}\right)^{2/3}s^{2/3},\qquad s\equiv 1-\eta_{1}\eta_{2},$ (76)
the fluxes read
$\mathcal{F}_{\rm dip}$ $=\mu\,\frac{2c^{5}k_{E}\Delta^{2}}{3G^{2}M^{3}m_{1}m_{2}}\,s^{-2}\,x_{q}^{4},$ (77)
$\mathcal{F}_{\rm quad}$ $=\frac{32}{5}\frac{c^{5}}{G}\,\nu^{2}\,x_{q}^{5}\,s^{-2},$ (78)
with $\nu=\mu/M$ and $\Delta=m_{1}m_{2}(\zeta_{2}-\zeta_{1})$, $\zeta_{A}=q_{A}/m_{A}$.
Equating the fluxes and solving for $x_{q}$ yields the gauge-invariant crossover value
$x_{q,\rm cross}=\frac{5}{48}(\eta_{2}-\eta_{1})^{2},$ (79)
where $\eta_{A}=\sqrt{k_{E}/G}\,\zeta_{A}$. All factors of $s$ cancel identically, so the crossover depends only on the charge-to-mass asymmetry.
Using
$\Omega=\frac{c^{3}}{GM}\,\frac{x_{q}^{3/2}}{s},$ (80)
the corresponding orbital frequency is
$\Omega_{\rm cross}=\frac{c^{3}}{GM}\frac{1}{s}\left(\frac{5}{48}\right)^{3/2}|\eta_{2}-\eta_{1}|^{3}.$ (81)
The dominant $m=2$ gravitational-wave frequency is therefore
$f_{\rm cross}=\frac{1}{\pi}\frac{c^{3}}{GM}\frac{1}{s}\left(\frac{5}{48}\right)^{3/2}|\eta_{2}-\eta_{1}|^{3}.$ (82)
Dipole radiation dominates for $x_{q}<x_{q,\rm cross}$, while quadrupole emission dominates for $x_{q}>x_{q,\rm cross}$. The crossover vanishes for equal charge-to-mass ratios ($\eta_{1}=\eta_{2}$), reflecting exact cancellation of the electric dipole moment, and in the neutral limit the quadrupole channel dominates at all frequencies. The dimensionless parameter
$x_{q,\rm cross}=\frac{5}{48}\,(\eta_{2}-\eta_{1})^{2}$ (83)
defines the frequency scale at which the leading dissipative driver of the inspiral transitions from electromagnetic dipole emission to gravitational quadrupole emission.
For $x_{q}\ll x_{q,\rm cross}$, the inspiral is dipole-dominated and the orbital evolution obeys the characteristic scaling
$\dot{\Omega}\sim\Omega^{3},$ (84)
whereas for $x_{q}\gg x_{q,\rm cross}$ the quadrupole channel dominates and the evolution asymptotes to the standard GR scaling
$\dot{\Omega}\sim\Omega^{11/3}.$ (85)
Numerically, this can be written as
$f_{\rm cross}\simeq 2.2\times 10^{3}\,\mathrm{Hz}\left(\frac{M_{\odot}}{M}\right)\frac{|\eta_{2}-\eta_{1}|^{3}}{1-\eta_{1}\eta_{2}}.$ (86)
For a $60\,M_{\odot}$ binary this becomes
$f_{\rm cross}\simeq 36\,\mathrm{Hz}\,\frac{|\eta_{2}-\eta_{1}|^{3}}{1-\eta_{1}\eta_{2}}.$ (87)
The cubic dependence on the charge-to-mass asymmetry implies that for $|\eta_{A}|\ll 1$ the crossover lies far below the ground-based detector band, so the inspiral remains quadrupole dominated; dipole radiation effects enter the LIGO/Virgo band only if the charge-to-mass asymmetry $|\eta_{2}-\eta_{1}|$ is of order unity, i.e. in highly asymmetric and near-extremal configurations.
Finally, this estimate is restricted to the leading-order dissipative PN truncation used here and to the regime $x_{q}\ll 1$ where the PN expansion is reliable.
X Dipole-driven circular inspiral through 2PN conservative order
We consider a quasi-circular charged binary in Einstein–Maxwell theory in the dipole-dominated regime, where electromagnetic dipole losses dominate over gravitational quadrupole emission. The conservative dynamics is retained through $2{\rm PN}$ order, encoded in the gauge-invariant circular binding energy $E_{b}(x_{q})$ derived from the ADM-type CoM Hamiltonian, while dissipation is included at leading $1.5{\rm PN}$ order via the radiation–reaction force.
We define the gauge-invariant frequency parameter
$x_{q}\equiv\left(\frac{GM\Omega}{c^{3}}\right)^{2/3}s^{2/3},\qquad s\equiv 1-\eta_{1}\eta_{2},\qquad\eta_{A}\equiv\frac{\sqrt{k_{E}}\,q_{A}}{\sqrt{G}\,m_{A}},\qquad k_{E}\equiv\frac{1}{4\pi\varepsilon_{0}},$ (88)
and the charge-asymmetry combination
$\Delta\equiv m_{1}q_{2}-m_{2}q_{1},\qquad M\equiv m_{1}+m_{2},\qquad\mu\equiv\frac{m_{1}m_{2}}{M},\qquad\nu\equiv\frac{\mu}{M}.$ (89)
For circular motion, the binding energy per unit reduced mass admits the $2{\rm PN}$ expansion
$\frac{E_{b}(x_{q})}{c^{2}}=-\frac{x_{q}}{2}+\frac{x_{q}^{2}}{24}\,s^{-2}\mathcal{E}_{1{\rm PN}}+\frac{x_{q}^{3}}{48}\,s^{-4}\mathcal{E}_{2{\rm PN}}+\mathcal{O}(x_{q}^{4}),$ (90)
where the explicit coefficients $\mathcal{E}_{1{\rm PN}}$ and $\mathcal{E}_{2{\rm PN}}$ are given in Sec. VIII. Differentiating,
$c^{-2}\frac{dE_{b}}{dx_{q}}=-\frac{1}{2}+\frac{x_{q}}{12}\,s^{-2}\mathcal{E}_{1{\rm PN}}+\frac{x_{q}^{2}}{16}\,s^{-4}\mathcal{E}_{2{\rm PN}}.$ (91)
Dissipation is driven by electromagnetic dipole radiation. For circular orbits, $(\hat{\mathbf{r}}\!\cdot\!\mathbf{v})=0$ so the relative radiation–reaction acceleration is antiparallel to the relative velocity, and the mechanical energy loss satisfies $\dot{E}=\mu\,\mathbf{v}\!\cdot\!\mathbf{a}_{\rm rel}^{\rm RR}$. Using the circular Kepler law $\Omega^{2}=GMs/r^{3}$ and eliminating $r$ in favor of $x_{q}$, the orbit-averaged dipole luminosity can be written as
$\mathcal{F}_{\rm dip}=-\dot{E}=\mu\,\frac{2c^{5}k_{E}\Delta^{2}}{3G^{2}M^{3}m_{1}m_{2}}\,s^{-2}\,x_{q}^{4}.$ (92)
In the dipole-dominated regime, energy balance reduces to
$\mu\,\frac{dE_{b}}{dx_{q}}\,\dot{x}_{q}=-\mathcal{F}_{\rm dip}.$ (93)
Solving for $\dot{x}_{q}$ and converting with $\dot{\Omega}/\Omega=(3/2)\dot{x}_{q}/x_{q}$, we obtain the closed-form inspiral law after a consistent expansion through relative $\mathcal{O}(x_{q}^{2})$ gives
$\frac{\dot{\Omega}}{\Omega}=\frac{2c^{3}k_{E}\Delta^{2}}{G^{2}M^{3}m_{1}m_{2}}\,s^{-2}\,x_{q}^{3}\left[1+\frac{x_{q}}{6}s^{-2}\mathcal{E}_{1{\rm PN}}+x_{q}^{2}s^{-4}\left(\frac{\mathcal{E}_{2{\rm PN}}}{8}+\frac{\mathcal{E}_{1{\rm PN}}^{2}}{36}\right)\right]+\mathcal{O}(x_{q}^{6}).$ (94)
Since $x_{q}=\left(GMs\,\Omega/c^{3}\right)^{2/3}$, the leading term implies the characteristic dipole-driven scaling $\dot{\Omega}=\Omega\,(\dot{\Omega}/\Omega)$ with $\dot{\Omega}$ growing as the cube of $\Omega$, in contrast with the quadrupole-driven GR scaling.
It is convenient to introduce the constant
$\Gamma\equiv\frac{2c^{3}k_{E}\Delta^{2}}{G^{2}M^{3}m_{1}m_{2}\,s^{2}},$ (95)
and rewrite Eq. (94) as
$\frac{\dot{\Omega}}{\Omega}=\Gamma\,x_{q}^{3}\left(1+a_{1}x_{q}+a_{2}x_{q}^{2}\right),\qquad a_{1}=\frac{\mathcal{E}_{1{\rm PN}}}{6s^{2}},\qquad a_{2}=\frac{\mathcal{E}_{1{\rm PN}}^{2}}{36s^{4}}+\frac{\mathcal{E}_{2{\rm PN}}}{8s^{4}}.$ (96)
Using $\dot{\Omega}/\Omega=(3/2)\dot{x}_{q}/x_{q}$, we obtain the equivalent chirp equation
$\dot{x}_{q}=\frac{2}{3}\Gamma x_{q}^{4}\left(1+a_{1}x_{q}+a_{2}x_{q}^{2}\right),$ (97)
and its perturbative inverse
$\frac{dt}{dx_{q}}=\frac{3}{2\Gamma}\,x_{q}^{-4}\left[1-a_{1}x_{q}+(a_{1}^{2}-a_{2})x_{q}^{2}\right]+\mathcal{O}(x_{q}^{-1}).$ (98)
Defining $\tau\equiv t_{c}-t$ and fixing the integration constant by $\tau\to 0$ as $x_{q}\to\infty$, one finds the time-to-coalescence relation
$\tau(x_{q})=\frac{1}{2\Gamma}\left[x_{q}^{-3}+\frac{3a_{1}}{2}x_{q}^{-2}+3(a_{2}-a_{1}^{2})x_{q}^{-1}\right],$ (99)
valid through relative $\mathcal{O}(x_{q}^{2})$. Inverting perturbatively yields an explicit chirp law,
$x_{q}(\tau)=(2\Gamma\tau)^{-1/3}\left[1+\frac{a_{1}}{2}(2\Gamma\tau)^{-1/3}+(a_{2}-a_{1}^{2})(2\Gamma\tau)^{-2/3}\right]+\mathcal{O}\!\left((2\Gamma\tau)^{-1}\right).$ (100)
Using $a_{2}-a_{1}^{2}=\mathcal{E}_{2{\rm PN}}/(8s^{4})$, the second correction depends only on the genuine $2{\rm PN}$ conservative coefficient.
Finally, the orbital frequency follows from the definition of $x_{q}$,
$\Omega(\tau)=\frac{c^{3}}{GMs}\,x_{q}^{3/2}(\tau),$ (101)
which gives
$\Omega(\tau)=\frac{c^{3}}{GMs}(2\Gamma\tau)^{-1/2}\left[1+\frac{\mathcal{E}_{1{\rm PN}}}{8s^{2}}(2\Gamma\tau)^{-1/3}+\frac{1}{s^{4}}\left(\frac{3\mathcal{E}_{2{\rm PN}}}{16}+\frac{\mathcal{E}_{1{\rm PN}}^{2}}{384}\right)(2\Gamma\tau)^{-2/3}\right]+\mathcal{O}\!\left((2\Gamma\tau)^{-1}\right).$ (102)
At leading order this implies $\Omega(\tau)=(c^{3}/GMs)(2\Gamma\tau)^{-1/2}$, so the dipole-driven inspiral exhibits $\Omega$ diverging as $(t_{c}-t)^{-1/2}$ and therefore $\dot{\Omega}$ scaling with the cube of $\Omega$.
XI Full circular inspiral at 2PN conservative order with dipole and quadrupole radiation
We consider a quasi-circular binary of masses $m_{1},m_{2}$ and charges $q_{1},q_{2}$, including electromagnetic dipole losses and gravitational quadrupole losses, while retaining conservative dynamics through 22PN order. We introduce the gauge-invariant frequency parameter
$x_{q}=\left(\frac{GM\Omega}{c^{3}}\right)^{2/3}s^{2/3},\qquad s\equiv 1-\eta_{1}\eta_{2},\qquad\eta_{A}=\sqrt{\frac{k_{E}}{G}}\frac{q_{A}}{m_{A}}.$ (103)
The conservative binding energy through 22PN order reads
$\frac{E_{b}(x_{q})}{c^{2}}=-\frac{x_{q}}{2}+\frac{x_{q}^{2}}{24}s^{-2}\mathcal{E}_{1\mathrm{PN}}+\frac{x_{q}^{3}}{48}s^{-4}\mathcal{E}_{2\mathrm{PN}}+\mathcal{O}(x_{q}^{4}),$ (104)
so that
$\mu\frac{dE_{b}}{dx_{q}}=-\frac{\mu c^{2}}{2}\left[1-ax_{q}-bx_{q}^{2}+\mathcal{O}(x_{q}^{3})\right],$ (105)
with
$a=\frac{1}{6}s^{-2}\mathcal{E}_{1\mathrm{PN}},\qquad b=\frac{1}{8}s^{-4}\mathcal{E}_{2\mathrm{PN}}.$ (106)
Energy balance for adiabatic inspiral,
$\mu\frac{dE_{b}}{dx_{q}}\dot{x}_{q}=-(\mathcal{F}_{\rm dip}+\mathcal{F}_{\rm quad}),$ (107)
requires the dipole and quadrupole luminosities expressed in $x_{q}$. The leading electromagnetic dipole flux is
$\mathcal{F}_{\rm dip}=\mu\frac{2c^{5}k_{E}\Delta^{2}}{3G^{2}M^{3}m_{1}m_{2}}\,s^{-2}x_{q}^{4},\qquad\Delta=m_{1}q_{2}-m_{2}q_{1},$ (108)
while the gravitational quadrupole flux, after eliminating $r$ using the charged Kepler law $\Omega^{2}=GMs/r^{3}$, becomes
$\mathcal{F}_{\rm quad}=\frac{32}{5}\frac{c^{5}}{G}\nu^{2}\,s^{-2}x_{q}^{5},\qquad\nu=\frac{\mu}{M}.$ (109)
Solving the balance equation yields the full inspiral law
$\dot{x}_{q}=\frac{2}{3}x_{q}^{4}(\Gamma_{\rm d}+\Gamma_{\rm q}x_{q})\left[1+ax_{q}+(a^{2}+b)x_{q}^{2}\right]+\mathcal{O}(x_{q}^{7}),$ (110)
where
$\Gamma_{\rm d}=\frac{2c^{3}k_{E}\Delta^{2}}{G^{2}M^{3}m_{1}m_{2}}s^{-2},\qquad\Gamma_{\rm q}=\frac{96}{5}\frac{c^{3}}{G}\frac{\mu}{M^{2}}s^{-2}.$ (111)
Using $\dot{\Omega}/\Omega=\tfrac{3}{2}\dot{x}_{q}/x_{q}$ gives
$\frac{\dot{\Omega}}{\Omega}=Ax_{q}^{3}+(B+Aa)x_{q}^{4}+\mathcal{O}(x_{q}^{5}),$ (112)
with
$A=\frac{3}{2}\frac{4c^{3}k_{E}\Delta^{2}}{3G^{2}M^{3}m_{1}m_{2}}s^{-2},\qquad B=\frac{3}{2}\frac{64}{5G}c^{3}\frac{\mu}{M^{2}}s^{-2}.$ (113)
The PN hierarchy is manifest: the $x_{q}^{3}$ term represents 1.51.5PN dipole radiation, the $x_{q}^{4}$ term contains both 2.52.5PN quadrupole radiation and the 11PN conservative correction to dipole emission, and higher powers correspond to mixed 22PN effects.
The dipole–quadrupole crossover occurs when $\mathcal{F}_{\rm dip}=\mathcal{F}_{\rm quad}$, equivalently when $\Gamma_{\rm d}=\Gamma_{\rm q}x_{q}$, yielding the gauge-invariant value
$x_{q,\rm cross}=\frac{\Gamma_{\rm d}}{\Gamma_{\rm q}}=\frac{5}{48}(\eta_{2}-\eta_{1})^{2}.$ (114)
Thus dipole radiation dominates for $x_{q}\ll x_{q,\rm cross}$, while quadrupole emission dominates for $x_{q}\gg x_{q,\rm cross}$. The crossover depends only on the charge-to-mass asymmetry and is independent of $s$.
The time-to-coalescence $\tau=t_{c}-t$ follows from
$\frac{dt}{dx_{q}}=\frac{3}{2}\frac{1-ax_{q}-bx_{q}^{2}}{x_{q}^{4}(\Gamma_{\rm d}+\Gamma_{\rm q}x_{q})},$ (115)
so that
$\tau(x_{q})=\frac{3}{2}\int_{x_{q}}^{\infty}\frac{1-ax^{\prime}-bx^{\prime 2}}{x^{\prime 4}(\Gamma_{\rm d}+\Gamma_{\rm q}x^{\prime})}\,dx^{\prime}.$ (116)
In the dipole-dominated regime ($x_{q}\ll x_{q,\rm cross}$),
$\tau\simeq\frac{1}{2\Gamma_{\rm d}}x_{q}^{-3},\qquad\dot{\Omega}\sim\Omega^{3},$ (117)
whereas in the quadrupole-dominated regime ($x_{q}\gg x_{q,\rm cross}$),
$\tau\simeq\frac{3}{8\Gamma_{\rm q}}x_{q}^{-4},\qquad\dot{\Omega}\sim\Omega^{11/3},$ (118)
recovering the standard GR scaling at high frequency. The transition between the two regimes is analytic and smooth, since the inspiral law depends on the single combination $\Gamma_{\rm d}+\Gamma_{\rm q}x_{q}$.
XII Scope and limitations
We have constructed an electromagnetic analogue of the post-Newtonian “Hamiltonian plus radiation–reaction” paradigm used in gravitational-wave physics. The conservative sector is truncated at the Darwin (11PN) level and dissipation is included through the leading 1.51.5PN dipole radiation–reaction force obtained via Landau–Lifshitz order reduction. Numerical examples employ parameter choices that make dissipative effects visible on accessible timescales; these do not alter the formal PN structure but are not intended for direct astrophysical calibration. Within this truncation, the canonical phase-space system exhibits monotonic energy loss, secular inspiral, and robust circularization, closely paralleling gravitational radiation reaction and suggesting structural universality in dissipative long-range dynamics.
We further explored inspiral laws for charged binaries in Einstein–Maxwell theory by combining the 22PN ADM-type conservative Hamiltonian with leading 1.51.5PN dipole dissipation. This yields gauge-invariant energy–frequency relations, analytic circular inspiral laws, and a dipole–quadrupole crossover scale separating electromagnetic and gravitational radiation dominance, all within a PN approximation. In special limits, the Landau–Lifshitz equations admit reductions to Painlevé transcendents [37, 38], and related universality structures have been conjectured for binary black-hole coalescence [39, 40, 41]. The present framework complements such approaches by providing an explicit, relativistic electromagnetic system with radiation reaction without reliance on special integrable limits.
XIII Conclusions and outlook
We have presented an explicit electromagnetic analogue of the PN “Hamiltonian plus radiation–reaction” framework widely used for compact-binary dynamics in GR and its ADM/EOB formulations [22, 27, 28, 29]. Radiation reaction is incorporated via Landau–Lifshitz order reduction of the Lorentz–Dirac self-force [7, 8, 9, 10, 13], yielding a causal second-order dynamics (see also [11, 12, 15]).
Our main result is a closed, directly implementable 11PN+1.5+1.5PN many-body phase-space system: conservative Darwin dynamics through $\mathcal{O}(c^{-2})$ supplemented by the leading near-zone dipole radiation–reaction force at $\mathcal{O}(c^{-3})$. In the conservative limit the Darwin Hamiltonian is preserved to numerical accuracy, while with radiation reaction the evolution exhibits monotonic energy loss and secular inspiral, in direct analogy with canonical PN radiation-reaction constructions in gravity [23, 24, 25, 26] and complementary to nonconservative action/EFT approaches [33, 34].
Specializing to binaries, we recover the characteristic dipole structure controlled by charge-to-mass asymmetry, including exact suppression when $q_{1}/m_{1}=q_{2}/m_{2}$ as in post-Coulombian analyses [35]. We derive analytic inspiral laws for circular and eccentric motion: a leading dipole chirp with $\dot{\Omega}\propto\Omega^{3}$ (and 11PN conservative corrections), and secular evolution for $(a,e)$ with an exactly integrable $a(e)$ relation and a closed-form time–eccentricity map. These results are supported by direct integration of the full phase-space equations, showing circularization and consistent radiative energy loss.
Embedding the discussion in Einstein–Maxwell theory, we employ the ADM-type CoM Hamiltonian through 22PN order and the leading 1.51.5PN dissipative acceleration from [36] to construct the gauge-invariant circular binding energy $E_{b}(x_{q})$ and a dipole-driven circular inspiral law through 22PN conservative order, in the same spirit as gauge-invariant PN/EOB diagnostics in GR [20, 22]. Including both electromagnetic dipole and gravitational quadrupole luminosities [26], we identify a gauge-invariant dipole–quadrupole crossover scale that depends only on the charge-to-mass asymmetry and yields a smooth transition from dipole-dominated evolution at low frequency to the standard GR-like quadrupole regime at high frequency.
The present work is intentionally limited to leading dipole dissipation and low-PN conservative dynamics. Natural extensions include higher-PN conservative terms [21, 22], higher-order radiation reaction [26, 33], and EOB-style resummations with consistent radiation-reaction prescriptions [27, 28, 30, 31, 32]. Beyond compact-binary motivations, the same explicit framework also provides a classical setup for radiation-reaction dynamics in regimes relevant to high-intensity laser experiments [17, 18, 19] and for exploring analytic structures (including Painlevé reductions) in dissipative long-range systems [37, 38, 39, 40, 41].
XIV Acknowledgments
RS thanks Suryateja Gavva, Anuradha Gupta, Sashwat Tanay and Leo C Stein for discussions on related problems in binary black hole dynamics. R.S is supported by DST INSPIRE Faculty fellowship, India (Grant No.IFA19-PH231). All authors acknowledge support from NFSG and OPERA Research Grant from Birla Institute of Technology and Science, Pilani (Hyderabad Campus).
XV Data Availability
The data that supports the findings of this study are available within the article.
Appendix A Runaway solution
Starting from the reduced one-dimensional Lorentz–Dirac equation,
$a(t)-t_{0}\,\dot{a}(t)=\frac{1}{m}F_{\mathrm{ext}}(t),\qquad t_{0}=\frac{2}{3}\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}m},$ (119)
we treat it as a first–order linear differential equation for the acceleration $a(t)$. Multiplying by the integrating factor $e^{-t/t_{0}}$ gives
$\frac{d}{dt}\!\left(e^{-t/t_{0}}a(t)\right)=-\frac{e^{-t/t_{0}}}{mt_{0}}\,F_{\mathrm{ext}}(t),$
which upon integration yields
$a(t)=e^{t/t_{0}}\!\left[a(0)-\frac{1}{mt_{0}}\!\int_{0}^{t}e^{-t^{\prime}/t_{0}}F_{\mathrm{ext}}(t^{\prime})\,dt^{\prime}\right].$
For a force that is switched on abruptly and then held constant, $F_{\mathrm{ext}}(t)=f\,\theta(t)$, the integral can be evaluated explicitly to give
$a(t)=e^{t/t_{0}}\left[b-\frac{f}{m}\left(1-e^{-t/t_{0}}\right)\theta(t)\right],$ (120)
where $b=a(0)$ is an integration constant and $\theta(t)$ is the Heaviside step function. The exponentially growing factor $e^{t/t_{0}}$ corresponds to the unphysical runaway mode, which is removed by the reduction–of–order procedure or by imposing suitable initial conditions that suppress it.
Appendix B Single–Particle Consistency Tests of Landau–Lifshitz Dynamics
(a) Trajectory in 3D.
(b) $r(\tau)$
(c) $x(\tau)$
(d) $y(\tau)$
(e) $z(\tau)$
Figure 2: Relativistic trajectory of a charged particle in a uniform magnetic field $\mathbf{B}=B\,\hat{\mathbf{z}}$, computed using the order–reduced Lorentz–Dirac equations. (a) Three–dimensional trajectory with initial (green) position. (b–e) Radius $r(\tau)$ measured from the relativistic guiding center and the spatial components $x(\tau)$, $y(\tau)$, and $z(\tau)$ at select time intervals. The motion shows damped cyclotron oscillations in the xxyy plane, while $z(\tau)$ stays constant for the chosen initial data. Parameters described in the main text.
In this section we numerically solve the order–reduced Lorentz–Dirac equations (Eq. 5) for a single charged particle in a uniform magnetic field, followed by three examples cases of increasing complexity. Unless stated otherwise, we choose the charge, mass, and permittivity to be $q=m=\varepsilon_{0}=1$. The radiation–reaction timescale $\tau_{0}=\frac{1}{4\pi\varepsilon_{0}}\!\left(\frac{2q^{2}}{3mc^{3}}\right)$ is evaluated using $c=30$ for the cyclotron example to enhance the dissipative effects over the integration interval. The particle is initially located at $\mathbf{r}_{0}=(0,0,0)$ with initial three–velocity $\mathbf{v}_{0}=(5,5,0)$, corresponding to a relativistic cyclotron orbit in the xxyy plane. We impose a uniform magnetic field $\mathbf{B}=(0,0,B_{c})$ with $B_{c}=10$, and no external electric field. The dynamical variables evolved are $(t(\tau),x(\tau),y(\tau),z(\tau))$ together with the four–velocity $u^{\mu}(\tau)$, integrated from proper time $\tau=0$ up to $\tau=\tau_{\mathrm{f}}=150$. These parameter choices produce long, slowly decaying cyclotron trajectories that illustrate the radiation–reaction–induced drift and the expected damping of the transverse motion, while the longitudinal motion remains constant due to the chosen initial conditions. As a consistency check, we verified that the instantaneous radiated power predicted by the Larmor formula matches the loss of kinetic energy obtained directly from the numerical solution of the order–reduced equations. Using the expression
$P_{\rm rad}(\tau)=\frac{2}{3}\,\frac{q^{2}}{4\pi\epsilon_{0}\,m^{2}c^{3}}\,\frac{dp_{\mu}}{d\tau}\frac{dp^{\mu}}{d\tau},$
with the proper–acceleration invariant $a^{2}(\tau)=a^{\mu}a_{\mu}$, we computed the total radiated energy $E_{\rm rad}=\int_{\tau_{\min}}^{\tau_{\max}}P_{\rm rad}(\tau)\,d\tau.$ For the parameter choices, the numerical integration yields
$E_{\rm rad}=1.51272,\qquad\Delta E_{\rm kin}=E_{\rm kin}(\tau_{\min})-E_{\rm kin}(\tau_{\max})=1.55527,$
corresponding to a fractional agreement $E_{\rm rad}/\Delta E_{\rm kin}=0.97.$ The small mismatch is expected and arises from finite–step integration error and the fact that radiation reaction is implemented through the order–reduced Landau–Lifshitz form rather than the full Lorentz–Dirac equation. For reference, the radiation–reaction timescale used in the simulation is
$\tau_{0}=\frac{1}{4\pi\varepsilon_{0}}\left(\frac{2q^{2}}{3mc^{3}}\right)=1.96488\times 10^{-6},$
which is much smaller than the dynamical timescale of the cyclotron orbit; this ensures that radiation damping accumulates slowly and provides a good numerical test of energy balance. The characteristic dynamical timescale of the motion is set by the cyclotron period. For a particle of charge $q$ and mass $m$ in a uniform magnetic field of magnitude $B_{c}$, the coordinate–time cyclotron frequency is $\omega_{\rm c}=qB_{c}/\gamma_{0}m$, so that
$T_{\rm c}^{(t)}=\frac{2\pi}{\omega_{\rm c}}=\frac{2\pi\gamma m}{qB_{c}}.$
With our choices $q=m=1$ and $B_{c}=10$ this gives
$T_{\rm c}^{(\tau)}\simeq 0.646534,$
while the corresponding coordinate–time period between successive peaks (as measured from the relativistically corrected center of the cyclotron orbit) is $T_{\rm c}^{(t)}=\gamma_{0}T_{\rm c}^{(\tau)}\simeq 0.646578$ for the initial Lorentz factor $\gamma_{0}\simeq 1.029$. Comparing with the radiation–reaction timescale
$\tau_{0}=\frac{1}{4\pi\varepsilon_{0}}\left(\frac{2q^{2}}{3mc^{3}}\right)=1.96488\times 10^{-6},$
we see that the radiation–reaction timescale is many orders of magnitude smaller than the orbital timescale, so that dissipation accumulates slowly over many revolutions. To quantify the radiative damping of the cyclotron orbit, we extract an effective decay timescale directly from the numerical trajectory. From the solution $x^{\mu}(\tau)$ we compute the instantaneous cyclotron radius $r(\tau)$ relative to the guiding center, and fit the resulting data to the exponential form
$r(\tau)\simeq A\,e^{-\beta\tau}+C,$ (121)
using a least–squares nonlinear model. The damping rate is then identified as $\beta$, giving the numerical timescale
$\tau_{\rm damp}^{\rm sim}\equiv\beta^{-1}.$ (122)
For comparison, the non-relativistic Landau–Lifshitz prediction for the cyclotron damping time,
$\tau_{\rm damp}^{\rm th}=\frac{1}{\tau_{0}\,\omega_{c}^{2}},\qquad\tau_{0}=\frac{2q^{2}}{3mc^{3}(4\pi\varepsilon_{0})},\qquad\omega_{c}=\frac{qB_{c}}{m},$ (123)
provides analytic benchmark. Evaluating these expressions for the parameters of the simulation yields $\tau_{\rm damp}^{\rm sim}\approx 4.40\times 10^{3}$ and $\tau_{\rm damp}^{\rm th}\approx 5.09\times 10^{3}$, in good quantitative agreement given the small-damping regime and the use of the non-relativistic theoretical estimate.
Case 1: Constant transverse $\mathbf{E}$ and $\mathbf{B}$ fields
(a) 3D trajectory
(b) $x(\tau)$
(c) $y(\tau)$
(d) $z(\tau)$
Case 2: Two Oscillating Electric and Magnetic Fields
(e) 3D trajectory
(f) $x(\tau)$
(g) $y(\tau)$
(h) $z(\tau)$
Case 3: Two crossed Electromagnetic Waves
(i) 3D trajectory
(j) $x(\tau)$
(k) $y(\tau)$
(l) $z(\tau)$
Figure 3: More examples: Landau–Lifshitz trajectories for three representative external-field configurations. Each row shows (left to right) the full 3D worldline projection followed by the coordinate evolution $x(\tau)$, $y(\tau)$, and $z(\tau)$.
B.1 More examples
To illustrate the structure of radiation–reacting worldlines predicted by the Landau–Lifshitz (LL) equation, we integrate the dynamics for three progressively more complex electromagnetic backgrounds, see Fig. 3 and Table 1. In all simulations the four–velocity is initialized relativistically as
$u^{\mu}=(\gamma c,\,\gamma v_{x},\,\gamma v_{y},\,\gamma v_{z}),$
with initial speeds chosen sufficiently small compared to the simulation light speed $c$. For the constant–field configurations we set $(v_{x},v_{y},v_{z})=(0,0,0)$. For the oscillatory and crossed–wave cases, symmetry is broken with small seed velocities
$v_{x}=c/100,\qquad v_{y}=c/1000,\qquad v_{z}=c/10,$
ensuring that the trajectory samples the full spatial structure of the fields. The time–dependent examples use amplitudes
$E_{0}=0.12\,c,\qquad B_{0}=43.$
Different numerical values of $c$ are used for stability in each configuration: $c=30$ for the cyclotron case, $c=5$ for Case 1, and $c=25$ for Cases 2 and 3.
The Cyclotron case (with $c=30$).
We begin with pure cyclotron motion in a uniform magnetic field $\mathbf{B}=(0,0,B_{0})$ and no electric field. This serves as a good test of LL damping in the absence of electric acceleration, and provides a reference point for the energy–balance diagnostics.
Case 1: Constant transverse fields (with $c=5$).
A uniform electric field $\mathbf{E}=(0,0,E_{0})$ together with a constant magnetic field $\mathbf{B}=(1,0,0)$ produces the standard transverse cycloidal trajectory with an $\mathbf{E}\!\times\!\mathbf{B}$ drift. This configuration serves as a test for validating LL damping in static external fields.
Case 2: Oscillating electric and magnetic fields (with $c=25$).
To probe LL radiation reaction in a time–dependent oscillating background, we impose independently oscillating fields
$\mathbf{E}(t)=\{E_{0}\sin[(\mathbf{k}\!\cdot\!\mathbf{r})-\omega t],0,0\},\qquad\mathbf{B}(t)=\{0,B_{0}\sin[(\mathbf{k}\!\cdot\!\mathbf{r})-\omega t],0\},$
with $\omega=c|\mathbf{k}|$. Because the amplitudes are unrelated, this configuration is not a plane electromagnetic wave For the oscillating-field setup we adopt the amplitudes $E_{0}=0.12\,c$ for the electric field and $B_{0}=43$ for the magnetic field.
Case 3: Two crossed electromagnetic waves (with $c=25$).
We next superpose two oscillatory fields with wavevectors $\mathbf{k}_{1}=(0,0,1/25)$ and $\mathbf{k}_{2}=(0,1/25,0)$ and corresponding frequencies $\omega_{1,2}=c|\mathbf{k}_{1,2}|$. Using the same amplitudes $E_{0}$ and $B_{0}$ as in Case 2, the resulting field has genuine three–dimensional interference, producing LL trajectories with multiscale drifts, sharp oscillations, and direction reversals characteristic of strong–field radiation–reacting motion. For each configuration we compute: (i) the relativistic kinetic energy
$K(\tau)=mc^{2}(\gamma(\tau)-1),$
(ii) the work done by the external fields,
$W_{\rm ext}(\tau)=\int_{\tau_{0}}^{\tau}q\,\mathbf{E}_{\rm ext}\!\cdot\!\mathbf{v}\,d\sigma,$
and (iii) the radiated energy,
$E_{\rm rad}(\tau)=\int_{\tau_{0}}^{\tau}P_{\rm cov}(\sigma)\,d\sigma,$
where $P_{\rm cov}$ is the covariant LL Larmor power. Energy balance requires
$\Delta K=W_{\rm ext}-E_{\rm rad},$
and we quantify this balance using the parameter
$R=\frac{|\Delta K-W_{\rm ext}|}{E_{\rm rad}}.$
Across the four simulations the LL energy balance agrees at a satisfactory numerical level:
$R_{\rm cyclotron}\simeq 1.03,\qquad R_{\rm Case\,1}\simeq 0.95,\qquad R_{\rm Case\,2}\simeq 1.00,\qquad R_{\rm Case\,3}\simeq 1.09.$
Even in the multidimensional, strongly nonlinear crossed–wave configuration (Case 3), the energy consistency remains within $\sim 8\%$, confirming that the Landau–Lifshitz order–reduced dynamics is energetically reliable across static, time–dependent, and fully multidimensional forcing environments.
Table 1: Energy–balance diagnostics for the cyclotron orbit and the three external–field configurations studied in Sec. III. Here $\Delta K=|K(t_{f})-K(0)|$, $W_{\rm ext}=\left|\int_{0}^{t_{f}}d\tau\,q\,\mathbf{E}\!\cdot\!\mathbf{v}\right|$, $E_{\rm rad}$ is the covariant radiated energy, and $R\equiv|\Delta K-W_{\rm ext}|/E_{\rm rad}$ tests the LL energy–balance relation.
Case $\Delta K$ $W_{\rm ext}$ $E_{\rm rad}$ $R$
Case 0: The Cyclotron case 1.55459 0 1.51272 1.02768
Case 1: Constant transverse $E$ and $B$ 0.00583564 0.005916 0.0000841766 0.954583
Case 2: Oscillating electric & magnetic fields 0.472655 0.296402 0.175510 1.00423
Case 3: Two crossed electromagnetic waves 39.5982 43.4668 3.55509 1.08818
Appendix C Details on the Darwin force correction at 1PN
In this section, we show the detailed steps leading to the 1PN conservative force in Eq. (17) of the main text. We start from the Darwin interaction Hamiltonian, along with the useful notations as presented below:
$H_{D,\mathrm{int}}=-\frac{1}{2}\sum_{a\neq b}\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}m_{a}m_{b}c^{2}}\;\frac{1}{r_{ab}}\,\mathcal{P}_{ab},\qquad\mathbf{r}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b},\quad r_{ab}=|\mathbf{r}_{ab}|$ (124)
where
$\mathcal{P}_{ab}:=\mathbf{p}_{a}\!\cdot\!\mathbf{p}_{b}+\frac{(\mathbf{p}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}}.$ (125)
Using the symmetry under $a\leftrightarrow b$ we can write
$H_{D,\mathrm{int}}=-\sum_{a<b}H_{ab},\qquad H_{ab}:=\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}m_{a}m_{b}c^{2}}\;\frac{1}{r_{ab}}\,\mathcal{P}_{ab}.$ (126)
Within 1PN accuracy we may replace $\mathbf{p}_{a}\simeq m_{a}\mathbf{v}_{a}$ and $\mathbf{p}_{b}\simeq m_{b}\mathbf{v}_{b}$, which yields
$\mathcal{P}_{ab}=m_{a}m_{b}\left[\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+\frac{(\mathbf{v}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{v}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}}\right].$ (127)
The masses thus cancel in the full expression, and with $k=1/(4\pi\varepsilon_{0})$ we obtain
$H_{ab}=-\frac{k\,q_{a}q_{b}}{2c^{2}}\,\frac{1}{r_{ab}}\left[\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+\frac{(\mathbf{v}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{v}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}}\right].$ (128)
The Darwin contribution to the canonical equation for $\mathbf{p}_{a}$ is thus
$\dot{\mathbf{p}}_{a}\Big|_{1\mathrm{PN}}=-\,\frac{\partial H_{D,\mathrm{int}}}{\partial\mathbf{x}_{a}}=-\sum_{b\neq a}\frac{\partial H_{ab}}{\partial\mathbf{x}_{a}}.$ (129)
Since $H_{ab}$ depends on $\mathbf{x}_{a}$ only through $\mathbf{r}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b}$, we may replace
$\frac{\partial}{\partial\mathbf{x}_{a}}\;=\;\frac{\partial}{\partial\mathbf{r}_{ab}}\equiv\bm{\nabla}_{\mathbf{r}},\qquad\mathbf{r}\equiv\mathbf{r}_{ab},\quad r\equiv r_{ab},\quad\hat{\mathbf{r}}\equiv\hat{\mathbf{r}}_{ab}=\frac{\mathbf{r}}{r}.$ (130)
For a single pair $(a,b)$, Eq. (128) becomes
$H_{ab}=-A\Big[S_{1}(\mathbf{r})+S_{2}(\mathbf{r})\Big],\qquad A:=\frac{k\,q_{a}q_{b}}{2c^{2}},$ (131)
with the associated quantities
$S_{1}(\mathbf{r}):=\frac{\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}}{r},\qquad S_{2}(\mathbf{r}):=\frac{(\mathbf{v}_{a}\!\cdot\!\mathbf{r})(\mathbf{v}_{b}\!\cdot\!\mathbf{r})}{r^{3}}.$ (132)
We can now write
$\dot{\mathbf{p}}_{a}^{(ab)}\Big|_{1\mathrm{PN}}=-\frac{\partial H_{ab}}{\partial\mathbf{x}_{a}}=-\bm{\nabla}_{\mathbf{r}}H_{ab}=A\,\bm{\nabla}_{\mathbf{r}}\big[S_{1}(\mathbf{r})+S_{2}(\mathbf{r})\big].$ (133)
Since $\mathbf{v}_{a}$ and $\mathbf{v}_{b}$ are independent of $\mathbf{r}$, we have
$\bm{\nabla}_{\mathbf{r}}S_{1}=(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b})\,\bm{\nabla}_{\mathbf{r}}\!\left(\frac{1}{r}\right)=(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b})\left(-\frac{\hat{\mathbf{r}}}{r^{2}}\right)=-\frac{\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}}{r^{2}}\,\hat{\mathbf{r}}$. (134)
We now define two intermediate scalars
$\alpha:=\mathbf{v}_{a}\!\cdot\!\mathbf{r},\qquad\beta:=\mathbf{v}_{b}\!\cdot\!\mathbf{r}.$ (135)
Then $S_{2}=\alpha\beta/r^{3}$ and
$\bm{\nabla}_{\mathbf{r}}S_{2}=\frac{1}{r^{3}}\bm{\nabla}_{\mathbf{r}}(\alpha\beta)+\alpha\beta\,\bm{\nabla}_{\mathbf{r}}\!\left(\frac{1}{r^{3}}\right).$ (136)
Because $\bm{\nabla}_{\mathbf{r}}\alpha=\mathbf{v}_{a}$ and $\bm{\nabla}_{\mathbf{r}}\beta=\mathbf{v}_{b}$, we have
$\bm{\nabla}_{\mathbf{r}}(\alpha\beta)=\beta\,\mathbf{v}_{a}+\alpha\,\mathbf{v}_{b}=(\mathbf{v}_{b}\!\cdot\!\mathbf{r})\,\mathbf{v}_{a}+(\mathbf{v}_{a}\!\cdot\!\mathbf{r})\,\mathbf{v}_{b}.$ (137)
Using the standard result
$\bm{\nabla}_{\mathbf{r}}\!\left(\frac{1}{r^{3}}\right)=-3\,\frac{\mathbf{r}}{r^{5}}=-\frac{3}{r^{4}}\,\hat{\mathbf{r}}$, (138)
we thus obtain
$\bm{\nabla}_{\mathbf{r}}S_{2}=\frac{1}{r^{3}}\Big[(\mathbf{v}_{b}\!\cdot\!\mathbf{r})\,\mathbf{v}_{a}+(\mathbf{v}_{a}\!\cdot\!\mathbf{r})\,\mathbf{v}_{b}\Big]-3\,\frac{(\mathbf{v}_{a}\!\cdot\!\mathbf{r})(\mathbf{v}_{b}\!\cdot\!\mathbf{r})}{r^{5}}\,\mathbf{r}.$ (139)
In order to reach Eq. (17) of the main text, it is convenient to express the scalar products with $\hat{\mathbf{r}}$:
$\mathbf{v}_{a}\!\cdot\!\mathbf{r}=r\,(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a}),\qquad\mathbf{v}_{b}\!\cdot\!\mathbf{r}=r\,(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b}).$ (140)
Using these, the previous expression becomes
$\bm{\nabla}_{\mathbf{r}}S_{2}$ $=\frac{1}{r^{3}}\Big[r(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})\,\mathbf{v}_{a}+r(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})\,\mathbf{v}_{b}\Big]-3\,\frac{r^{2}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})}{r^{5}}\,\mathbf{r}$
$=\frac{1}{r^{2}}\Big[\mathbf{v}_{a}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})-3\,\hat{\mathbf{r}}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})\Big].$ (141)
Adding Eqs. (134) and (141) we find
$\bm{\nabla}_{\mathbf{r}}\big(S_{1}+S_{2}\big)$ $=-\frac{\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}}{r^{2}}\,\hat{\mathbf{r}}$
$\quad+\frac{1}{r^{2}}\Big[\mathbf{v}_{a}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})-3\,\hat{\mathbf{r}}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})\Big]$
$=\frac{1}{r^{2}}\Big[\mathbf{v}_{a}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})-\hat{\mathbf{r}}\big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}_{b})\big)\Big].$
Therefore the Darwin contribution to $\dot{\mathbf{p}}_{a}$ from particle $b$ is
$\dot{\mathbf{p}}_{a}^{(ab)}\Big|_{1\mathrm{PN}}=\frac{k\,q_{a}q_{b}}{2c^{2}r_{ab}^{2}}\Big[\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})-\hat{\mathbf{r}}_{ab}\big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\big)\Big].$ (142)
Summing over $b\neq a$ gives the full Darwin 1PN correction
$\dot{\mathbf{p}}_{a}\Big|_{1\mathrm{PN}}=k\sum_{b\neq a}\frac{q_{a}q_{b}}{2c^{2}r_{ab}^{2}}\Big[\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})-\hat{\mathbf{r}}_{ab}\big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\big)\Big],$ (143)
which is Eq. (17) in the main text. The Darwin 1PN correction has a simple physical interpretation. The terms proportional to $\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})$ and $\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})$ represent the magnetostatic, or current–current, interaction between moving charges: each particle experiences a velocity–dependent force generated by the motion of the other charge. The remaining contribution, proportional to $\hat{\mathbf{r}}_{ab}\big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\big)$, ensures that the interaction has the unique Poincaré–consistent form required at order $v^{2}/c^{2}$.
The velocity–dependent conservative 1PN force derived in Eq. (17) of the main text is the classical, spin–independent component of the Darwin interaction. It represents the leading relativistic correction to Coulomb’s law obtained by expanding the full electromagnetic interaction between two charges to order $v^{2}/c^{2}$. This correction encodes conservative retardation effects in the near zone and admits three equivalent interpretations.
First, starting from classical electrodynamics, the retarded scalar and vector potentials of a point charge $q_{b}$ moving with velocity $\mathbf{v}_{b}$ are
$\Phi_{b}(\mathbf{x},t)=k\,\frac{q_{b}}{R_{b}-\mathbf{R}_{b}\!\cdot\!\bm{\beta}_{b}},\qquad\mathbf{A}_{b}(\mathbf{x},t)=\frac{k\,q_{b}}{c}\,\frac{\mathbf{v}_{b}}{R_{b}-\mathbf{R}_{b}\!\cdot\!\bm{\beta}_{b}},$ (144)
where $\mathbf{R}_{b}=\mathbf{x}-\mathbf{x}_{b}(t_{\rm ret})$ and $\bm{\beta}_{b}=\mathbf{v}_{b}/c$. Expanding in the near zone for small velocities and small retardation, and retaining terms through order $v^{2}/c^{2}$, one finds schematically
$\frac{1}{R_{b}-\mathbf{R}_{b}\!\cdot\!\bm{\beta}_{b}}=\frac{1}{r_{ab}}\left[1+\frac{\mathbf{v}_{b}\!\cdot\!\mathbf{r}_{ab}}{c\,r_{ab}}+\frac{v_{b}^{2}}{2c^{2}}-\frac{(\mathbf{v}_{b}\!\cdot\!\mathbf{r}_{ab})^{2}}{2c^{2}r_{ab}^{2}}+\mathcal{O}(c^{-3})\right].$ (145)
Constructing the electric field $\mathbf{E}_{b}=-\nabla\Phi_{b}-\partial\mathbf{A}_{b}/\partial t$, the field of particle $b$ evaluated at the position of particle $a$ contains, in addition to the Coulomb term, a velocity–dependent 1PN contribution proportional to
$\frac{kq_{b}}{c^{2}r_{ab}^{2}}\Big[\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})+\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})-\hat{\mathbf{r}}_{ab}(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b})-3\hat{\mathbf{r}}_{ab}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\Big].$
Multiplication by $q_{a}$ reproduces exactly the Darwin force in Eq. (17). Thus, the Darwin interaction corresponds to the $v^{2}/c^{2}$ expansion of the full retarded electromagnetic interaction when radiation emission is neglected.
Second, the same interaction arises naturally from an effective–action viewpoint. After integrating out the electromagnetic field (e.g. in Coulomb gauge) and expanding the near–zone effective action to order $v^{2}/c^{2}$, one obtains an instantaneous current–current interaction,
$S=-\sum_{a}m_{a}c^{2}\!\int\!dt\,\sqrt{1-\frac{v_{a}^{2}}{c^{2}}}+\frac{1}{8\pi k}\!\int\!d^{3}x\,d^{3}x^{\prime}\frac{\mathbf{J}(\mathbf{x},t)\!\cdot\!\mathbf{J}(\mathbf{x}^{\prime},t)}{|\mathbf{x}-\mathbf{x}^{\prime}|}+\cdots,$ (146)
where $\mathbf{J}$ is the charge current. Expanding the matter action to 1PN order and substituting the point–particle current $\mathbf{J}(\mathbf{x},t)=\sum_{a}q_{a}\mathbf{v}_{a}\,\delta^{3}(\mathbf{x}-\mathbf{x}_{a})$ yields the Darwin Lagrangian
$L_{\rm Darwin}=\frac{k}{2c^{2}}\sum_{a\neq b}\frac{q_{a}q_{b}}{r_{ab}}\Big[\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+(\mathbf{v}_{a}\!\cdot\!\hat{\mathbf{r}}_{ab})(\mathbf{v}_{b}\!\cdot\!\hat{\mathbf{r}}_{ab})\Big].$ (147)
The corresponding Hamiltonian (via a Legendre transform) coincides with $H_{D,\mathrm{int}}$, and differentiation with respect to $\mathbf{x}_{a}$ reproduces Eq. (17), making explicit that the Darwin term originates purely from the instantaneous, non–radiative sector of the electromagnetic field.
Third, in quantum electrodynamics the spin–independent part of the Breit Hamiltonian is
$H_{\rm Breit}=\frac{kq_{a}q_{b}}{2m_{a}m_{b}c^{2}r_{ab}}\left[\mathbf{p}_{a}\!\cdot\!\mathbf{p}_{b}+\frac{(\mathbf{p}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}}\right].$ (148)
which has exactly the same operator structure as the classical Darwin Hamiltonian. In the $\hbar\!\to\!0$ limit, identifying $\mathbf{p}_{a}=m_{a}\mathbf{v}_{a}$ reduces this expression to the interaction (128), and hence to the same equations of motion.
In all three approaches, the resulting velocity–dependent force takes the universal form
$\dot{\mathbf{p}}_{a}\Big|_{1\mathrm{PN}}\propto\mathbf{v}_{a}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})+\mathbf{v}_{b}(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})-\hat{\mathbf{r}}_{ab}\Big(\mathbf{v}_{a}\!\cdot\!\mathbf{v}_{b}+3(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{a})(\hat{\mathbf{r}}_{ab}\!\cdot\!\mathbf{v}_{b})\Big),$ (149)
demonstrating that the Darwin interaction is the complete conservative correction to the classical two–body problem at order $v^{2}/c^{2}$, coinciding with the standard Darwin/Breit interaction in both classical and quantum electrodynamics.
Appendix D Simulation results: the Charge Neutral Binary
D.1 The conservative sector
(a) Trajectories.
(b) Darwin Hamiltonian evolution.
Figure 4: Two oppositely charged blobs, each containing ten particles, evolved with the conservative Darwin Hamiltonian at 1PN order. Panel (a) shows the three–dimensional trajectories (blue: positive charges; black: negative charges), and panel (b) the corresponding evolution of the Darwin Hamiltonian $H(\tau)$.
(a) 3D trajectories with initial (green) and final (red) markers.
(b) Parametric projection in the xxyy plane.
(c) $x(\tau)$ vs. $\tau$.
(d) $y(\tau)$ vs. $\tau$.
(e) $z(\tau)$ vs. $\tau$.
(f) Hamiltonian $H(\tau)$ showing radiative energy loss.
Figure 5: Charge neutral binary of same mass starting from diametrically opposite locations with equal and opposite momenta, with radiation reaction from the truncated post Newtonian approach: (a) Full 3D trajectories; (b) parametric xxyy plane; (c–e) components $x(\tau)$, $y(\tau)$, $z(\tau)$; (f) Hamiltonian $H(\tau)$, decreasing due to radiation. The initial particle positions are shown in green, and the final positions at $\tau=\tau_{f}$ are shown in red.
(a) 3D trajectories with initial (green) and final (red) markers.
(b) Parametric projection in the xxyy plane.
(c) $x(\tau)$ vs. $\tau$.
(d) $y(\tau)$ vs. $\tau$.
(e) $z(\tau)$ vs. $\tau$.
(f) Hamiltonian $H(\tau)$ showing radiative energy loss.
Figure 6: Charge neutral binary of extreme mass ratio, one heavy and one light (with circular orbit initial conditions), with radiation reaction from the truncated post Newtonian approach: (a) Full 3D trajectories; (b) parametric xxyy plane; (c–e) components $x(\tau)$, $y(\tau)$, $z(\tau)$; (f) Hamiltonian $H(\tau)$, decreasing due to radiation. The initial particle positions are shown in green, and the final positions at $\tau=\tau_{f}$ are shown in red.
(a) 3D trajectories with initial (green) and final (red) markers.
(b) Parametric projection in the xxyy plane.
(c) Hamiltonian $H(\tau)$ showing radiative energy loss.
(d) $x(\tau)$ vs. $\tau$.
(e) $y(\tau)$ vs. $\tau$
(f) $\frac{\mathcal{I}}{\mathcal{I}_{0}}$ vs No. of orbits
Figure 7: Charge neutral binary, one heavy and one light, with elliptic orbit initial conditions. Radiation reaction from the truncated post Newtonian approach shows clear signature of circularization: (a) Full 3D trajectories; (b) parametric xxyy plane; (c) Hamiltonian $H(\tau)$, decreasing due to radiation (d-e) components $x(\tau)$, $y(\tau)$ and (f) plot showing $\mathcal{I}=a(1-e^{2})/e^{4/3}$ is preserved to sub-percent accuracy, with $\mathcal{O}(1/c^{2})$ corrections as expected . The initial particle positions are shown in green, and the final positions at $\tau=\tau_{f}$ are shown in red.
As a stringent validation of the conservative sector of our framework, we consider a system of two spatially separated three–dimensional charge blobs of opposite sign, evolved using the Darwin Hamiltonian alone, with all radiation–reaction terms explicitly switched off. Each blob consists of multiple point charges initialized within a compact spherical region, with the two charge clouds separated by a distance large compared to their individual sizes.
Figure 4(a) displays the resulting trajectories. Positively charged particles are shown in blue and negatively charged particles in black, with filled markers indicating the initial positions. The evolution is fully conservative, with the velocity–dependent Darwin interaction producing nontrivial many–body dynamics while preserving the overall phase–space structure.
The corresponding Hamiltonian $H(\tau)$ is shown in Fig. 4(b). As expected for purely conservative dynamics, no secular drift is observed. The small bounded fluctuations visible in $H(\tau)$ arise from unavoidable numerical stiffness during close encounters between particles and provide a diagnostic of numerical accuracy. The absence of any systematic energy loss confirms that the Darwin sector has been implemented consistently and provides a reliable starting point for the inclusion of dissipative radiation–reaction effects in the next section. We now present representative solutions of the dissipative 1.5PN accurate NN–body framework Eq. (16). Throughout this section we use $\tau$ to denote the ordinary coordinate time (not the proper time). The conservative interactions are evolved with the Darwin Hamiltonian, retaining all terms through 1PN order, $\mathcal{O}(1/c^{2})$, while dissipation is incorporated as a non-Hamiltonian Landau–Lifshitz radiation–reaction force at leading 1.5PN order, $\mathcal{O}(1/c^{3})$, implemented via order reduction. This canonical-plus-dissipative framework allows us to test how radiative losses modify otherwise integrable Coulombic motion, producing secular inspiral and circularization. The examples below progress from a symmetric equal-mass neutral binary to an extreme-mass-ratio “hydrogen-like” configuration and finally to an initially eccentric orbit that circularizes under radiation reaction; in each case we monitor trajectories and the Hamiltonian $H(\tau)$ to quantify the cumulative energy loss.
D.2 Charge neutral binary of same mass
To validate our construction of the conservative plus dissipative system governed by Eq. (16), we first consider a symmetric charge-neutral bound configuration consisting of two point charges of equal mass $m=1$ and opposite charge $\pm q$, interacting via an attractive Coulomb potential V(r)=k/rV(r)=-k/r. The particles are initially placed at diametrically opposite positions,
$\mathbf{r}_{1}(0)=(1,0,0),\qquad\mathbf{r}_{2}(0)=(-1,0,0),$
so that the interparticle separation is $r=2R=2$. For a circular orbit the centripetal balance condition,
$\frac{mv^{2}}{R}=\frac{\alpha}{(2R)^{2}},$
requires a tangential speed $v=\tfrac{1}{2}$ for $m=k=R=1$. We therefore choose the initial momenta
$\mathbf{p}_{1}(0)=(0,\,0.5,\,0),\qquad\mathbf{p}_{2}(0)=(0,\,-0.5,\,0),$
which supply the equal and opposite tangential velocities needed to sustain the circular motion. The corresponding total energy,
$E_{\rm tot}=mv^{2}-\frac{k}{2R}=-\,\frac{1}{4},$
is negative as expected for a bound nonrelativistic Coulomb orbit. These initial data are chosen so that the nonrelativistic Coulomb dynamics would support a circular orbit of radius unity. When the Landau–Lifshitz self–force is included, however, the system no longer admits an exact periodic solution: radiation reaction steadily removes mechanical energy, producing a gradual reduction in orbital radius and an associated inward spiral, as shown in Fig. 5. To make the dissipative dynamics visible on numerically tractable time scales, we deliberately choose $c$ such that the leading radiation– reaction force, scaling as $\mathcal{O}(1/c^{3})$, is suppressed by only $\sim 10^{-2}$ relative to the conservative 1PN corrections $\mathcal{O}(1/c^{2})$. This choice does not affect the formal consistency of the PN expansion, but simply rescales the inspiral time.
The six–panel figure displays the full three-dimensional worldlines, the parametric projection in the xxyy plane, and the individual coordinate functions $x(\tau)$, $y(\tau)$, and $z(\tau)$. The final panel shows the monotonic decrease of the Hamiltonian $H(\tau)$, which directly reflects the radiative energy loss predicted by our order-reduced evolution. For clarity, the initial particle positions are rendered in green, while their final locations at $\tau=\tau_{f}$ are shown in red, making the cumulative orbit shrinkage visually explicit. This two-body example provides a clean test of the LL radiation–reaction scheme and illustrates how even a perfectly symmetric bound configuration evolves irreversibly once electromagnetic self-interaction is taken into account.
D.3 Charge neutral binary of extreme mass ratio
As a second application of our N–body framework Eq. (16), we consider an unequal–mass, oppositely charged system designed to mimic a classical “hydrogen–like” configuration. For a pair of point charges $(q_{1},m_{1})$ and $(q_{2},m_{2})$ interacting through the Coulomb potential, a circular orbit may be constructed by imposing the usual centripetal balance in the relative two–body system. Denoting by $\mathbf{r}=\mathbf{x}_{2}-\mathbf{x}_{1}$ the separation vector, the magnitude of the Coulomb force is $F_{C}=k\,|q_{1}q_{2}|/r^{2}$ with $k=1/(4\pi\varepsilon_{0})$. In the centre–of–mass frame the relative coordinate evolves as a particle of reduced mass $\mu=m_{1}m_{2}/(m_{1}+m_{2})$ in this central potential. A circular orbit of radius $r_{0}$ requires
$\frac{\mu v_{\rm rel}^{2}}{r_{0}}=k\,\frac{|q_{1}q_{2}|}{r_{0}^{2}},\qquad\Rightarrow\qquad v_{\rm rel}=\sqrt{\,k\,\frac{|q_{1}q_{2}|}{\mu r_{0}}\,}$, (150)
where $v_{\rm rel}$ is the relative orbital speed. One convenient choice of initial data is to place the two charges on the xx–axis at
$\mathbf{x}_{1}(0)=-\frac{m_{2}}{m_{1}+m_{2}}\,(r_{0},0,0),\qquad\mathbf{x}_{2}(0)=\frac{m_{1}}{m_{1}+m_{2}}\,(r_{0},0,0),$ (151)
with momenta
$\mathbf{p}_{1}(0)=-\mathbf{p}_{2}(0),\qquad|\mathbf{p}_{2}(0)|=\mu~v_{\rm rel},$ (152)
and the velocity of particle 2 chosen along the +y+y direction to produce a counterclockwise orbit in the $xy$ plane. In the heavy–nucleus limit $m_{1}\gg m_{2}$ these expressions reduce to the familiar hydrogenic initial data $\mathbf{x}_{1}(0)=(0,0,0)$, $\mathbf{x}_{2}(0)=(r_{0},0,0)$, and $\mathbf{p}_{2}(0)=(0,\,\sqrt{m_{2}k|q_{1}q_{2}|/r_{0}},\,0)$. We choose charges $q_{1}=+1$, $q_{2}=-1$ and masses $m_{1}\gg m_{2}$ so that particle 1 acts as a heavy Coulomb center, while particle 2 executes a bound orbit. This initialization guarantees that, in the absence of radiation reaction, the system would remain on a circular Coulomb orbit. When LL radiation reaction is included, however, the light particle gradually spirals inward as it continuously loses energy to electromagnetic radiation. Once again, for numerical illustration, we choose the speed of light $c$ to be moderately small so that the radiation–reaction force, which enters at $\mathcal{O}(1/c^{3})$, is only suppressed by a factor $\sim 10^{-3}$ relative to the conservative 1PN terms; this exaggerates dissipative effects and allows the inspiral to be resolved on accessible time scales without altering the underlying PN ordering.
Figure 6 shows the resulting inspiral. Panel (a) displays the full 3D trajectories, with green and red markers denoting the initial and final positions. The parametric projection in the xxyy plane (b) makes the secular inward drift clearly visible. The coordinate components $x(\tau)$, $y(\tau)$, and $z(\tau)$ plotted in panels (c)–(e) exhibit both the fast orbital oscillations and the slow amplitude decay associated with radiative backreaction. Finally, the Hamiltonian $H(\tau)$ in panel (f) shows a clean and monotonic decrease, confirming that the truncated post–Newtonian LL scheme captures the expected secular energy loss. This hydrogen–like example illustrates how the LL self–force naturally produces inspiral dynamics in a bound Coulomb system, providing a classical analogue of radiation–reaction–driven orbital decay.
D.4 Charge neutral binary of comparable mass showing circularization of initial elliptic orbits
In Fig. 1 we examine a two–body configuration demonstrating the radiative circularization predicted by the truncated post-Newtonian scheme Eq. (16). We consider oppositely charged particles with masses $(m_{1},m_{2})=(5,1)$ initialized in the center–of–mass frame at a separation $r_{0}=1$,
$\mathbf{r}_{1}(0)=(0,0,0),\qquad\mathbf{r}_{2}(0)=\frac{m_{1}}{m_{1}+m_{2}}(r_{0},0,0),$
and assign equal–magnitude and opposite momenta $\mathbf{p}_{1}(0)=-\mathbf{p}_{2}(0)$ with $|\mathbf{p}_{1}(0)|=\mu\,v_{\rm rel}$,
$v_{\rm rel}=\sqrt{\frac{k\,|q_{1}q_{2}|}{\mu\,r_{0}}},\qquad\mu=\frac{m_{1}m_{2}}{m_{1}+m_{2}},$
so that, in the absence of radiation reaction, the orbit is elliptic. When evolved in our scheme governed by Eq. (16), the orbit gradually shrinks and circularizes, as seen in the parametric xxyy projection in Fig. 1(b). Panels (c)–(e) display the individual coordinate components $x(\tau)$, $y(\tau)$, and $z(\tau)$, showing the expected modulation of the radial amplitude and no out–of–plane drift. The Hamiltonian $H(\tau)$ in Fig. 1(f) decreases monotonically, matching the cumulative Larmor power by construction. As with earlier examples, the initial particle positions are marked in green and their final locations at $\tau=\tau_{f}$ are indicated in red.